A Complete Bibliography of Bell Labs Technical Journal, 2000–2009

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

14 October 2017 Version 1.08

Title word cross-reference

80 × 107 [SLJ^{+}09].

0 [Bi05].

1 [SLJ^{+}09]. 100G [HSS^{+}09]. 100Gb [WHK^{+}09]. 100Gb/s [WHK^{+}09]. 17799/27001 [MBC^{+}07]. 1x [LS07, Bi05, BVYY07, BCV^{+}07, CDM^{+}07, CP07, DHJ^{+}07, MWZ07]. 1xEV [BBC^{+}02, BCG^{+}03]. 1xEV-DO [BBC^{+}02, BCG^{+}03].

2001 [LGTM02].

3g [CMCP06, BCG^{+}03, BRG^{+}05, FK04, JGA04, RLP05, UVMES^{+}04]. 3G1X [BCL^{+}02]. 3GPP [BGDT07, vdGHM04]. 3GPP-based [vdGHM04].

40Gb [WHK^{+}09]. 40Gb/s [WHK^{+}09]. 4G [HSP03, KAGS09, RS09, SR09, SBEW09]. 4th [KGT06].

5ESS(R) [BHH^{+}00].

6 [LGTM02].

7-nines [McP06]. 7R [ME00]. 7R/E^{T M} [ME00].

8 [SLJ^{+}09]. 802.11 [DJR07, RLP05]. 802.11-based [DJR07]. 802.11/3G [RLP05]. 802.16e [BCC^{+}07, BCG^{+}09, DKR^{+}06]. 802.16e-based [BCC^{+}07, BCG^{+}09].

911 [AMBK^{+}07]. 911-NOW [AMBK^{+}07]. 95 [Ahm02].

Absence [Dan08]. accelerated [AJK06]. access [BDSS07, BJM^{+}02, CMMZ08, CMPW02, Cla09, DGH^{+}02, DLM^{+}02, De 09, FYJB09, GMW07, HP08, HPvV^{+}08, LBF^{+}04, Maz06, NO08, OVN08, PS07, PRS01, PZBT03, SR09, SLW^{+}09, TBVV08, Ton02, TB04, UGDS01, VFRO08, WvLS08]. access-independent [BDSS07]. Accounting [FHW01]. Accurately [AGK^{+}03]. Achieving [Koc07, YJY06, McP06]. across [AEM^{+}06, KTW00, LHMS05, LBF^{+}04, OC08, TB08, ŽBLvB05]. action [Gol05]. active [PUV07]. activity [Ded03]. adapter [Wan00]. Adapting [JGA04]. Adaptive [ABBE08, Rad04, RTM^{+}04, Lau00]. added [ARS^{+}06, RTM^{+}04]. addressing [DMO00, Koc02]. administrative [NS03]. admission [ABF^{+}07, HKUW03, OVN08]. Advanced [CCH^{+}01, IBPF08, OJ04, DEK^{+}05, IGNC^{+}05, KRA06a, MBB00]. Advances [GDS^{+}00, Koc02, Kol09, KSO03, QW05, CGJ^{+}02, KP09]. advantage [BBTD05]. Advantages [Maz06]. advertisement [ZMS08]. aeronautics [Tri02]. against [DDL08, HLS09, WRL^{+}09]. agent [Sun06, VW08]. agents [AG04]. aggregated [Rao06]. aggregation [HHVST07]. agile [Cho04b]. agreements [Har05, MvHdMvE03]. aided [BYA09]. air [BGDT07]. Alcatel [Ano07a, Ano07b, Ano07c, Ano07d, Ano08a, Ano08b, Ano08c, Ano08d, Ano09a, Ano09b, Ano09c, Ano09d, BDE^{+}07, RCB^{+}09]. alerting [Fre06]. algorithm [HKUW03, JLN03]. algorithms [AMG05, Ahm02, BGW09, CKR05, Nak03, ŽKN^{+}08]. all-IP [SJM^{+}04]. Alliance [BGTU05].

Alternative [LS07]. aluminum [CMC^{+}05]. Ambient [TGC^{+}08]. American [KT07, Tri02]. among [LUV08]. Analysis [EM08, Rao06, Tsa00, BCC^{+}07, BCG^{+}09, BCL^{+}02, BCSW09, Bru05, CLOU06, FDC^{+}07, HJvO^{+}06, JLD^{+}06, LUV08, Mac04, PSST07, TM05, UVMES06, vVWBH05]. analytics [BH08]. analyze [GM07]. Analyzing [MA06, ABC^{+}05, BF06]. ANIQUE [KT07]. antenna [BYA09]. antennas [LCM^{+}02]. API [GGHH^{+}03]. APIs [BMOP00]. Application [SBSS01, ABF^{+}07, BVYY07, CCC^{+}02, DSW08, Flo06, GA00, GM07, JGA04, JLN03, Lar08, LS01, MHH05, Rad04, Yan02]. Application-level [SBSS01]. Applications [AH08, CRVX02, BMvB06, BCV09, BCD^{+}08, BYA09, CGLU03, GFB05, HV06, KKL^{+}02, Lab08, LDFK05, MGS^{+}05, MBB00, NO08, RKKM05, SSM^{+}00, Sie01, SRR08, SBLD08, TB04, UGDS01, Vla03, UVMES06, WXT^{+}08]. applied [Ram07]. Applying [HSP03, KM03, RHSG03]. Approach [Mac04, Aiz05, BBK^{+}09, CMCP06, CMMZ08, CRVX02, HHS07, HLS09, KRA06a, Moo04, SMZ^{+}09, SSC^{+}08]. Approaches [HH04, Gol06]. April [LGTM02]. arbitrary [LCM^{+}02]. arbitration [PPN09]. architect [MGKO08]. architectural [HKH05]. architecture [BLS06, Bas03, BH08, BMM^{+}02, BU07, DHK00, DGJ02, DNPQ01, Dos03, Fis07, GB00, GR03, HPvV^{+}08, IAM02, JCL07, KUV03, LSF05, LBF^{+}04, Maz06, MMD00, PCR^{+}08, PRT05, Rao06, SSSK00, SJM^{+}04, TB04, UVMES^{+}04, UVMES06, VFRO08, VVL^{+}09, XLVN01, BUV03]. Architectures [BWT^{+}04, Bel04, DEK^{+}05, DHM^{+}03, DER^{+}03, FLM^{+}06, HK04, KRA06a, KSA08, NO08, SDC08, SPNR03, VDR08]. area [BD00]. Areas [Koc04, DJR07]. array [CMC^{+}05]. art [Pyr03a]. ASON [HJvO^{+}06]. ASON/GMPLS [HJvO^{+}06]. Aspects [GMW07, SŽVdLvW08, DKR^{+}06]. assessment [BBR^{+}09, CC04]. assisted [RBW08]. associated [CGJ^{+}02]. assurance [BLP02, MWW06]. Asynchronous [Ton02]. ATM [MMD00, Sch00]. attacks [HLS09, LB07]. Augmenting [DJM^{+}07]. Authenticating [CGV09]. Authentication [MD04, ŽBLvB05, BBG^{+}02]. authoring [DDHS01]. Authorization [CKR^{+}04]. auto [LB07]. auto-updating [LB07]. Automated [BFMT00, HB09, LRR^{+}04]. automatically [XLVN01]. Automating [HS00, LWM08]. automation [MD02, RNMP08]. Autonomous [Cla09]. availability [ABSZ06, BF06, JK02, MA06, McP06, MC06, PMCR06, UVMES06]. avatar [CKT08]. aware [GHW08, NDC^{+}09].

backbone [CP04, HJM^{+}04]. backhaul [BYA09, LS07]. backpressure [MvHdMvE03]. backup [LRR^{+}04]. band [SSSK00]. band/multi [FEK02]. bandwidth [BDM^{+}06, VSL^{+}08, NDM^{+}06]. banking [LLZH09, Ram07]. base [BBK^{+}07, BBM^{+}09, Fis07, KKL^{+}02]. based [AMT06, BCC^{+}07, BCG^{+}09, BLS06, BR03, BBC^{+}09, Cho04a, CMPW02, CEL^{+}04, DPB09, DSA^{+}09, DJR07, FK04, GKP^{+}02, HHS08, HKUW03, HLS^{+}06, KTW00, KAGS09, KUV03, LSF05, LBF^{+}04, LDFK05, Ma07, Maz06, MC06, MKL^{+}05, NQR^{+}01, PMCR06, PS07, RFBJ04, RNMP08, SAN03, TSC^{+}03, VVL^{+}09, Wan07, XLVN01, vdGHM04, NDC^{+}09]. battery [LSF05]. beam [PWW^{+}05]. bearer [AMT06, ABK^{+}00]. become [KS06]. becomes [Net00]. behavior [BCD^{+}08, DRC^{+}05, HSP03]. behind [BKLW00]. Bell [Ano00a, Ano00b, Ano00c, Ano00d, Ano01a, Ano01b, Ano02a, Ano02b, Ano03a, BRMS00, Das00, Geh03, GM07, Hol00, MBC^{+}07, OJ04]. benefiting [Bur08]. best [BCV^{+}07, HL03]. between [AMG05, RTM^{+}04, SRR08]. Beyond [ARS^{+}06, HAH^{+}05, CMCP06, Net00, SBSS01]. Biased [DFR06]. Bio [Aiz05]. Bio-inspired [Aiz05]. bit [LGC06, SLJ^{+}09]. bit/s/Hz [SLJ^{+}09]. BLAST [LCM^{+}02]. Blended [HV06, CJKS06, PUV07]. Blending [BDE^{+}07]. block [LLCM02]. book [PUV07]. boosted [VDR08]. Brainpower [BBTD05]. breach [MCP^{+}03]. bridge [SSSK00, AMT03]. briefs [Ano00a, Ano00b, Ano00c, Ano00d, Ano01a, Ano01b, Ano02a, Ano02b, Ano03a]. brink [MCP^{+}03]. broadband [BKMV09, De 09, HKH05, PZBT03, RWGS09, TBVV08]. Broadcast [BVYY07, GSCF09, AdLvWH^{+}07, SLKAM09]. Broadcast-multicast [BVYY07]. broker [DEK^{+}05]. brokering [KRA06a]. Broker^{T M} [BDE^{+}07, KRA06b]. Building [GCSB07, KM03, SSSK00, BDE^{+}07, SBC^{+}01]. burst [LGTM02, SLK^{+}09]. business [Bar06, GA00, JMS04, PWHH00, RGAK01].

C [Tri02, DLM^{+}07]. CA [KAGS09]. cache [BGW09]. Caching [KSA08, Rad04]. Calculation [JK02]. Call [Bru05, CLW^{+}06, CLC07, CGR02, FDC^{+}07, GA00, KSO03, MD04, VW08]. Call-routing [Bru05]. calls [SSSK00]. campus [RLP05]. campus-wide [RLP05]. capabilities [Mal04]. capability [ABC^{+}05, KRA06b]. Capacitive [EPL^{+}05]. capacity [AFRZ07, CMMZ08, DJM^{+}07, FTV06, KMSW02, NNT05, RBW08, UVMES06, DC09]. capital [LUV08]. Capitalizing [Koc03]. capture [BHH^{+}00]. Carbon [Ram05, HSK^{+}09]. care [GA00]. carrier [HHH^{+}06, LE09]. carrier-grade [HHH^{+}06]. cascade [MGS^{+}05]. Case [Ram07, GEOT04, MA06]. Catastrophe [Fre06]. CCMS [IB03]. CDMA [CMMZ08, CR04, CDL07, CEL^{+}04, DHJ^{+}07, GKP^{+}02, Lau00, Ma07, MWZ07]. CDMA-based [GKP^{+}02]. CDMA2000 [BFL^{+}02, BCL^{+}02, CP07, DLM^{+}07, DLM^{+}02, DHJ^{+}07, MKL^{+}05, SGRP04].

CDMA2000-1x [DHJ^{+}07]. cell [AMG05, RCB^{+}09]. cellular [BBD^{+}05, FHM^{+}05, NT09, SLKAM09, Fre06]. CELNET [BRG^{+}05]. center [Bes07, Ma07]. centers [NNT05]. central [LMR09]. CEO [Rus04]. certificates [KAGS09]. certificates-based [KAGS09]. Chairman [Pow04, Rus04]. challenge [MGKO08]. Challenges [ASK04, BBE^{+}08, GJ04a, JJM^{+}05, NO08, AH08, ELM00, HSS^{+}09, Jos08]. change [AMS00, IB03, TA04]. changes [MW00]. Changing [CCH^{+}01]. channel [ABK^{+}00, CDL07, Lau00]. channels [LLCM02]. characterization [BF06, Sun05]. charging [CLW^{+}06, LC08, Maz06]. China [WXT^{+}08]. chip [HAH^{+}05]. circuit [AEM^{+}06, CEL^{+}04, ELOW02, JLD^{+}06]. circuits [DRC^{+}05, WHK^{+}09]. city [TGC^{+}08]. civilian [VWC^{+}04]. class [DFR06, PSST07]. clean [BBK^{+}09]. client [IBPF08]. close [FM08]. cluster [Ma07]. co [Bes06]. co-locating [Bes06]. coding [Lau00, SGK09]. collaborative [AdLvWH^{+}07, BCG^{+}09]. collection [TA04]. Com [Lab08]. commerce [CKR^{+}04]. commercialization [PWHH00]. Commission [Pow04]. Common [CP07, CE06, LC08]. communicate [CCH^{+}01]. communicating [Lab08]. communication [Aro06, BCC^{+}07, BBLS08, BCSW09, CCC^{+}02, HB09, Koc05, KP09, NQR^{+}01, CV08]. Communications [Pow04, AH08, BKLW00, DKK00, DSA^{+}09, HLS^{+}06, LGTM02, LRR^{+}04, MHM06, Rau04, RKR06, SPNR03, VK05, ZTH^{+}00]. community [TGC^{+}08]. comparative [TM05]. Comparing [HMM07]. Comparison [ESHKW06, EFH04]. Compensation [BXK^{+}09]. Compensation/mitigation [BXK^{+}09]. competency [TM05]. competitive [BBLS08, BBTD05]. competitiveness [MDMF03]. complex [Rad07]. complexity [DHK00, HSP03, Wan00]. Component [Bas03, BR03, MDMF03]. compound [GHW08]. comprehensive [RKR06]. Computer [Pyr03a]. concept [CHS08]. concepts [HP08, HMM07, Mal04]. concerns [RP08]. concurrent [IB03]. conduction [Sha07]. configuration [IB03, LWM08]. congestion [EFH04]. connecting [DJR07]. connection [ABC^{+}05, JOR00]. connection-oriented [ABC^{+}05]. considerations [AHL^{+}02, ABSZ06, BUV03, GJ04b, MESG06]. considering [Wan06]. consistency [LHMS05]. constrained [VSL^{+}08]. constructing [AG04]. consumer [BBLS08]. Contemplating [GJ04a]. Content [VDR08, BBE^{+}08, BGW09, RBW08, RNMP08, SSC^{+}08, ZMS08, vdGB08]. context [BCD^{+}08, BKMB06, MRL07, TF07]. context-free [TF07]. Continuing [KP09]. continuity [JMS04, MGKO08]. continuous [DRNY04]. control [ARS^{+}06, AAG^{+}09, AMT06, ABF^{+}07, CLW^{+}06, CGLU03, CLC07, CGR02, FM08, HSF^{+}05, HHS08, HKUW03, KMSW02, Kro04, LMMV06, MD04, NQR^{+}01, OVN08, SBSS01, SLM02, TB04, WPK^{+}07, XLVN01, CEJ^{+}03]. Controllable [SvWdBvdG02]. controlled [Aiz05]. controller [LC08]. controls [Cho04b]. Converged [JJM^{+}05, LC08, MGKO08, ABBST08, AKN^{+}07, AH08, ABBE08, BMOP00, BLP02, BDSS07, Bel04, CLC07, DCHY00, DHK00, HLS^{+}06, SM05, SJM^{+}04]. convergence [BCRL05, ELOW02, GHVH06]. cool [Sha07]. cooperation [BGW09]. coordination [HSP03]. COPS [IAM02]. CORBA [BR03]. core [AdLvWH^{+}07, CPH09, CEL^{+}04, DLM^{+}02, DDHS01, ELOW02, FHW01, GMW07, HHVST07, IGNC^{+}05]. correctness [LLZH09]. corrosion [FS06]. Cost [CDL07, AMS00, DJR07, FTV06, GM06, MC06, PS07, Tsa00, Wan06]. costs [DMO00]. count [AMG05]. counter [HLS09]. coverage [CPH09, MZP^{+}09, SFK^{+}07]. CPE [YJY06]. Craftsmanship [Pyr03a]. creation [KMS^{+}02, LPST01]. creativity [Koc04]. criteria [Yan02]. Critical [CLOU06, AJK06, HKO^{+}04, OC08]. Cross [TGC^{+}08]. Cross-media [TGC^{+}08]. crown [Geh03]. crystallization [Aiz05]. crystals [DRC^{+}05]. CSCF [LUV08]. current [AK03]. customer [BLH^{+}05]. CyberCarrier [BCGM00]. cycle [DK04]. cycling [JLD^{+}06].

dash [DSA^{+}09]. Data [MB03, ABC^{+}05, Bel04, BFL^{+}02, BFL04, Bes07, BCL^{+}02, BHLN02, BBC^{+}02, BCV^{+}07, Bru05, DLM^{+}07, DLM^{+}02, DCHY00, DHK00, FDC^{+}07, GKP^{+}02, HB09, JLD^{+}06, KTW00, KFL00, LHMS05, LCBM09, LMR^{+}06, MA06, MKL^{+}05, RR01, SSM^{+}00, Vla03, Wis04]. data-based [KTW00]. Data-driven [MB03]. DDos [HHS07]. decades [BRMS00]. decision [CS04, LGH07, Yan02]. decision-making [Yan02]. decoding [Ton02]. Dedication [Ano04a]. Defining [Har05]. DEFINITY [GA00]. Delivering [GMS^{+}00]. delivery [ABBST08, BMOP00, VSL^{+}08, Hao08, LS01, SFK^{+}07, SSC^{+}08, ZMS08]. demodulation [AdLvWH^{+}07]. denial [HLS09]. deployable [AGK^{+}03]. Deploying [Bur08]. Deployment [HC09, BDM^{+}06, Cla09, FCG^{+}06, LUV08, NNT05, OC08, Van05]. deployments [BH08, DC09]. Design [AdLvWH^{+}07, BCC^{+}07, BCG^{+}09, CCE^{+}06, LLZH09, MRD^{+}05, SW09, ABN^{+}03, BBK^{+}09, BDM^{+}06, DKLS01, FHM^{+}05, Lar08, LMR09, MC06, Moo04, RKKM05, UVMES^{+}04, LDFK05, WDC^{+}05]. Designing [Aro06, Sol07, ASK04, AGK^{+}03, Rub09, Ale07]. Detecting [LCBM09]. detection [CKR05, LCM^{+}02, MCP^{+}03, Sha04, TF07, Ton02]. Deterministic [Ma07]. Deterministic-based [Ma07]. development [AdLvWH^{+}07, BMOP00, BEM^{+}05, BDM^{+}06, BFMT00, DK04, Hol00, MD02, PWHH00, VW08, Wan00]. device [GT07, RKKM05, WHK^{+}09]. devices [GFB05, LCBM09, RKKM05, SBC^{+}01, TB08, WDC^{+}05]. Devising [SSSK00]. DHBT [WHK^{+}09]. DHT [HHS08]. DHT-based [HHS08]. Diabelli [CEJ06]. diagnostics [SDC08]. diameter [Wan07]. dictionaries [KTW00]. different [Maz06]. digital [AdLvWH^{+}07, MPGS08, VDR09, WAB^{+}08, ŽKN^{+}08]. Dimensioning [LE09]. directions [MSSS00, NO08]. Disaster [DRNY04, AMBK^{+}07, BWT^{+}04, GEOT04, HMR04, Kro04]. disasters [CCRV04]. disparate [RFBJ04]. displayed [CGV09]. dissemination [BBE^{+}08]. Distributed [AAG^{+}09, FYJB09, Moo04, ADH^{+}01, IB03, LHMS05, MB03, Nak03, XLVN01]. distribution [BGW09, ESHKW06, FM08, PBDS09, Rao06, RBW08, She07, vdGB08]. division [FYJB09]. DNS [Bes07]. do [Bi05, BVYY07, CDM^{+}07, BBC^{+}02, BCG^{+}03, MWZ07, LS07, BCV^{+}07, CP07, DLM^{+}07]. Documenting [HK04]. Domain [WSS^{+}06, GM07, HRB^{+}02, PDDS09, Pyr03b, SW09]. domains [TBVV08]. dots [DRC^{+}05]. downlink [BHT09, DGH^{+}02]. downtime [KZFB06]. driven [BDSS07, DNPQ01, MB03, SMZ^{+}09]. Driving [ARS^{+}06]. drop [VSL^{+}08]. DSL [VSL^{+}08]. Dual [TAS^{+}07]. due [CMC^{+}05]. durations [BF06]. during [Kro04]. DVB [Rub09, SLKAM09, SLK^{+}09]. DVB-SH [Rub09, SLKAM09, SLK^{+}09]. DWDM [FTV06, SLJ^{+}09]. Dynamic [BBD^{+}05, KMSW02, MVJ^{+}01, AAG^{+}09, BBG^{+}02, CEJ^{+}03, Lub00, NT09, SŽVdLvW08]. dynamics [BBLS08, HSF^{+}05].

e-business [RGAK01]. early [GM07]. Economic [FFG^{+}07a, FFG^{+}07b, PDDS09, PSST07, vVWBH05]. Economics [BIGKR07, WZ03]. ecosystem [FM08, SBLD08]. edge [MKL^{+}05]. effect [DM03]. Effective [KKKC08, Kro04, CDL07, Har05]. effects [AMS00, KMSW02]. efficiencies [DLM^{+}07]. Efficiency [vdGB08, SLJ^{+}09]. Efficient [NT09, SLKAM09, AMG05, HHVST07, NNT05, SGK09]. effort [AdLvWH^{+}07, BCV^{+}07]. Effortless [JOR00]. efforts [BFL04]. eGEM [HJvO^{+}06]. Eight [RKR06]. Electrically [KTKH05]. electronic [RKKM05]. element [ELM00, MDMF03, EJV00]. elements [LHMS05]. emergency [AMBK^{+}07, HKH05]. emergent [HSP03]. Emerging [ZJP03, AH08, ABF^{+}07, BCSW09, NO08, VDR09, VVL^{+}09]. EMS [ME00]. emulation [VW08]. emulator [HJvO^{+}06]. enable [CRVX02]. enabled [BBD^{+}00, BLV00, CJKS06, DDHS01, SBLD08, SJM^{+}04]. Enabling [Bar06, BKMB06, FZ09, LMR07, SRR08, SJ01, FEK02, HKQ^{+}04, HSK^{+}09]. End [ABSZ06, SGRP04, SDC08, Bis05a, CRVX02, JCL07]. End-to-end [ABSZ06, SGRP04, SDC08, JCL07]. endeavors [Das00]. endpoint [ACEM04]. endpoints [BH05, BHH05]. engine [ABBST08, BH08]. engineering [BCG^{+}03, CEJ^{+}03, HL03, MR01, Pyr03a, Yan02]. engineers [Rub09]. enhance [MBC^{+}07]. Enhanced [BLV00, PUV07, ZS08, BVYY07, CKR^{+}04, DLM^{+}07, HSK^{+}09, KFL00, SLK^{+}09, LKL00]. Enhancement [SR09, AHKK06, CMC^{+}05, HMS^{+}07]. Enhancing [RHSG03, RKR06]. ensuring [MVX^{+}04]. Enterprise [Mil00, Ram07]. environment [CCC^{+}02, LPST01, Ram07, RPB^{+}04]. environments [ABBE08]. epitaxy [PWW^{+}05]. equipment [GEOT04, LMR09]. error [SLK^{+}09]. eSAE [LPST01]. ESD [FS06]. essential [DMO00]. essentials [vVWBH05]. estimation [GHW08, KT07]. Ethernet [BIGKR07, Gol06, HHVST07, MZP^{+}09, MvHdMvE03]. E^{T M} [ME00]. European [BBK^{+}09, VFRO08]. EV [Bi05, BVYY07, CDM^{+}07, BCV^{+}07, CP07, DLM^{+}07, LS07, MWZ07]. EV-do [Bi05, BVYY07, CDM^{+}07, BCV^{+}07, CP07, DLM^{+}07, MWZ07, LS07]. Evaluation [HHS08, AK03, Ahm02, Kol09]. event [LGTM02, Nak03]. every [RCB^{+}09]. everywhere [AKW00]. Evolution [BGK^{+}06, BD00, DGH^{+}02, GKP^{+}02, MMD00, AKW00, BGDT07, CCD^{+}00, CV08, DHJ^{+}07, HSF^{+}05, HJM^{+}04, JJM^{+}05, LS01, RS09, SJ01, ZTH^{+}00, DLM^{+}07]. evolutionary [RWGS09]. evolutions [VDR09]. Evolving [Aus06, DMO00]. Evros [SFK^{+}07]. examples [FFG^{+}07b]. excel [JLN03]. exchange [BBG^{+}02, CRVX02]. excursion [DRC^{+}05]. existing [HC09, TAS^{+}07]. expenses [LUV08]. experience [ADH^{+}09, Van05]. experiences [TGC^{+}08]. experiments [WHK^{+}09]. Experium^{T M} [BBTD05]. exploiting [BKMB06]. exploration [AMG05]. extended [DK04]. Extending [BH05, SFK^{+}07]. extensible [AG04]. Extreme [Pyr03b].

fabricated [PWW^{+}05]. fabrication [WDC^{+}05]. factor [Sha07]. fading [LLCM02]. failure [MCP^{+}03, Rao06]. fairness [SR09]. Fast [ADH^{+}01, Lub00, Nak03, PPN09]. FASTeR [ANPP04]. fault [CKR05, IGNC^{+}05, MB03, NQR^{+}01]. fault-tolerant [NQR^{+}01]. feature [DDHS01, HS00]. features [CCD^{+}00]. Federal [Pow04]. feedbacks [LLCM02]. femtocell [CHS08, CPH09, DC09]. femtocells [BKMV09, HC09]. Fi [GMW07]. fiber [GDS^{+}00, LGC06, WZ03, vVWBH05]. fiber-to-the-premises [vVWBH05]. field [BHH^{+}00, Bi05, McP06]. file [She07]. finance [Der04]. fingerprinting [GT07]. firewalls [SvWdBvdG02]. Flexent(R) [DLM^{+}02, ELOW02]. Flexent^{T M} [EJV00]. flexible [DKLS01]. Flow [NDC^{+}09]. Flow-based [NDC^{+}09]. flowing [DFR06]. focus [PCR^{+}08]. footprint [HSK^{+}09]. force [LDFK05]. forecasting [AJK06]. form [Sha07]. form-factor [Sha07]. Formal [FHOS04]. Forward [BCL^{+}02, Bi05]. foundations [ARS^{+}06]. FPGA [DK04]. Framework [BF06, AG04, ABBE08, BGK^{+}06, BR03, BCSW09, CC04, GB00, GM07, HMH04, LGH07, MVX^{+}04, MBC^{+}07, MB03, MK01, RKR06, Wan01]. Free [WDC^{+}05, Fre06, TF07]. Free-space [WDC^{+}05]. frequency [Fis07, PBDS09]. Friends [CJKS06, MBB00]. function [ABF^{+}07, Kol09, LGH07, LC08]. Functionality [DEK^{+}05, HSK^{+}09]. Future [DNPQ01, VK05, BBK^{+}09, BBD^{+}05, DKK00, HKH05, Koc02, KS06].

GaAs [PWW^{+}05]. gains [RBW08]. gaming [NO08]. GaN [FEK02]. gas [DFR06]. gateway [HHH05, HHH^{+}06]. Gb [SLJ^{+}09]. Gb/s [SLJ^{+}09]. gene [HAH^{+}05]. general [ADH^{+}01]. Generalized [XLVN01]. generating [SJ01]. generation [AP05, Ale07, AMT03, BBDK04, BBR^{+}09, BM03, BBC^{+}02, BU07, Bur08, CE06, CGLU03, CMPW02, De 09, DSW08, Fis07, FHM^{+}05, Hao08, HPvV^{+}08, KMSW02, KKL^{+}02, KMS^{+}02, Koc05, Koc07, KP09, KGT06, LPW02, LKL00, LPST01, MMBDK08, NO08, PCR^{+}08, RP08, SBC^{+}01, SLW^{+}09, SJ01, SJM^{+}04, WvLS08, SLM02]. generator [Wan01]. Generic [LGH07, Liu08]. geographically [IB03]. Gigabit [MZP^{+}09]. Global [IAM02, Cow07, FFG^{+}07a, FFG^{+}07b, PWHH00, RWGS09, Van05]. GMPLS [HJvO^{+}06]. GoS [DC09]. grade [HHH^{+}06, LE09]. grammar [TF07]. grid [BCV09]. ground [Aus06]. group [Rao06]. growing [Koc03]. growth [AJK06]. guide [Rub09].

handling [Ma07]. handoff [Ahm02].

handover [DKR^{+}06]. handset [DJM^{+}07]. happening [Aus06]. hardening [Bur08, Moo04]. hardware [AJK06, Moo04]. harmonized [Maz06]. haul [FTV06]. Health [GA00]. healthier [CKR05]. heterogeneous [RPB^{+}04, ŽBLvB05]. hierarchical [PPN09, Rao06]. High [DSA^{+}09, BBC^{+}02, BDL09, BXK^{+}09, DGH^{+}02, FTV06, GGHH^{+}03, GKP^{+}02, LGC06, MGS^{+}05, Vla03]. high-performance [GGHH^{+}03]. high-power [MGS^{+}05]. high-speed [BBC^{+}02, BDL09, BXK^{+}09, DGH^{+}02, GKP^{+}02, Vla03]. higher [DLM^{+}07]. historical [KTW00]. Hoadley [JLN03]. home [TBVV08, VVL^{+}09, WZ03]. homeland [BBDK04]. homomorphism [TF07]. hop [AAG^{+}09, SLW^{+}09]. hosted [CKT08]. hosting [CCC^{+}02]. HSDPA [BJK^{+}07, GSCF09]. HSPA [RWGS09]. HSUPA [BJK^{+}07]. human [Mac04]. Hunsaker [Tri02]. Hybrid [EACM00, RLP05, SLKAM09, ZS08]. Hz [SLJ^{+}09].

identification [Mac04, TF07]. idle [CLdJH08]. IEC [BFL04]. IEEE [BCC^{+}07, BCG^{+}09, DKR^{+}06, DJR07]. IETF [BFL04]. IFEC [SLK^{+}09]. II [FFG^{+}07b]. imaging [JBR^{+}09]. immunity [Moo04]. impact [HC09, Kol09]. impacts [UVMES06]. imperative [PWHH00]. imperfect [Wan06]. implementation [PS07, RLP05, Wan01]. implementations [CKR^{+}04]. Implementing [CR04, GB00, GGHH^{+}03, HH04]. implications [DC09]. improved [GM06, Mal04]. Improving [Bes06, Bes07]. IMS [AMT06, AFG^{+}06, ABSZ06, AEM^{+}06, BGK^{+}06, Bes06, Bes08, BKMB06, CJKS06, CFL^{+}06, CEJ06, CEL^{+}04, ELOW02, GL06, GMW07, HV06, HMS^{+}07, IBPF08, JCL07, Jos08, KRA06b, LUV08, LMR^{+}06, Liu08, Maz06, PMCR06, PCR^{+}08, PSST07, SBLD08, UVMES06, WRL^{+}09, WSS^{+}06, WC08]. IMS-based [Maz06, PMCR06]. IMS-enabled [SBLD08]. increase [CPH09]. Increased [MDMF03]. independent [BDSS07, DFS^{+}05]. indexing [RNMP08]. industry [BGTU05]. inference [HB09]. Inferring [MRL07]. Information [MBC^{+}07, BKMB06, Cho04b]. infrared [JBR^{+}09, MGS^{+}05]. infrastructure [AKN^{+}07, BCC^{+}04, CLOU06, FM08, KRLV08, LRR^{+}04, Mac04, Rau04, RKR06, Sha04, SBEW09, VDR09]. infrastructures [HKO^{+}04, OC08]. ingredients [RKR06]. initiation [DGJ02, GL03, GJ04b, Zen04]. innovation [Ano00a, Ano00b, Ano00c, Ano00d, Ano01a, Ano01b, Ano02a, Ano02b, Ano03a, Koc04, TA04]. innovations [BRMS00, HPvV^{+}08]. innovative [SBLD08]. InP [DSA^{+}09, WHK^{+}09]. InP-based [DSA^{+}09]. IN(R) [CCC^{+}02]. insight [Pan06]. inspired [Aiz05]. installation [Moo04]. integrated [CC04, HRB^{+}02, Pan06, SLM02]. Integrating [MC06, SM05, ASK04, TAS^{+}07]. integration [GA00, MHH05, WDC^{+}05]. integrity [MD04]. Intelligence [FM08, Cho04b, MK01]. Intelligence-ready [FM08]. Intelligent [BJM^{+}02, DHM^{+}03, HHH05, HHH^{+}06, Jai02, Pan06, ZMS08, BH05, VSL^{+}08, GMS^{+}00, STB07, Sun06]. Intelligent^{T M} [DHM^{+}03, Dos03, GR03, SPNR03]. Inter [BCC^{+}04, PDDS09]. inter-domain [PDDS09]. Inter-infrastructure [BCC^{+}04]. interaction [KRA06b]. interactions [SRR08, TGC^{+}08]. interactive [NO08, RNMP08, ZMS08]. intercept [FS06]. interface [Ale07, AMT03, BGDT07]. interleaved [SW09]. international [AdLvWH^{+}07]. Internet [LS01, BBK^{+}09, GM07, PRS01, SBC^{+}01, SBSS01, UGDS01]. interoperability [CFL^{+}06, JOR00]. intranet [Bes07]. Introduction [Kor06, AHKK06]. Intrusion [Sha04]. intrusive [KT07]. Intuitive [BCD^{+}08]. invocation [Liu08]. iOptimize [ABC^{+}05]. IP [AHL^{+}02, AK03, BLS06, Bar06, BDSS07, BCV^{+}07, CLW^{+}06, CGLU03, CLC07, CMPW02, CLdJH08, DM03, Koc03, MWW06, PS07, SBSS01, SJM^{+}04, TAS^{+}07, VDR09]. IP-based [BLS06, CMPW02, PS07]. IP-driven [BDSS07]. IPsec [KAGS09]. IPTV [BDE^{+}07, DDL08, ESHKW06, EM08, HHVST07, KSA08, SDC08, VDR08]. IPv4 [CFL^{+}06]. IPv4/IPv6 [CFL^{+}06]. IPv6 [CFL^{+}06]. IS-95 [Ahm02]. ISO [BFL04, MBC^{+}07]. issue [PT07]. Issues [ACEM04]. ISUP [EFH04]. Iterative [LCM^{+}02]. ITU [BFL04, Ram07]. ITU-T [Ram07].

Java [BFMT00, GGHH^{+}03]. Jerome [Tri02]. jewel [Geh03]. Jini [JOR00]. just [CMMZ08]. just-in-time [CMMZ08].

key [BBG^{+}02]. km [SLJ^{+}09]. Knowledge [BBTD05].

Labs [GM07, OJ04, Ano00a, Ano00b,

Ano00c, Ano00d, Ano01a, Ano01b, Ano02a, Ano02b, Ano03a, BRMS00, Das00, Geh03, Hol00, MBC^{+}07, OJ04]. LambdaRouter [NQR^{+}01]. LambdaUnite(R) [BDM^{+}06]. LambdaXtreme(R) [FTV06, FCG^{+}06]. LAN [HVKL03]. landscape [PWHH00]. language [BFMT00, KSO03]. LANs [Sha04]. large [TB04]. large-scale [TB04]. laser [MGS^{+}05]. lasers [BKLW00, DSA^{+}09]. last [VSL^{+}08]. Law [Bis05a]. layer [AHKK06, Aus06, FHM^{+}05, HMS^{+}07, HKQ^{+}04, STB07]. layers [DM03].

Learning [BCD^{+}08, TAS^{+}07]. legacy [AEM^{+}06]. Letter [OJ04, Pow04, Rus04]. level [BCG^{+}03, Har05, LWM08, MvHdMvE03, SBSS01, SŽVdLvW08]. levels [RTM^{+}04]. life [AJK06, Der04, DK04, Geh03, Van05]. life-cycle [DK04]. lifestyle [CJKS06, PUV07]. lightpaths [DKLS01]. line [Hol00, MPGS08, WAB^{+}08, ŽKN^{+}08]. link [BDE^{+}07, BCL^{+}02, Bi05, CEJ^{+}03, DM03, GKP^{+}02, Lau00]. live [TAS^{+}07]. locating [Bes06]. Location [SAN03, HLS^{+}06]. Location-based [SAN03, HLS^{+}06]. locked [DSA^{+}09]. long [BGDT07, FTV06]. loss [DDL08]. low [DJR07, FTV06]. low-cost [DJR07]. LTE [RWGS09]. Lucent [BDE^{+}07, Ano06b, Ano06c, Ano06d, Ano01c, Ano02c, Ano02d, Ano02e, Ano03b, Ano03c, Ano03d, Ano03e, Ano04b, Ano04c, Ano04d, Ano04e, Ano05a, Ano05b, Ano05c, Ano05d, Ano06a, Ano07a, Ano07b, Ano07c, Ano07d, Ano08a, Ano08b, Ano08c, Ano08d, Ano09a, Ano09b, Ano09c, Ano09d, AHKK06, KRA06b, NQR^{+}01, Rus04, RCB^{+}09].

macrocellular [HC09]. Macroscopic [Sun05]. made [PUV07]. MAGNET [CMCP06]. Maintaining [VSL^{+}08, LHMS05]. maintenance [BLH^{+}05, Bel04, DHM^{+}03, Har05, Wan06]. making [Yan02]. malware [TF07]. Man [LB07]. Man-in-the-middle [LB07]. manage [EJV00, HMR04]. Managed [HVKL03, NDC^{+}09]. Management [EJV00, MBC^{+}07, ABN^{+}03, Bas03, BDSS07, BDM^{+}06, BJM^{+}02, BCGM00, BJK^{+}07, BCRL05, BRG^{+}05, Cho04a, CP07, DKR^{+}06, Ded03, DHM^{+}03, DPB09, EACM00, ELM00, GB00, HL06, HSK^{+}09, HRB^{+}02, IGNC^{+}05, IB03, KRA06b, KAGS09, LMR^{+}06, LKL00, MDMF03, MB03, MSSS00, NDM^{+}06, QW05, Ram04, RHSG03, RPB^{+}04, Sch00, SBS00, SŽVdLvW08, Van05, WSS^{+}06]. manager [BUV03, LKL00]. Manager^{T M} [ME00]. Managing [ECGL01, KTW00, MvHdMvE03, TBVV08, TA04, SM05, TB04]. manufacturing [WDC^{+}05]. MapWeb [HLS^{+}06]. market [BBLS08]. markets [BBLS08]. mash [SMZ^{+}09]. mash-up [SMZ^{+}09]. MASS [Flo06]. massive [DC09]. materials [RKKM05]. Mathematical [FFG^{+}07a]. Maximizing [MPGS08]. measurement [FDC^{+}07, HKUW03]. measurement-based [HKUW03]. measurements [Bi05]. Measures [BHLN02]. Measuring [AMS00, SAN03]. mechanism [Bes07, SLKAM09]. mechanisms [CKR^{+}04, MD04]. Media [SSC^{+}08, ZZLD04, HHH05, HHH^{+}06, SvWdBvdG02, TGC^{+}08]. mediation [BH05, BGUV03, DDHS01, GMS^{+}00, KUV03]. meet [BCV09]. meets [ADH^{+}09]. memory [PPN09]. MEMS [EPL^{+}05, FEK02, GFB05, JBR^{+}09, LDFK05]. MEMS-based [LDFK05]. merchant [MGKO08]. mesh [ANPP04, ADH^{+}01, DJR07, Nak03]. message [BMvB06, CDL07, HHS08, MD04]. messaging [HV06]. methodology [DFS^{+}05]. Metrics [TM05, HSP03]. metro [GR03, MvHdMvE03]. microphones [EPL^{+}05]. microprojector [RCB^{+}09]. mid [MGS^{+}05]. mid-infrared [MGS^{+}05]. middle [LB07]. Middleboxes [SvWdBvdG02]. migrate [SSSK00]. migration [FZ04, PSST07, WC08]. mile [VSL^{+}08]. millennium [Mil00]. MIMO [AP05, BHT09]. minimize [CPH09, DK04]. MiRingBack [BGK^{+}06]. mitigation [BXK^{+}09]. mixed [ACEM04, DFR06]. Mobile [AHS01, BGTU05, Flo06, Tor02, AK03, BBR^{+}09, BFL^{+}02, BVYY07, CKR^{+}04, LLZH09, Ma07, MA06, MRL07, RWGS09, SRR08, SBLD08, Tor00, UGDS01, ZTH^{+}00]. mobiles [CEL^{+}04]. Mobility [RPB^{+}04, AEM^{+}06, CPH09, IAM02, QW05]. mode [CLdJH08, DSA^{+}09, TAS^{+}07]. mode-locked [DSA^{+}09]. model [BR03, GM06, HKO^{+}04, JMS04, Keb00, VW08, Yan02]. model-based [BR03]. Modeling [BBLS08, KZFB06, BCC^{+}04, CP04, CLOU06, DSW08, FFG^{+}07a, FFG^{+}07b, FHOS04, JLD^{+}06, LWM08, Ma07]. models [Bar06, FFG^{+}07a, HHVST07, KM03, UVMES^{+}04, Wan06, KKKC08]. modern [TBVV08]. Modular [BBG^{+}04]. module [NS03]. molecular [PWW^{+}05]. Monitoring [AKN^{+}07, HP08, TSC^{+}03]. Moore [Bis05a]. Moving [Koc05]. MPE [SLK^{+}09]. MPE-IFEC [SLK^{+}09]. MPLS [IGNC^{+}05, NQR^{+}01, TSC^{+}03, XLVN01]. MPLS-based [NQR^{+}01, XLVN01]. ms [ANPP04]. MSPPS [KGT06]. Multi [GR03, Nak03, SLW^{+}09, SBEW09, ASK04, AAG^{+}09, BLS06, BEM^{+}05, BHT09, BKMV09, FEK02, HHS07, HRB^{+}02, LWM08, NT09, PPN09, Sun06, VFRO08, Yan02]. multi-agent [Sun06]. multi-band [FEK02]. multi-band/multi-standard [FEK02]. multi-criteria [Yan02]. Multi-event [Nak03]. Multi-hop [SLW^{+}09, AAG^{+}09]. multi-level [LWM08]. multi-operations [ASK04]. multi-operator [BKMV09]. multi-perspective [HHS07]. Multi-radio [SBEW09]. Multi-service [GR03, BLS06, BKMV09, VFRO08]. multi-standard [FEK02]. multi-step [NT09]. multi-Tb [PPN09]. multi-Tb/s [PPN09]. multi-technology [HRB^{+}02]. multi-user [BHT09]. multi-vendor [BEM^{+}05]. multicarrier [Ton02]. multicast [BVYY07, GSCF09, SGK09]. multicomponent [Lub00]. Multimedia [RNMP08, AHL^{+}02, ACEM04, Bar06, BGK^{+}06, CLW^{+}06, CGLU03, CLC07, De 09, SSSK00, TAS^{+}07, TB08]. multiple [BYA09, FYJB09, Ton02, TB08]. multiplexing [BCG^{+}09]. multipoint [BBC^{+}09]. multipriority [MR01]. multiservice [DDHS01, Jai02, MR01, SM05]. multiuse [GLM^{+}02]. MUSE [VFRO08]. My [Der04].

names [CGV09]. nano [WDC^{+}05].

nano-optical [WDC^{+}05]. Nanophotonics [DRC^{+}05]. nanoscale [Aiz05]. nanostructured [KTKH05, LSF05]. Nanostructures [PWW^{+}05]. Nanotechnology [Bis05a]. nanotubes [Ram05]. narrowband [KT07]. NAT [LCBM09]. national [CLOU06, FCG^{+}06, HKO^{+}04, KT07]. Native [CLdJH08]. natural [KSO03]. nature [Net00]. near [BHT09]. near-optimum [BHT09]. need [CGR02, HHS07]. needs [Koc02, RCB^{+}09]. neglected [Ded03]. NEMS [GFB05]. Network [BHH05, CEJ^{+}03, CKT08, DFS^{+}05, EFH04, HJvO^{+}06, KAGS09, MESG06, NNT05, Sie01, SGK09, VVL^{+}09, AMBK^{+}07, ABN^{+}03, AMT03, Bas03, Bes06, BCD^{+}08, BMM^{+}02, BD00, BCGM00, CMMZ08, CKR05, Cho04a, CC04, CPH09, CS04, CEL^{+}04, De 09, DHM^{+}03, DPB09, DER^{+}03, DKK00, DJR07, EJV00, ESHKW06, EACM00, FHW01, FCG^{+}06, FDC^{+}07, FM08, GMW07, GT07, HHH05, Har05, HL06, HHVST07, HPvV^{+}08, HRB^{+}02, HLS09, HKO^{+}04, HJM^{+}04, JJM^{+}05, Jos08, LE09, LWM08, LRR^{+}04, LC08, MA06, MMBDK08, MVX^{+}04, MC06, MK01, MKL^{+}05, NQR^{+}01, NO08, PS07, RLP05, RKR06, RHSG03, RPB^{+}04, Rub09, Sch00, She07, SBS00, SPNR03, SSC^{+}08, TSC^{+}03, Tor00, TB04, UVMES^{+}04, VDR09, VW08, Vla03, ZS08]. Network-hosted [CKT08]. network-provided [DKK00]. Network-supported [Sie01]. networking [Aus06, CMCP06, CV08, FCG^{+}06, Mil00, Net00, NDC^{+}09, SSM^{+}00]. Networks [BCV09, KS06, AMG05, AK03, AKN^{+}07, AAG^{+}09, AGK^{+}03, ANPP04, ABC^{+}05, ADH^{+}01, ACEM04, AEM^{+}06, BLS06, BLP02, BLV00, BDSS07, BMN^{+}00, BBE^{+}08, BBC^{+}09, Bel04, BJM^{+}02, BBG^{+}02, BBD^{+}05, BGW09, BU07, Bur08, BRG^{+}05, CMCP06, CMMZ08, CE06, CGLU03, CR04, CP04, CCE^{+}06, CEJ^{+}03, CMPW02, CP07, Cla09, CLdJH08, DKLS01, DHK00, DSW08, ELOW02, ECGL01, FK04, Fre06, FZ04, GHW08, Gol06, GR03, GMS^{+}00, GJ04b, HHH^{+}06, HL06, HP08, HH04, HKH05, HSP03, HB09, HC09, HRB^{+}02, IGNC^{+}05, Jai02, JGA04, JMS04, KRLV08, Koc02, KMS^{+}02, Koc03, Koc05, Koc07, KP09, KAGS09, KSA08, KH02, LCBM09, LMR^{+}06, LKL00, MHM06, MWW06, MvHdMvE03, MRD^{+}05, MVJ^{+}01, MR01, NT09, OVN08, PBDS09, Pan06, PMCR06, PCR^{+}08, PRT05, QW05, Rad07, RGAK01, RFBJ04, RR01, RTM^{+}04, RP08, RBW08]. networks [SW09, Sch00, SM05, SLW^{+}09, SGRP04, SBSS01, SLM02, SPNR03, STB07, SJ01, SJM^{+}04, TBVV08, TAS^{+}07, Tsa00, UGDS01, UVMES^{+}04, VWC^{+}04, VVL^{+}09, Vla03, Wan01, WPK^{+}07, WvLS08, XLVN01, ZJP03, ZZLD04, ŽBLvB05, vVWBH05, vdGHM04]. Next [DSW08, Fis07, KGT06, NO08, WvLS08, AP05, Ale07, AMT03, BBR^{+}09, BM03, BU07, Bur08, CE06, CGLU03, De 09, FHM^{+}05, Hao08, HPvV^{+}08, KMS^{+}02, Koc05, Koc07, KP09, LKL00, LPST01, Net00, PCR^{+}08, RP08, SBC^{+}01, SLW^{+}09, SJ01, SJM^{+}04, SLM02]. Next-generation [DSW08, Fis07, KGT06, NO08, WvLS08, AP05, Ale07, AMT03, BBR^{+}09, BM03, BU07, Bur08, CE06, CGLU03, De 09, FHM^{+}05, Hao08, HPvV^{+}08, KMS^{+}02, Koc05, Koc07, KP09, LKL00, LPST01, PCR^{+}08, RP08, SBC^{+}01, SLW^{+}09, SJ01, SJM^{+}04, SLM02]. night [CJKS06]. nines [McP06]. Nobel [Das00]. Noble [Das00]. node [MKL^{+}05]. nodes [Ma07]. noise [LGTM02]. nomadic [FM08]. non [KT07, SR09, VWC^{+}04]. non-civilian [VWC^{+}04]. non-intrusive [KT07]. non-orthogonal [SR09]. notification [ABBE08]. Novel [AMG05, KKL^{+}02, HLS09, KRA06a, Vla03, WRL^{+}09]. NRZ [SLJ^{+}09]. NRZ-VSB [SLJ^{+}09]. number [LCM^{+}02]. nurse [GA00]. NZDSF [SLJ^{+}09].

OAM&P [MB03]. object [EJV00, MB03]. object-oriented [MB03]. Ocelot(R) [DFS^{+}05]. Octavia [BM03]. ODM [YJY06]. OEM [YJY06]. OEM/ODM [YJY06]. OFDM [BDL09]. OFDMA [BCC^{+}07]. off [PS07]. office [LMR09]. offs [DDL08, LUV08]. One [Hol00]. OneLink [ME00]. online [CLW^{+}06]. Open [BMN^{+}00, Bes08, GGHH^{+}03, GMS^{+}00, GJ04a, UGDS01, BGTU05]. opening [KFL00]. operation [Bel04, Har05, LE09]. operations [AMBK^{+}07, ASK04, CC04, Tsa00, TM05]. operator [BKMV09]. operators [WXT^{+}08]. opportunistic [BKMV09]. Opportunities [HB09, JBR^{+}09, MHH05, JJM^{+}05]. Optical [AMT03, BDL09, GHVH06, LGC06, AKW00, AGK^{+}03, ADH^{+}01, Aus06, BXK^{+}09, BCRL05, CCE^{+}06, DKLS01, DRC^{+}05, DSA^{+}09, ECGL01, FCG^{+}06, GR03, HL06, HP08, HPvV^{+}08, HSF^{+}05, LMMV06, MC06, MESG06, NDM^{+}06, Sha07, SLM02, SPNR03, WDC^{+}05]. Optical/packet [GHVH06, BCRL05]. optics [GDS^{+}00, AKW00]. Optimal [LLCM02, OC08, PMCR06, RR01, Rao06, Ton02, Ton02]. Optimization [CPH09, BBD^{+}05, BRG^{+}05, DRNY04, DFS^{+}05, KSA08, ŽKN^{+}08]. optimized [DLM^{+}07]. optimizing [ABC^{+}05, VSL^{+}08]. optimum [BHT09]. options [HC09, LUV08]. OPUCE [SMZ^{+}09]. organizing [BGW09, HSP03]. oriented [ABC^{+}05, BU07, DPB09, KRLV08, MB03]. origin [MD04]. orthogonal [SR09]. OSS [PRT05]. other [NO08]. outage [BF06]. overhead [HHS08]. overload [EFH04]. Overview [BFL04, Bis05b, CRT04, HT08, KL06, KW03, Man02a, Man02b, MB05, SBG09, Sah00, Swe03, CHS08, Das00, PT07, WXT^{+}08].

P2P [HHS08, ZS08]. pack [JLD^{+}06]. package [BDE^{+}07]. packaging [FS06]. Packet [DHJ^{+}07, MWW06, AMT03, BCRL05, VSL^{+}08, CEL^{+}04, DGH^{+}02, DLM^{+}02, DDL08, ELOW02, GHW08, GHVH06, HSS^{+}09, HJM^{+}04, LE09, MKL^{+}05, PBDS09, PPN09, SSSK00, Tsa00, Wan01, CCC^{+}02]. PacketStar(R) [MD02]. paging [CDL07, NT09]. paradigm [ELM00, Tor00]. parallel [She07]. Parlay [BUV03, KUV03]. partial [LLCM02]. partnership [HMR04]. party [CGR02, DDHS01, FM08, SRR08, SBLD08]. patents [Ano01c, Ano02c, Ano02d, Ano02e, Ano03b, Ano03c, Ano03d, Ano03e, Ano04b, Ano04d, Ano05c, Ano05d, Ano06a, Ano06b, Ano06c, Ano06d, Ano07a, Ano07b, Ano07c, Ano07d, Ano08a, Ano08b, Ano08c, Ano08d, Ano09a, Ano09b, Ano09c, Ano09d, Ano04c, Ano04e, Ano05a, Ano05b]. path [CEL^{+}04, Lab08, MWW06, RR01]. patterns [HK04]. Peer [RBW08, SSC^{+}08]. Peer-assisted [RBW08]. peer-to-peer [SSC^{+}08]. peering [Jos08]. penetration [BU07]. Performance [BBR^{+}09, BBC^{+}02, BCV^{+}07, CGJ^{+}02, WAB^{+}08, ŽKN^{+}08, AMG05, Ahm02, Aro06, BCG^{+}09, BHH^{+}00, Bes06, BCL^{+}02, BHLN02, Bi05, BCSW09, CEE04, DSA^{+}09, DC09, GGHH^{+}03, HL03, LUV08, Ma07, MPGS08, RBW08, SAN03, SW09]. performance-measuring [SAN03]. personal [CMCP06, IAM02, ZTH^{+}00]. personalized [BCSW09, vdGB08]. perspective [HHS07, JBR^{+}09]. perspectives [LMMV06]. phone [PUV07, RCB^{+}09]. phones [CGJ^{+}02]. photonic [DRC^{+}05]. Physical [FHM^{+}05, DM03]. Physical-layer [FHM^{+}05]. physics [Der04, WHK^{+}09].

Piecewise [JLD^{+}06]. PINT [KFL00]. plane [ARS^{+}06, LMMV06]. planned [KZFB06]. planning [CMMZ08, MESG06]. plasmon [CMC^{+}05]. Plastic [RKKM05]. platform [ABK^{+}00, BBG^{+}04, BM03, DLM^{+}02, FS06, HLS^{+}06, MBB00, NS03, SFK^{+}07, WXT^{+}08]. platforms [ABBST08, DCHY00, LPW02, LS01]. play [ABF^{+}07]. pluggable [Sha07]. PMD [BXK^{+}09]. policies [HMH04]. Policy [Cho04a, FK04, HKQ^{+}04, LGH07, QW05, RHSG03]. Policy-based [Cho04a, FK04]. Polymorphic [TF07]. Portable [VW08]. Portal [CCD^{+}00]. ports [BCC^{+}04]. posture [VWC^{+}04]. power [HSF^{+}05, KMSW02, LLCM02, MGS^{+}05, OC08]. Practical [Rub09, MESG06]. practice [AAG^{+}09, HL03]. practices [KKKC08]. precoders [WAB^{+}08]. precoding [BHT09]. Predicting [MW00, MRL07]. prediction [CKR05]. premises [vVWBH05]. prepaid [CKR^{+}04]. preparedness [DRNY04, JMS04]. preparing [HMR04]. Presence [UVMES06, BKMB06, Cow07, FLM^{+}06, Koc03]. President [OJ04]. Preventative [BLH^{+}05]. preventing [GT07]. preventive [Wan06]. preview [WvLS08]. principles [RHSG03]. prioritization [BMvB06]. priority [CR04]. private [HVKL03, MVJ^{+}01, RGAK01, RTM^{+}04]. Prize [Das00]. Prize-winning [Das00]. proactive [BLH^{+}05]. probing [Wan07]. problem [HHS07]. process [DRNY04, GM06, Hao08]. processing [ABK^{+}00, HSS^{+}09]. processor [CS04, GLM^{+}02, MKL^{+}05, Vla03]. processor-based [MKL^{+}05]. product [KRA06b, LS01, MD02]. production [Wis04]. products [GCSB07, PWHH00, YJY06]. profile [CE06, Ma07]. profiling [ABBST08]. Programmable [DCHY00, ABK^{+}00, BMN^{+}00, DDHS01]. programming [Pyr03a, Pyr03b, ADH^{+}09]. project [Van05, VFRO08]. promising [BDL09]. Promoting [SBLD08]. proof [LLZH09]. proposal [HKH05, Jos08]. proposition [HMS^{+}07, IGNC^{+}05]. propositions [PDDS09]. protected [DKLS01]. Protecting [DDL08, Rau04, Mac04]. protection [FS06, SLK^{+}09]. Protocol [KFL00, Aro06, DGJ02, GCK01, GL03, GJ04b, LLZH09, Wan07, Zen04]. protocols [Nak03]. prototype [CJKS06]. provided [DKK00]. provider [Bar06, BH08, FZ04, HMR04, HJM^{+}04, Koc02, ZJP03]. providers [CRVX02, MESG06]. providing [KRLV08, vVWBH05, MZP^{+}09]. Provisioning [JCL07, Rad07, ARS^{+}06, DKLS01, Liu08, Nak03, Tor00, Van05]. proxy [BUV03]. PSAX [MD02]. pseudo [STB07]. pseudo-wires [STB07]. PSTN [KFL00, WC08]. PTT [LS07]. Public [VDR09].

Q3 [Wan00]. QMP [JLN03]. QoE [GHW08]. QoS [AK03, BCG^{+}03, DER^{+}03, DM03, JCL07, PZBT03, SGRP04, WPK^{+}07]. Quality [CDM^{+}07, CMPW02, KH02, vdGHM04, VSL^{+}08, EM08, FK04, KT07, Kro04, MC06]. quant [Der04]. Quantifying [IGNC^{+}05, JMS04]. quantitative [HMS^{+}07]. quantum [DRC^{+}05, DSA^{+}09, MGS^{+}05]. query [Ma07].

R&D [FTV06]. Radio [BJK^{+}07, Ale07, BJM^{+}02, CMMZ08, CMPW02, FEK02, Fis07, KKL^{+}02, LGTM02, PS07, SBEW09]. RADIOSTAR [MZP^{+}09]. randomized [SW09]. Rapid [BMOP00, BEM^{+}05, PWHH00, ARS^{+}06, LPST01, VW08]. rate [Lau00, LGC06]. reach [SFK^{+}07]. readings [TA04]. ready [FM08]. real [ABC^{+}05, PUV07, SLKAM09, Van05]. real-life [Van05]. real-time [SLKAM09]. reality [KS06, Tor02]. realization [EM08]. receive [LCM^{+}02]. recommendation [Ram07]. Recommendations [VFRO08]. Reconfigurable [AP05, FEK02]. recovering [CCRV04]. recovery [AMBK^{+}07, BWT^{+}04, DRNY04, HMR04, Kro04, MCP^{+}03, NNT05]. reduce [GEOT04]. reduced [HSK^{+}09]. Reducing [DHK00, Wan00, CCRV04]. reduction [GM06]. redundant [BBC^{+}09]. Refactoring [NS03]. reflections [Der04, FLM^{+}06]. regulatory [Gol05]. relationship [Sun05]. Relay [BYA09, SLW^{+}09]. Relay-aided [BYA09]. release [Ram04]. releases [KTW00]. Reliability [AJK06, CP04, Gol06, MHM06, PCR^{+}08, RP08, Bes06, CC04, CLOU06, DFR06, HMM07, MC06, RKR06, Sun05, Sun06, YJY06]. reliability-based [MC06]. reliable [HHVST07]. remanufactured [GEOT04]. Remote [HL06]. repair [SLKAM09, Wan06]. requirements [BCV09, Gol05, HMM07, JGA04, KKKC08, KM03, PRT05]. rescue [Mal04]. research [Cow07, Das00]. Reserve [LSF05]. reserves [OC08]. residency [GL03]. resiliency [GL03]. Resilient [BBC^{+}09]. resolve [McP06]. resonances [CMC^{+}05]. Resource [BBTD05, OVN08, ABF^{+}07, BDSS07, BJK^{+}07]. resources [BJM^{+}02, HLS09]. Respond [McP06]. response [AMBK^{+}07, HKH05]. restoration [ANPP04, ADH^{+}01, GEOT04, Nak03]. restore [McP06]. results [WAB^{+}08]. reuse [BKMV09]. Rev [Bi05]. revenue [SJ01]. revenue-generating [SJ01]. reverse [GKP^{+}02, Lau00]. Revision [BVYY07, CDM^{+}07, BCV^{+}07, DLM^{+}07]. revisited [WZ03]. revolution [BKLW00, CV08]. Revolutionary [RWGS09]. RF [FEK02, GFB05, Moo04, Rub09].

RF-MEMS [FEK02]. rise [Tri02]. risk [Ded03, MW00]. risks [JMS04]. roaming [IAM02]. robust [Nak03]. role [BCRL05, CCRV04, FZ04, RGAK01, STB07]. rollout [FCG^{+}06]. routed [MMBDK08]. router [BBK^{+}07, MKL^{+}05]. routing [Bru05, GCK01, KSO03, RR01]. rural [DJR07].

s [SLJ^{+}09, PPN09, SLJ^{+}09, WHK^{+}09]. safeguarding [LRR^{+}04]. savings [RBW08]. Scalable [MMBDK08, ADH^{+}01]. scale [Fre06, TB04]. scale-free [Fre06]. schedules [SW09]. scheduling [LLCM02]. scheme [BBG^{+}02, PPN09]. science [FZ09, Pyr03a, Ram05]. SDH [BIGKR07, GR03]. Seamless [AEM^{+}06, RFBJ04, BFL^{+}02, GMS^{+}00]. search [Mal04]. searching [AHS01]. second [Net00]. Secure [BBM^{+}09, She07, Bel04, CMCP06, GCSB07, IBPF08, Kol09]. Securing [WRL^{+}09]. Security [BLS06, PRS01, VWC^{+}04, AFG^{+}06, BBDK04, BFL04, BMM^{+}02, CP04, CC04, CP07, Flo06, GMW07, HMM07, KAGS09, LLZH09, MCP^{+}03, MVX^{+}04, MBC^{+}07, MWZ07, PMCR06, PS07, RTM^{+}04, RKR06, Sol07, SFK^{+}07, Vla03, Wan07, GM07]. selection [HHH05, HHH^{+}06]. selective [NDM^{+}06]. Self [BGW09, CPH09, Cla09, HSP03, NDC^{+}09]. self-deployment [Cla09]. self-managed [NDC^{+}09]. Self-Optimization [CPH09]. Self-organizing [BGW09, HSP03]. Semantic [Lar08, NDC^{+}09, Sun06]. semiconductor [DSA^{+}09, MGS^{+}05]. sensor [LDFK05, VVL^{+}09]. separation [KAGS09]. SEQUIN [TSC^{+}03]. server [FLM^{+}06, JGA04, RBW08]. servers [Bes06, DDHS01]. Service [BLP02, BGUV03, BU07, CGLU03, Cho04b, DDHS01, DPB09, KMS^{+}02, KRA06b, LKL00, Sch00, ABBST08, AMT06, AHKK06, BLS06, BLH^{+}05, Bar06, BGK^{+}06, BH08, BH05, BDE^{+}07, BCGM00, BKMV09, CR04, CRVX02, CDM^{+}07, CDL07, CJKS06, CMPW02, DEK^{+}05, DCHY00, ELM00, EM08, FK04, FZ04, GR03, GMS^{+}00, Hao08, Har05, HMR04, HLS09, HJM^{+}04, HMS^{+}07, KRLV08, Koc02, KRA06a, KUV03, KH02, LPW02, LS01, Liu08, LPST01, MGKO08, MvHdMvE03, MESG06, SMZ^{+}09, SBS00, SFK^{+}07, TAS^{+}07, Tor00, Tsa00, UGDS01, Van05, VDR08, VFRO08, VVL^{+}09, WPK^{+}07, ZJP03, vdGHM04, DHM^{+}03, Dos03, GR03, GL03, SPNR03, UVMES^{+}04]. service-based [AMT06]. service-delivery [SFK^{+}07]. service-level [MvHdMvE03]. service-mediation-enabled [DDHS01]. Service-oriented [BU07, DPB09, KRLV08]. Services [AHL^{+}02, Gol05, ARS^{+}06, AKN^{+}07, ABF^{+}07, ADH^{+}09, ABSZ06, ACEM04, BWT^{+}04, BMOP00, BBDK04, BBD^{+}00, BEM^{+}05, BLV00, BM03, BVYY07, BCSW09, BGTU05, BU07, BKMB06, BRG^{+}05, CRVX02, CCC^{+}02, CLC07, CCD^{+}00, CCH^{+}01, Dan08, DLM^{+}07, De 09, DCHY00, DHK00, DGJ02, DPB09, DNPQ01, DKK00, GHW08, GMS^{+}00, Hao08, HHVST07, HVKL03, HH04, HKH05, HB09, HKQ^{+}04, Koc02, KP09, KFL00, LPW02, LBF^{+}04, LPST01, MHH05, MBB00, MSSS00, NS03, PUV07, PDDS09, RTM^{+}04, RNMP08, SLKAM09, SSSK00, STB07, SJ01, Tor02, VDR09, VW08, VVL^{+}09, ZZLD04]. serving [MKL^{+}05]. Session [DGJ02, GL03, Zen04, CLW^{+}06, GJ04b, HMH04, TB08]. sessions [RPB^{+}04, SGRP04, vdGHM04]. SH [Rub09, SLKAM09, SLK^{+}09]. shared [ANPP04, ADH^{+}09, GL06, PPN09]. shared-experience [ADH^{+}09]. sharing [FYJB09]. shift [ELM00, Tor00]. short [CDL07]. shortest [RR01]. shortest-path [RR01]. SiC [FEK02]. SiC/GaN [FEK02]. Signaling [SLM02, SPNR03, IBPF08, Jos08, SR09]. signalling [CPH09]. significance [GHW08]. significance-aware [GHW08]. silicon [DRC^{+}05]. simple [DKLS01, LE09]. Simulation [Ahm02, HSF^{+}05, CEJ06, Lub00]. simulations [Bi05]. SIP [AG04, ACEM04, BMvB06, CE06, CEE04, CEL^{+}04, EFH04, FZ04, GCK01, GJ04a, HH04, HMH04, JGA04, LBF^{+}04, RFBJ04, RPB^{+}04, SGRP04, SJM^{+}04, UVMES^{+}04, ZJP03, ZZLD04, vdGHM04]. SIP-based [CEL^{+}04, LBF^{+}04, RFBJ04]. SIP-enabled [SJM^{+}04]. sites [Moo04]. sizing [Keb00]. SLAs [CCRV04]. slate [BBK^{+}09]. slits [CMC^{+}05]. small [DRC^{+}05, Sha07]. smart [BHH05]. SNMP [TSC^{+}03]. SNMP-based [TSC^{+}03]. Social [CV08]. soft [CGJ^{+}02]. Software [Hol00, Ram04, Aro06, ABC^{+}05, AMS00, BBD^{+}00, BEM^{+}05, BM03, BR03, BFMT00, HS00, JOR00, Keb00, KRA06a, LB07, MBB00, MW00, NS03, Sun05, Ded03, FFG^{+}07b]. Solar [LGTM02]. solution [ASK04, BFL^{+}02, HRB^{+}02, LRR^{+}04, Van05, Wis04, Yan02, KFL00]. solutions [AK03, CKR^{+}04, FHW01, GMW07, GCSB07, HPvV^{+}08, RWGS09]. solve [HHS07]. some [GJ04a]. SONET [BIGKR07, GR03]. SONET/SDH [BIGKR07, GR03]. space [Koc03, LLCM02, WDC^{+}05]. space-time [LLCM02]. Spanning [TB08]. spatial [BCG^{+}09, FYJB09]. spectral [DLM^{+}07, SLJ^{+}09]. spectrum [BKMV09, FYJB09, SŽVdLvW08, ŽKN^{+}08]. Speech [BBD^{+}00, KT07]. Speech-enabled [BBD^{+}00]. speed [BBC^{+}02, BDL09, BXK^{+}09, DGH^{+}02, GKP^{+}02, Vla03, ABK^{+}00]. SPIDER [DKLS01]. SPIRITS [KUV03]. SPIRITS-based [KUV03]. SS7 [Bru05]. standard [BGDT07, FEK02, KT07]. standardization [BFL04, QW05]. standards [ABF^{+}07, BGUV03, SSM^{+}00]. Starburst [SBC^{+}01]. Static [FS06]. station [BBK^{+}07, Fis07, KKL^{+}02]. stations [BBM^{+}09]. Statistical [CKR05]. statistics [JK02]. Status [LMMV06]. step [CEL^{+}04, NT09]. storage [DJM^{+}07, VDR08]. strategic [TA04]. strategies [BWT^{+}04, KSA08, Sch00]. strategy [BBG^{+}04, DER^{+}03, FFG^{+}07a, FFG^{+}07b, Wan07]. streaming [GL06, She07, SSC^{+}08, WPK^{+}07, ZS08]. structure [DEK^{+}05]. structured [SRR08]. study [Bi05, MA06, Ram07]. sub [Ton02, ANPP04]. Sub-50 [ANPP04]. sub-optimal [Ton02]. Subscriber [LMR^{+}06, Ma07, MPGS08, WAB^{+}08, ŽKN^{+}08]. subscribers [LCBM09]. subsystem [AHL^{+}02, Bar06, CLW^{+}06, CLC07, TAS^{+}07]. subwavelength [CMC^{+}05]. Success [GLM^{+}02]. SuperDHLR [IAM02]. superhydrophobic [KTKH05, LSF05]. Support [CEL^{+}04, BHH05, CDM^{+}07, CFL^{+}06, CLdJH08, Gol05, GKP^{+}02, MMD00, SGRP04, UGDS01, VDR09]. supported [Sie01]. Supporting [De 09, PZBT03, ACEM04, BGTU05, IBPF08]. surface [CMC^{+}05]. surfaces [KTKH05, LSF05]. survey [Gol06]. survivability [HKO^{+}04]. survivable [BWT^{+}04]. switch [BHH^{+}00, ME00]. switched [PBDS09, Wan01, XLVN01]. switches [GR03, PPN09]. switching [GEOT04, Ma07, MMD00, NDM^{+}06, NNT05, PPN09, SLM02]. System [DKR^{+}06, BCC^{+}07, Bas03, BCL^{+}02, Bi05, CDL07, DLM^{+}07, DHJ^{+}07, FFG^{+}07b, FTV06, Flo06, Fre06, GB00, GA00, HHS08, IB03, JK02, KMSW02, LKL00, MCP^{+}03, TSC^{+}03, WHK^{+}09, Yan02, BBTD05, EJV00, MBC^{+}07, SLM02]. systematic [RKR06]. systems [ASK04, AP05, Ale07, AKW00, BCG^{+}09, BBR^{+}09, BR03, BBC^{+}02, BVYY07, BCV^{+}07, BCG^{+}03, BCRL05, CDM^{+}07, DMO00, FYJB09, FHM^{+}05, Fre06, GKP^{+}02, HSF^{+}05, IGNC^{+}05, KTW00, KGT06, LGTM02, LCM^{+}02, Lub00, MDMF03, McP06, RFBJ04, RS09, SLKAM09, SLK^{+}09, SR09, Sun06].

T [Ram07]. tandem [Tsa00]. tasks [LWM08]. Tb [SLJ^{+}09]. Tb/s [PPN09, SLJ^{+}09]. TDM [Tsa00]. technical [AHL^{+}02, PDDS09, vVWBH05]. technique [BHT09, FYJB09]. Techniques [MR01, BXK^{+}09, HMM07, LS07, LGC06, SAN03, DDL08]. Technologies [Ano04c, Ano04e, Ano05a, Ano05b, AH08, Ano03b, Ano03c, Ano03d, Ano03e, Ano04b, Ano06b, Ano06c, Ano06d, EJV00, ESHKW06, FEK02, HSS^{+}09, HJM^{+}04, KKL^{+}02, KP09, LGC06, Maz06, Ano01c, Ano02c, Ano02d, Ano02e, Ano04d, Ano05c, Ano05d, Ano06a, OJ04, Rus04]. Technology [HPvV^{+}08, ZTH^{+}00, AMS00, BDL09, CGJ^{+}02, DK04, DPB09, DNPQ01, FZ09, FS06, GFB05, HRB^{+}02, JOR00, Koc02, MDMF03, PT07, Ram05, ZS08]. Telco [ADH^{+}09, SMZ^{+}09, SSC^{+}08]. telco-driven [SMZ^{+}09]. telecom [Koc03, OC08, SSC^{+}08, WXT^{+}08]. telecommunication [LMR09, Moo04, PWHH00]. telecommunications [BWT^{+}04, BFL04, BCC^{+}04, JMS04, Mac04, McP06, Pyr03b]. telephony [BDE^{+}07, CGV09, GM07, MMD00]. TelePortal^{T M} [BBD^{+}00]. temperature [JLD^{+}06]. term [BGDT07]. terminal [CKT08, SRR08]. terrestrial [AdLvWH^{+}07]. Test [MD02, BR03, FFG^{+}07a, FFG^{+}07b, GM06]. Testing [ABN^{+}03, AJK06, DFR06, Wan07]. their [ABF^{+}07, HC09, RGAK01]. theory [AAG^{+}09]. there [FLM^{+}06]. Thermal [HSK^{+}09, LMR09]. things [DRC^{+}05]. Third [BBDK04, BBC^{+}02, CGR02, CMPW02, DDHS01, FM08, KMSW02, KKL^{+}02, LPW02, SRR08, SBLD08]. Third-generation [BBDK04, BBC^{+}02, CMPW02, KMSW02, KKL^{+}02, LPW02]. third-party [CGR02, DDHS01, FM08]. threats [WRL^{+}09]. throughput [SR09]. Time [PBDS09, ABC^{+}05, CMMZ08, GEOT04, LLCM02, SW09, SLKAM09]. time-domain [SW09]. TMN [DPB09]. TMN-based [DPB09]. tolerant [NQR^{+}01]. tomorrow [HB09]. tool [CEJ06, DKLS01, DFS^{+}05]. tools [ABN^{+}03, CCE^{+}06, KKKC08]. topology [CEJ^{+}03, MMBDK08]. traceability [KKKC08]. trade [DDL08, LUV08, PS07]. trade-off [PS07]. trade-offs [DDL08, LUV08]. Tradeoff [RTM^{+}04]. tradeoffs [AMG05]. traffic [BCG^{+}03, DSW08, DNPQ01, JK02, MR01, NDC^{+}09, UVMES06, Wan01]. traffic-aware [NDC^{+}09]. transceivers [AP05, Sha07]. transcend [Wis04]. transformation [Koc07]. Transition [CFL^{+}06]. Transmission [CMC^{+}05, GSCF09, LGC06, SLJ^{+}09, WHK^{+}09]. transmissions [BHT09, BYA09]. transparent [CCE^{+}06]. Transport [GJ04b, BDL09, BXK^{+}09, ESHKW06, FTV06, GHVH06, HSF^{+}05, HLS09, KGT06, LE09, LS07, MC06, SSSK00, SPNR03, STB07, XLVN01]. treatment [CDL07]. trees [BBC^{+}09]. trends [BGTU05, DNPQ01]. trials [DFS^{+}05]. triple [ABF^{+}07]. triple-play [ABF^{+}07]. tunable [KTKH05]. TV [BDE^{+}07]. TV-link [BDE^{+}07]. types [ACEM04].

ultimate [MRD^{+}05]. Ultra

[BKMV09, FTV06]. Ultra-broadband [BKMV09]. UMTS [BBK^{+}07, BFL^{+}02, BJM^{+}02, DGH^{+}02, DC09, FHW01, KH02, SAN03, UGDS01, WPK^{+}07]. uncooled [JBR^{+}09]. Understanding [DMO00, GT07, MK01]. unified [CRVX02, Hao08]. unit [AMT03]. update [Ma07]. updates [Keb00]. updating [LB07]. upgrading [CP04]. use [Sha07]. User [BCG^{+}03, AG04, BCD^{+}08, BHT09, CEJ^{+}03, SRR08, CKT08]. User-level [BCG^{+}03].

User-terminal [CKT08]. users [BFL^{+}02, CRVX02, FM08, MRL07]. uses [ZJP03]. Using [GEOT04, MBC^{+}07, BMOP00, BBD^{+}00, BDE^{+}07, Bi05, Bru05, BRG^{+}05, DFS^{+}05, FDC^{+}07, GMW07, GMS^{+}00, HJvO^{+}06, HSP03, JK02, KTW00, KRA06b, LHMS05, LWM08, LCBM09, MvHdMvE03, WC08, EJV00].

value [ARS^{+}06, HMS^{+}07, IGNC^{+}05, RTM^{+}04]. value-added [ARS^{+}06, RTM^{+}04]. Variable [Lau00]. Variable-rate [Lau00]. various [Gol06]. VBA [JLN03]. vendor [BEM^{+}05, DFS^{+}05, HMR04]. vendor-independent [DFS^{+}05]. verification [ABN^{+}03, HS00, Hol00, MD04]. Versatile [BDM^{+}06, MGS^{+}05]. versus [Pyr03a]. very [DRC^{+}05]. via [BKMV09, SBLD08, TF07]. video [AdLvWH^{+}07, VSL^{+}08, GL06, NO08, VDR09, ZS08, vVWBH05]. VillageNet [DJR07]. Virtual [RGAK01, DJM^{+}07, HVKL03, MVJ^{+}01, RTM^{+}04, Tor02]. virtualization [CKT08]. virtualized [KRLV08]. vision [BBG^{+}04, SBS00]. voice [Bel04, BCV^{+}07, CCD^{+}00, DCHY00, DHK00, KFL00, MMD00, NO08, SSSK00]. voice-band [SSSK00]. voice/data [DCHY00, DHK00, KFL00]. VoiceXML [BBD^{+}00]. VoIP [AFRZ07, BH05, BHH05, DER^{+}03, FZ04, GMW07, HHH05, HHH^{+}06, HKUW03, LMR07, LS07, MHH05, PMCR06, PRT05, RFBJ04]. volume [DNPQ01]. VPN [BFL^{+}02, RTM^{+}04]. vs [PS07]. VSB [SLJ^{+}09]. vulnerabilities [CCRV04, GM07, Mac04, Wan07]. vulnerability [FHOS04].

WAN [DNPQ01]. WAP [BLV00]. WAP-enabled [BLV00]. Warranty [Wan06]. Wavelength [NDM^{+}06, SW09]. WaveTunnel [AGK^{+}03]. web [AHS01, EJV00, Lab08, ADH^{+}09, BU07, CCC^{+}02, LPW02]. wheels [AMBK^{+}07]. Wi [GMW07]. Wi-Fi [GMW07]. wide [BD00, RLP05]. WiMAX [JCL07]. WINNER [Ale07]. winning [Das00]. wired [PZBT03]. Wireless [BMM^{+}02, FDC^{+}07, HJM^{+}04, Mal04, PT07, SSM^{+}00, Van05, AMG05, AHL^{+}02, AAG^{+}09, AP05, BBDK04, BBR^{+}09, BHLN02, BBC^{+}02, BBG^{+}02, BYA09, BCG^{+}03, BRG^{+}05, CMCP06, CGLU03, CR04, CP04, Cla09, CLdJH08, EJV00, FHM^{+}05, GA00, HKH05, JJM^{+}05, KKL^{+}02, KH02, LGTM02, LMR07, LCBM09, LBF^{+}04, MZP^{+}09, MKL^{+}05, NT09, Pan06, PRS01, PZBT03, RFBJ04, SR09, Sch00, Sha04, SLW^{+}09, UVMES^{+}04, VK05, Wan00]. wireline [JJM^{+}05, LBF^{+}04]. wireline-wireless [JJM^{+}05]. wires [STB07]. within [MB03]. WLAN [BFL^{+}02]. working [CJKS06]. world [AH08, DFR06]. worm [She07].

X [MRD^{+}05]. X.805 [Ram07]. XML [BFMT00, LHMS05, MMBDK08, Wan07]. XML-based [Wan07]. XML/XSLT [LHMS05]. Xplorer [BRG^{+}05]. XSLT [LHMS05].

years [Net00].

References

Akyol:2009:DDC

[AAG^{+}09] Umut Akyol, Matthew Andrews, Piyush Gupta, John D. Hobby, Iraj Saniee, and Alexander Stolyar. Distributed dynamic control of multi-hop wireless networks: From theory to practice. Bell Labs Technical Journal, 14 (3):139–155, Autumn 2009. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Arlein:2008:ANF

[ABBE08] Robert M. Arlein, Stéphane Betgé-Brezetz, and J. Robert Ensor. Adaptive notification framework for converged environments. Bell Labs Technical Journal, 13 (2):155–159, Summer 2008. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Aghasaryan:2008:PEC

[ABBST08] Armen Aghasaryan, Stéphane Betgé-Brezetz, Christophe Senot, and Yann Toms. A profiling engine for converged service delivery platforms. Bell Labs Technical Journal, 13(2):93–103, Summer 2008. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Asghar:2005:ISC

[ABC^{+}05] Naeem Asghar, Randeep S.

Bhatia, Rakesh Chandwani, Colin Corcoran, Fang Hao, Joseph Karwisch, Pramod V. Koppol, T. V. Lakshman, Michael P. Siesta, and Stephen M. Zlatos. iOptimize: A software capability for analyzing and optimizing connection-oriented data networks in real time. Bell Labs Technical Journal, 9(4): 67–81, Winter 2005. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anderson:2007:ERA

[ABF^{+}07] Thomas W. Anderson, Peter Busschbach, Igor Faynberg, Hui-Lan Lu, and Dong Sun. The emerging resource and admission control function standards and their application to the new triple-play services. Bell Labs Technical Journal, 12 (1):5–21, Spring 2007. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Andrews:2000:SPP

[ABK^{+}00] Marilyn G. Andrews, Charles C. Byers, Prudence T. Z. Kapauan, Trent R. Nanke, and John L. Shepherd. The SPEED platform for programmable bearer channel processing. Bell Labs Technical Journal, 5(3):76– 97, Autumn 2000. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Alicherry:2003:TVN

[ABN^{+}03] Mansoor Alicherry, Randeep Singh Bhatia, Harsha Nagesh, Chitra Phadke, and Vishy Poosala. Testing and verification of network management and design tools. Bell Labs Technical Journal, 8(3):3–13, Autumn 2003. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Asthana:2006:EEA

[ABSZ06] Abhaya Asthana, Eric J.

Bauer, Meenakshi Sharma, and Xuemei Zhang. Endto-end availability considerations for services over IMS. Bell Labs Technical Journal, 11(3):199– 210, Autumn 2006. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Azada:2004:ISM

[ACEM04] Maria R. G. Azada, Jim Calme, Richard P. Ejzak, and John J. MacNamara. Issues in supporting multimedia services in SIP networks with mixed endpoint types. Bell Labs Technical Journal, 9(3): 183–197, Autumn 2004. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Austin:2001:FSD

[ADH^{+}01] Gary P. Austin, Bharat T.

Doshi, Christopher J. Hunt, Ramesh Nagarajan, and M. Akber Qureshi. Fast, scalable, and distributed restoration in general mesh optical networks. Bell Labs Technical Journal, 6(1):67– 81, Spring 2001. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Arlein:2009:TMW

[ADH^{+}09] Robert M. Arlein, Dennis R. Dams, Richard B. Hull, John Letourneau, and Kedar S. Namjoshi. Telco meets the Web: Programming sharedexperience services. Bell Labs Technical Journal, 14 (3):167–185, Autumn 2009. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Ashikhmin:2007:DDT

[AdLvWH^{+}07] Alexei Ashikhmin, Adriaan J. de Lind van Wijngaarden, Zhao Haibo, Bertrand M. Hochwald, Thomas L. Marzetta, Vinay Purohit, Chen Qinghong, Paul A. Wilford, Sheng-Rong Zhou, Michael A. Zuniga, and Edward S. Zuranski. Design and development of a terrestrial digital video broadcast demodulation core: An international collaborative effort. Bell Labs Technical Journal, 12(2):97– 118, Summer 2007. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Azada:2006:SMA

[AEM^{+}06] Maria R. G. Azada, Richard P. Ejzak, John J. MacNamara, Donna Sand, and Robin Thompson. Seamless mobility across IMS and legacy circuit networks. Bell Labs Technical Journal, 10(4):25– 38, Winter 2006. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anderlind:2006:IS

[AFG^{+}06] Erik E. Anderlind, David W.

Faucher, Eric H. Grosse, Daniel N. Heer, Andrew R. McGee, David P. Strand, and Robert J. Thornberry Jr. IMS security. Bell Labs Technical Journal, 11 (1):37–58, Spring 2006. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Abraham:2007:CV

[AFRZ07] Santosh Abraham, Peretz M.

Feder, Michael C. Recchione, and Honghai Zhang. The capacity of VoIP over 802.11. Bell Labs Technical Journal, 11(4):253–271, Winter 2007. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Arlein:2004:EFC

[AG04] Robert M. Arlein and Vijay K. Gurbani. An extensible framework for constructing SIP user agents. Bell Labs Technical Journal, 9(3): 87–100, Autumn 2004. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Alicherry:2003:WAD

[AGK^{+}03] Mansoor Alicherry, Sadanand Gogate, Rohit Khanna, Harsha Nagesh, Chitra Phadke, Vishy Poosala, and Sathya Srinivasan. WaveTunnel — Accurately designing deployable optical networks. Bell Labs Technical Journal, 8(1): 219–232, Spring 2003. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Aidan:2008:ACC

[AH08] Bruno Aïdan and Richard B. Hull. Applications for the converged communications world: Challenges and emerging technologies. Bell Labs Technical Journal, 13 (2):1–4, Summer 2008. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anupam:2006:ILS

[AHKK06] Vinod Anupam, Richard B. Hull, Sarbmeet S. Kanwal, and Bharat Kumar. An introduction to Lucent’s service enhancement layer. Bell Labs Technical Journal, 10 (4):179–196, Winter 2006. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Adamek:2002:STC

[AHL^{+}02] John G. Adamek, Eric H. Henrikson, Harold A. Lassers, Anne Y. Lee, and Ronald B. Martin. Services and technical considerations for the wireless IP multimedia subsystem. Bell Labs Technical Journal, 7(2):91–104, Summer 2002. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Ahmed:2002:SPE

[Ahm02] Walid K. M. Ahmed. Simulation and performance evaluation of IS-95 handoff algorithms. Bell Labs Technical Journal, 7(1):165– 181, Spring 2002. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

August:2001:MWS

[AHS01] Kit G. August, Mark H.

Hansen, and Elizabeth Shriver. Mobile web searching. Bell Labs Technical Journal, 6(2): 84–98, Summer 2001. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Aizenberg:2005:BIA

[Aiz05] Joanna Aizenberg. A Bioinspired approach to controlled crystallization at the nanoscale. Bell Labs Technical Journal, 10(3):129– 141, Autumn 2005. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Acevedo:2006:RGF

[AJK06] Pablo E. Acevedo, Donald S.

Jackson, and Robert W. Kotlowitz. Reliability growth and forecasting for critical hardware through accelerated life testing. Bell Labs Technical Journal, 11 (3):121–135, Autumn 2006. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Agarwal:2003:ECQ

[AK03] Avesh Kumar Agarwal and R. Krishna Kumar. An evaluation of current QoS solutions for mobile IP networks. Bell Labs Technical Journal, 8(2):3–13, Summer 2003. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Agrawal:2007:MIC

[AKN^{+}07] Shipra Agrawal, C. N. Kanthi, K. V. M. Naidu, Jeyashankher Ramamirtham, Rajeev Rastogi, Scott Satkin, and Anand Srinivasan. Monitoring infrastructure for converged networks and services. Bell Labs Technical Journal, 12(2):63–77, Summer 2007. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Alferness:2000:EOS

[AKW00] Rod C. Alferness, Herwig Kogelnik, and Thomas H. Wood. The evolution of optical systems: Optics everywhere. Bell Labs Technical Journal, 5(1):188– 202, Spring 2000. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Alexiou:2007:WDN

[Ale07] Angeliki Alexiou. WINNER:

Designing a new radio interface for next-generation systems. Bell Labs Technical Journal, 12(2):19–35, Summer 2007. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Abusch-Magder:2007:NNW

[AMBK^{+}07] David Abusch-Magder, Peter Bosch, Thierry E. Klein, Paul A. Polakos, Louis G. Samuel, and Harish Viswanathan. 911-NOW: A network on wheels for emergency response and disaster recovery operations. Bell Labs Technical Journal, 11(4): 113–133, Winter 2007. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Abusch-Magder:2005:NAE

[AMG05] David Abusch-Magder and John M. Graybeal. Novel algorithms for efficient exploration of the tradeoffs between cell count and performance in wireless networks. Bell Labs Technical Journal, 10(2):81–98, Summer 2005. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Atkins:2000:MTE

[AMS00] David L. Atkins, Audris Mockus, and Harvey P. Siy. Measuring technology effects on software change cost. Bell Labs Technical Journal, 5(2): 7–18, Summer 2000. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anderson:2003:OIU

[AMT03] Jeffrey G. Anderson, Steven L. Makowski, and Stanley Tamras. Optical interface unit — Bridge to the next-generation packet network. Bell Labs Technical Journal, 8(1):65– 81, Spring 2003. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Alfano:2006:ISB

[AMT06] Frank M. Alfano, Peter J. McCann, and Thomas T. Towle. IMS service-based bearer control. Bell Labs Technical Journal, 10(4): 151–166, Winter 2006. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2000:BLIa

[Ano00a]  Anonymous.  Bell Labs

innovation briefs. Bell Labs Technical Journal, 5 (1):215, Spring 2000. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2000:BLIb

[Ano00b] Anonymous. Bell Labs innovation briefs. Bell Labs Technical Journal, 5(2):189– 190, Summer 2000. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2000:BLIc

[Ano00c] Anonymous. Bell Labs innovation briefs. Bell Labs Technical Journal, 5(3):217– 218, Autumn 2000. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2000:BLId

[Ano00d] Anonymous. Bell Labs innovation briefs. Bell Labs Technical Journal, 5(4):185– 186, Winter 2000. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2001:BLIa

[Ano01a] Anonymous. Bell Labs innovation briefs. Bell Labs Technical Journal, 6(1):229– 230, Spring 2001. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2001:BLIb

[Ano01b] Anonymous. Bell Labs innovation briefs. Bell Labs Technical Journal, 6(2):170– 171, Summer 2001. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2001:RLT

[Ano01c] Anonymous. Recent Lucent Technologies patents. Bell Labs Technical Journal, 6(2): 172–197, Summer 2001. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2002:BLIa

[Ano02a] Anonymous. Bell Labs innovation briefs. Bell Labs Technical Journal, 7(1):227– 228, Spring 2002. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2002:BLIb

[Ano02b] Anonymous. Bell Labs innovation briefs. Bell Labs Technical Journal, 7(2):195– 196, Summer 2002. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2002:RLTa

[Ano02c] Anonymous. Recent Lucent Technologies patents. Bell Labs Technical Journal, 7(1): 229–240, Spring 2002. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2002:RLTb

[Ano02d] Anonymous. Recent Lucent Technologies patents. Bell Labs Technical Journal, 7(2):

197–205, Summer 2002. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2002:RLTc

[Ano02e] Anonymous. Recent Lucent Technologies patents. Bell Labs Technical Journal, 7(3): 225–238, Autumn 2002. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2003:BLI

[Ano03a] Anonymous. Bell Labs innovation briefs. Bell Labs Technical Journal, 8(2):129– 130, Summer 2003. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2003:RLTa

[Ano03b] Anonymous. Recent Lucent technologies patents. Bell Labs Technical Journal, 7(4): 187–196, Winter 2003. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2003:RLTb

[Ano03c] Anonymous. Recent Lucent technologies patents. Bell Labs Technical Journal, 8(1): 239–246, Spring 2003. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2003:RLTc

[Ano03d] Anonymous. Recent Lucent technologies patents. Bell Labs Technical Journal, 8(2): 131–139, Summer 2003. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2003:RLTd

[Ano03e] Anonymous. Recent Lucent technologies patents. Bell Labs Technical Journal, 8(3): 135–143, Autumn 2003. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2004:D

[Ano04a] Anonymous. Dedication. Bell Labs Technical Journal, 9(2): fmvii, Summer 2004. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2004:RLTa

[Ano04b] Anonymous. Recent Lucent technologies patents. Bell Labs Technical Journal, 8(4): 231–241, Winter 2004. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2004:RLTb

[Ano04c] Anonymous. Recent Lucent Technologies Patents. Bell Labs Technical Journal, 9(1): 211–218, Spring 2004. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2004:RLTc

[Ano04d] Anonymous. Recent Lucent Technologies patents. Bell Labs Technical Journal, 9(2): 205–213, Summer 2004. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2004:RLTd

[Ano04e] Anonymous. Recent Lucent Technologies Patents. Bell Labs Technical Journal, 9(3): 271–280, Autumn 2004. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2005:RLTa

[Ano05a] Anonymous. Recent Lucent Technologies Patents. Bell Labs Technical Journal, 9(4): 201–210, Winter 2005. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2005:RLTb

[Ano05b] Anonymous. Recent Lucent Technologies Patents. Bell Labs Technical Journal, 10(1):201–207, Spring 2005. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2005:RLTc

[Ano05c] Anonymous. Recent Lucent Technologies patents. Bell Labs Technical Journal, 10 (2):173–178, Summer 2005. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2005:RLTd

[Ano05d] Anonymous. Recent Lucent Technologies patents. Bell Labs Technical Journal, 10 (3):235–240, Autumn 2005. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2006:RLTa

[Ano06a] Anonymous. Recent Lucent Technologies patents. Bell Labs Technical Journal, 10 (4):271–277, Winter 2006. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2006:RLTb

[Ano06b] Anonymous. Recent lucent technologies patents. Bell Labs Technical Journal, 11 (1):215–221, Spring 2006. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2006:RLTc

[Ano06c] Anonymous. Recent lucent technologies patents. Bell Labs Technical Journal, 11 (2):241–248, Summer 2006. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2006:RLTd

[Ano06d] Anonymous. Recent lucent technologies patents. Bell Labs Technical Journal, 11 (3):237–245, Autumn 2006. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2007:RALa

[Ano07a] Anonymous. Recent Alcatel– Lucent patents. Bell Labs Technical Journal, 11(4): 327–333, Winter 2007. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2007:RALb

[Ano07b] Anonymous. Recent Alcatel– Lucent patents. Bell Labs Technical Journal, 12(1): 263–274, Spring 2007. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2007:RALc

[Ano07c] Anonymous. Recent Alcatel– Lucent patents. Bell Labs Technical Journal, 12(2): 161–171, Summer 2007. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2007:RALd

[Ano07d] Anonymous. Recent Alcatel– Lucent patents. Bell Labs Technical Journal, 12(3): 205–213, Autumn 2007. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2008:RALa

[Ano08a] Anonymous. Recent Alcatel– Lucent patents. Bell Labs Technical Journal, 12(4): 213–221, Winter 2008. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2008:RALb

[Ano08b] Anonymous. Recent Alcatel– Lucent patents. Bell Labs Technical Journal, 13(1): 263–272, Spring 2008. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2008:RALc

[Ano08c] Anonymous. Recent Alcatel– Lucent patents. Bell Labs Technical Journal, 13(2): 237–246, Summer 2008. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2008:RALd

[Ano08d] Anonymous. Recent Alcatel– Lucent patents. Bell Labs Technical Journal, 13(3): 167–178, Autumn 2008. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2009:RALa

[Ano09a] Anonymous. Recent Alcatel– Lucent patents. Bell Labs Technical Journal, 13(4): 277–286, Winter 2009. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2009:RALb

[Ano09b] Anonymous. Recent Alcatel– Lucent patents. Bell Labs Technical Journal, 14(1): 283–294, Spring 2009. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2009:RALc

[Ano09c] Anonymous. Recent Alcatel– Lucent patents. Bell Labs Technical Journal, 14(2): 251–260, Summer 2009. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:2009:RALd

[Ano09d] Anonymous. Recent Alcatel– Lucent patents. Bell Labs Technical Journal, 14(3): 207–216, Autumn 2009. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Alicherry:2004:FSM

[ANPP04] Mansoor Alicherry, Harsha Nagesh, Chitra Phadke, and Vishy Poosala. FASTeR: Sub-50 ms shared restoration in mesh networks. Bell Labs Technical Journal, 9(1): 99–110, Spring 2004. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Alexiou:2005:RMT

[AP05] Angeliki Alexiou and Constantinos B. Papadias. Reconfigurable MIMO transceivers for next-generation wireless systems. Bell Labs Technical Journal, 10(2):139– 156, Summer 2005. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Arora:2006:DCP

[Aro06] Shivani Arora. Designing communication protocol software for performance. Bell Labs Technical Journal, 11 (1):203–207, Spring 2006. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Acharya:2006:BRP

[ARS^{+}06] Swarup Acharya, Pankaj Risbood, Anurag Srivastava, Vijay P. Kumar, and Walter Rothkegel. Beyond rapid provisioning: Driving valueadded services on control plane foundations. Bell Labs Technical Journal, 11 (2):203–214, Summer 2006. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Agron:2004:CDI

[ASK04] Minerva Agrón, Ronald C. Silacci, and Peter H. M. Klein. Challenges in designing and integrating a multioperations systems solution.

Bell Labs Technical Journal, 9(1):111–120, Spring 2004. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Austin:2006:EOL

[Aus06] Gary P. Austin. Evolving optical layer networking: What’s happening on the ground. Bell Labs Technical Journal, 11(2):3– 12, Summer 2006. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Barnett:2006:ENS

[Bar06] William J. Barnett, Jr. Enabling new service provider business models with the IP multimedia subsystem. Bell Labs Technical Journal, 10 (4):7–15, Winter 2006. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Bass:2003:CAN

[Bas03] Neil R. Bass. Component architecture in a network management system. Bell Labs Technical Journal, 8(3):51– 65, Autumn 2003. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Bi:2002:PDT

[BBC^{+}02] Qi Bi, Ronald R. Brown, Dongzhe Cui, Asif D. Gandhi, Ching-Yao Huang, and Stan Vitebsky. Performance of 1xEV-DO third-generation wireless high-speed data systems. Bell Labs Technical Journal, 7(3):97–107, Autumn 2002. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Bejerano:2009:RMN

[BBC^{+}09] Yigal Bejerano, Italo Busi, Laurent Ciavaglia, Enrique Hernandez-Valencia, Pramod Koppol, Vincenzo Sestito, and Martin Vigoureux. Resilient multipoint networks based on redundant trees. Bell Labs Technical Journal, 14(2):113–130, Summer 2009. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Ball:2000:SES

[BBD^{+}00] Thomas Ball, Veta Bonnewell, Peter Danielsen, Peter Mataga, and Kenneth Rehor. Speech-enabled services using TelePortal^{T M} software and VoiceXML. Bell Labs Technical Journal, 5(3):98– 111, Autumn 2000. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Borst:2005:DOF

[BBD^{+}05] Simon C. Borst, Arumugam Buvaneswari, Lawrence M. Drabeck, Michael J. Flanagan, John M. Graybeal, Georg K. Hampel, Mark Haner, William M. Mac-Donald, Paul A. Polakos, George Rittenhouse, Iraj Saniee, Alan Weiss, and Philip A. Whiting. Dynamic optimization in future cellular networks. Bell Labs Technical Journal, 10(2):99– 119, Summer 2005. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Balachandran:2004:TGW

[BBDK04] Krishna Balachandran, Kenneth C. Budka, Tewfik L. Doumi, and Joseph H. Kang. Third-generation wireless services for homeland security. Bell Labs Technical Journal, 9(2):5– 21, Summer 2004. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Beck:2008:NCC

[BBE^{+}08] Andre Beck, Sem Borst, Bob Ensor, Jairo O. Esteban, Volker Hilt, Ivica Rimac, and Anwar Walid. New challenges in content dissemination networks. Bell Labs Technical Journal, 13(3):5– 11, Autumn 2008. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Blumenthal:2002:SAD

[BBG^{+}02] Uri Blumenthal, Milind M.

Buddhikot, Juan A. Garay, Scott C. Miller, Sarvar Patel, Luca Salgarelli, and Dorothy Stanley. A scheme for authentication and dynamic key exchange in wireless networks. Bell Labs Technical Journal, 7(2):37– 48, Summer 2002. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Battaglia:2004:MPV

[BBG^{+}04] Patricia A. Battaglia, Charles C. Byers, Leslie A. Guth, Albert Holliday, Claudio Spinelli, and Jason J. Tong. Modular platform vision and strategy. Bell Labs Technical Journal, 9(1):121–142, Spring 2004. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Bauer:2007:UBS

[BBK^{+}07] Markus Bauer, Peter Bosch, Nidal Khrais, Louis G. Samuel, and Peter Schefczik. The UMTS base station router. Bell Labs Technical Journal, 11(4):93–111, Winter 2007. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Banniza:2009:EAC

[BBK^{+}09] Thomas-Rolf Banniza, Dietrich Boettle, Ralf Klotsche, Peter Schefczik, Michael Soellner, and Klaus Wuenstel. A European approach to a clean slate design for the future Internet. Bell Labs Technical Journal, 14 (2):5–22, Summer 2009. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Baryshnikov:2008:MMD

[BBLS08] Yuliy Baryshnikov, James Borger, Wonsuck Lee, and Abdol Saleh. Modeling market dynamics in competitive communication consumer markets. Bell Labs Technical Journal, 13(2): 193–208, Summer 2008. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Bosch:2009:SBS

[BBM^{+}09] Peter Bosch, Alec Brusilovsky, Rae McLellan, Sape Mullender, and Paul Polakos. Secure base stations. Bell Labs Technical Journal, 13 (4):227–243, Winter 2009. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Balachandran:2009:PAN

[BBR^{+}09] Krishna Balachandran, Qi Bi, Ashok Rudrapatna, James Seymour, Robert Soni, and Andreas Weber. Performance assessment of nextgeneration wireless mobile systems. Bell Labs Technical Journal, 13(4):35–58, Winter 2009. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Byrnes:2005:EKR

[BBTD05] Michael T. Byrnes, James M. Bromfield, Dogan Taskent, and Patricia M. Doane. The Experium^{T M} Knowledge Resource System — Brainpower as competitive advantage. Bell Labs Technical Journal, 9(4):145–154, Winter 2005. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Beyeler:2004:IIM

[BCC^{+}04] Walter E. Beyeler, Stephen H. Conrad, Thomas F. Corbet, Gerard P. O’Reilly, and David D. Picklesimer. Inter-infrastructure modeling — ports and telecommunications. Bell Labs Technical Journal, 9(2):91– 105, Summer 2004. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Balachandran:2007:DAI

[BCC^{+}07] Krishna Balachandran, Doru Calin, Fang-Chen Cheng, Niranjan Joshi, Joseph H. Kang, Achilles Kogiantis, Kurt Rausch, Ashok Rudrapatna, James P. Seymour, and Jonqyin Sun. Design and analysis of an IEEE 802.16e-based OFDMA communication system. Bell Labs Technical Journal, 11 (4):53–73, Winter 2007. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Bila:2008:INA

[BCD^{+}08] Nilton Bila, Jin Cao, Robert Dinoff, Tin Kam Ho, Bharat Kumar, Daniel Lieuwen, and Paulo Santos. Intuitive network applications: Learning user context and behavior. Bell Labs Technical Journal, 13(2):31–47, Summer 2008. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Borst:2003:ULQ

[BCG^{+}03] Simon C. Borst, Kenneth L.

Clarkson, John M. Graybeal, Harish Viswanathan, and Philip A. Whiting. Userlevel QoS and traffic engineering for 3G wireless 1xEV-DO systems. Bell Labs Technical Journal, 8(2):33– 47, Summer 2003. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Balachandran:2009:DPA

[BCG^{+}09] Krishna Balachandran, Doru Calin, Nandu Gopalakrishnan, Joseph H. Kang, Achilles Kogiantis, Shupeng Li, Lawrence Ozarow, Sudhir Ramakrishna, Ashok N. Rudrapatna, and Russell Sun. Design and performance analysis of collaborative spatial multiplexing for IEEE 802.16e-based systems. Bell Labs Technical Journal, 13(4):97–117, Winter 2009. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Brenner:2000:CSN

[BCGM00] Michael R. Brenner, Ming Chu, George Gross, and Manu Malek. CyberCarrier service and network management. Bell Labs Technical Journal, 5(4):44– 62, Winter 2000. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Bi:2002:FLP

[BCL^{+}02] Qi Bi, Pi-Chun Chen, Patrick Li, Stan Vitebsky, and Yang Yang. Forward link performance analysis of the CDMA2000 3G1X data system. Bell Labs Technical Journal, 7(3):83–96, Autumn 2002. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Busschbach:2005:RMS

[BCRL05] Peter Busschbach, Steve Corum, Humberto La Roche, and Patrice Lamy. The role of management systems in optical/packet convergence. Bell Labs Technical Journal, 10(1):105–117, Spring 2005. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Borst:2009:NFP

[BCSW09] Sem Borst, Jason Collins, Iraj Saniee, and Anwar Walid. A new framework for performance analysis of emerging personalized communication services. Bell Labs Technical Journal, 14 (3):127–138, Autumn 2009. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Bi:2007:PED

[BCV^{+}07] Qi Bi, Pi-Chun Chen, Stan Vitebsky, Yang Yang, and Qinqing Zhang. Performance of 1x EV-do revision a systems with best effort data and voice over IP. Bell Labs Technical Journal, 11 (4):217–235, Winter 2007. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Berde:2009:NMR

[BCV09] Bela Berde, Agostino Chiosi, and Dominique Verchere. Networks meet the requirements of grid applications. Bell Labs Technical Journal, 14(1):173–184, Spring 2009. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Bosco:2000:EWA

[BD00] Harry L. Bosco and Douglas C. Dowden. Evolution of the wide area network. Bell Labs Technical Journal, 5(1): 46–72, Spring 2000. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Beck:2007:BTI

[BDE^{+}07] Andre Beck, Robert E. Daugherty, J. Robert Ensor, Jairo O. Esteban, F. Theodore Freuler, Sugato Ganguly, Kristin F. Kocan, and William D. Roome. Blending telephony and IPTV: Building the TV-link service package using the Alcatel–Lucent Service Broker^{T M}. Bell Labs Technical Journal, 12(1): 23–39, Spring 2007. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Buchali:2009:OOP

[BDL09] Fred Buchali, Roman Dischler, and Xiang Liu. Optical OFDM: A promising high-speed optical transport technology. Bell Labs Technical Journal, 14(1):125– 146, Spring 2009. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Berger:2006:VBM

[BDM^{+}06] Matthias Berger, Klaus Dechet, Gerhard Meyer, Walter Rothkegel, Wolfram Sturm, and Bjoern Wilke. Versatile bandwidth management: The design, development, and deployment of LambdaUnite(R). Bell Labs Technical Journal, 11(2):65– 81, Summer 2006. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Bauer:2007:IDA

[BDSS07] Markus Bauer, Armin Dekorsy, Peter Schefczik, and Michael Soellner. IP-driven access-independent resource management in converged access networks. Bell Labs Technical Journal, 12(2):37– 61, Summer 2007. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Bellows:2004:ASO

[Bel04] Douglas H. Bellows. Architectures for secure operation and maintenance of converged voice and data networks. Bell Labs Technical Journal, 8(4):55–59, Winter 2004. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Bass:2005:RSD

[BEM^{+}05] Elijah W. Bass, Bilgehan Erman, David P. Mongeau, Mengqi Wu, and Chen Xie. Rapid software development for multi-vendor services. Bell Labs Technical Journal, 9(4):155–169, Winter 2005. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Bessis:2006:IPR

[Bes06]  Thierry Bessis.  Improv-

ing performance and reliability of an IMS network by co-locating IMS servers. Bell Labs Technical Journal, 10(4):167–178, Winter 2006. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Bessis:2007:IDM

[Bes07] Thierry Bessis. Improving the DNS mechanism in a data center intranet. Bell Labs Technical Journal, 12 (1):131–144, Spring 2007. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Bessis:2008:OI

[Bes08] Thierry Bessis. Open IMS. Bell Labs Technical Journal, 13(3):159–166, Autumn 2008. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Bauer:2006:FAC

[BF06] Eric J. Bauer and Paul H. Franklin. Framework for availability characterization by analyzing outage durations. Bell Labs Technical Journal, 11(3):39–46, Autumn 2006. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Benenati:2002:SMV

[BFL^{+}02] David Benenati, Peretz M. Feder, Nancy Y. Lee, Silvia Martin-Leon, and Reuven Shapira. A seamless mobile VPN data solution for CDMA2000 , UMTS, and WLAN users. Bell Labs Technical Journal, 7(2):143– 165, Summer 2002. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Bertine:2004:ODT

[BFL04] Herbert Bertine, Igor Faynberg, and Hui-Lan Lu. Overview of data and telecommunications security standardization efforts in ISO, IEC, ITU, and IETF. Bell Labs Technical Journal, 8(4): 203–229, Winter 2004. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Bruns:2000:ASD

[BFMT00] Glenn R. Bruns, Alan E.

Frey, Peter A. Mataga, and Susan J. Tripp. Automated software development with XML and the Java language. Bell Labs Technical Journal, 5(2):32–43, Summer 2000. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Bachl:2007:LTE

[BGDT07] Rainer Bachl, Peter Gunreben, Suman Das, and Said Tatesh. The long term evolution towards a new 3GPP air interface standard. Bell Labs Technical Journal, 11 (4):25–51, Winter 2007. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Batni:2006:EMS

[BGK^{+}06] Ramachendra Batni, Katherine Guo, Anand Kagalkar, Sarit Mukherjee, and Ranjan Sharma. Evolution of MiRingBack service to multimedia MiRingBack service in the IMS framework. Bell Labs Technical Journal, 10 (4):53–70, Winter 2006. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Brenner:2005:OMA

[BGTU05] Michael R. Brenner, Michel L. F. Grech, Mohammad Torabi, and Musa R. Unmehopa. The Open Mobile Alliance and trends in supporting the mobile services industry. Bell Labs Technical Journal, 10(1):59– 75, Spring 2005. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Bennett:2003:SMS

[BGUV03] Andrew J. Bennett, Michel L. F. Grech, Musa R. Unmehopa, and Kumar V. Vemuri. Service mediation standards. Bell Labs Technical Journal, 7(4):77– 90, Winter 2003. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Borst:2009:SOA

[BGW09] Sem Borst, Varun Gupta, and Anwar Walid. Selforganizing algorithms for cache cooperation in content distribution networks. Bell Labs Technical Journal, 14 (3):113–125, Autumn 2009. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Beck:2005:ESM

[BH05] Andre Beck and Markus Hofmann. Extending service mediation to intelligent VoIP endpoints. Bell Labs Technical Journal, 10 (1):11–15, Spring 2005. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Batni:2008:AEA

[BH08] Ramachendra P. Batni and John F. Heck. An analytics engine architecture for service provider deployments. Bell Labs Technical Journal, 13(3):129– 141, Autumn 2008. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Barshefsky:2000:RFS

[BHH^{+}00] Alvin Barshefsky, Fred Hu, Sean Hu, Lisa Bergen Ohm, and Hemi Trickey. 5ESS(R) field switch performance capture. Bell Labs Technical Journal, 5(2):19–31, Summer 2000. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Beck:2005:NSS

[BHH05] Andre Beck, Volker Hilt, and Markus Hofmann. Network support for smart VoIP endpoints. Bell Labs Technical Journal, 10(1): 5–9, Spring 2005. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Bi:2002:MWD

[BHLN02] Qi Bi, Ching-Yao Huang, Patrick Li, and Mark E. Newbury. Measures of wireless data performance. Bell Labs Technical Journal, 7(3):219– 223, Autumn 2002. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Boccardi:2009:NOP

[BHT09] Federico Boccardi, Howard Huang, and Matteo Trivellato. A near-optimum precoding technique for downlink multi-user MIMO transmissions. Bell Labs Technical Journal, 13(4):79–95, Winter 2009. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Bi:2005:FLP

[Bi05]  Qi Bi.  A forward link

performance study of the 1x EV-do Rev. 0 system using field measurements and simulations. Bell Labs Technical Journal, 10(2):5– 19, Summer 2005. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Borger:2007:EES

[BIGKR07] Jim Borger, Alina Ionescu-Graff, Samrat Kulkarni, and Narayan Raman. Economics of Ethernet over SONET/ SDH. Bell Labs Technical Journal, 12(1):187– 206, Spring 2007. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Bishop:2005:NEM

[Bis05a] David Bishop. Nanotechnology and the end of Moore’s Law? Bell Labs Technical Journal, 10(3):23– 28, Autumn 2005. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Bishop:2005:O

[Bis05b] David Bishop. Overview.

Bell Labs Technical Journal, 10(3):1–3, Autumn 2005. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Brueck:2007:RRM

[BJK^{+}07] Stefan Brueck, Enrico Jugl, Hans-Juergen Kettschau, Michael Link, Jens Mueckenheim, and Andrei Zaporozhets. Radio resource management in HSDPA and HSUPA. Bell Labs Technical Journal, 11(4):151– 167, Winter 2007. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Bernhard:2002:IMR

[BJM^{+}02] Urs Bernhard, Enrico Jugl, Jens Mueckenheim, Holger Pampel, and Michael Soellner. Intelligent management of radio resources in UMTS access networks. Bell Labs Technical Journal, 7(3):109– 126, Autumn 2002. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Brinkman:2000:LBC

[BKLW00] William F. Brinkman, Thomas L. Koch, David V. Lang, and Daniel P. Wilt. The lasers behind the communications revolution. Bell Labs Technical Journal, 5(1): 150–167, Spring 2000. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Brok:2006:ENS

[BKMB06] Jacco Brok, Bharat Kumar, Erik Meeuwissen, and Harold J. Batteram. Enabling new services by exploiting presence and context information in IMS. Bell Labs Technical Journal, 10(4):83– 100, Winter 2006. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Buddhikot:2009:UBF

[BKMV09] Milind M. Buddhikot, Irwin Kennedy, Frank Mullany, and Harish Viswanathan. Ultrabroadband femtocells via opportunistic reuse of multioperator and multi-service spectrum. Bell Labs Technical Journal, 13(4):129– 143, Winter 2009. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Barker:2005:PMP

[BLH^{+}05] William O. Barker, Juliet R.

Lane, David P. Holbrook, Narasimha R. Vadrevu, and Lorraine T. Padalino. Preventative maintenance: A proactive customer service. Bell Labs Technical Journal, 9(4):187–200, Winter 2005. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Banke:2002:SAC

[BLP02] Richard T. Banke, Thomas K.

Lybarger, and Michael L. Patton. Service assurance for converged networks. Bell Labs Technical Journal, 7(1): 99–114, Spring 2002. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Balyan:2006:SAI

[BLS06] Avneesh Balyan, Karthic Loganathan, and Sridhar Sripathi. Security architecture for IP-based multiservice networks. Bell Labs Technical Journal, 11(1): 59–78, Spring 2006. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Batni:2000:ESW

[BLV00] Ramachendra P. Batni, Chinmei C. Lee, and Douglas W. Varney. Enhanced services in WAP-enabled networks. Bell Labs Technical Journal, 5(3):145–152, Autumn 2000. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Batteram:2003:OSP

[BM03] Harold J. Batteram and Erik Meeuwissen. Octavia: A software platform for nextgeneration services. Bell Labs Technical Journal, 7(4):139– 154, Winter 2003. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Blumenthal:2002:WNS

[BMM^{+}02] Uri Blumenthal, Michael Marcovici, Semyon Mizikovsky, Sarvar Patel, Ganapathy S. Sundaram, and Marcus Wong. Wireless network security architecture. Bell Labs Technical Journal, 7(2): 19–36, Summer 2002. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Bear:2000:OPN

[BMN^{+}00] Christopher L. Bear, Warren A. Montgomery, Diana B. Nolte, Suzanne K. Rajchel, and Michael C. Silva. Open, programmable networks. Bell Labs Technical Journal, 5(3): 30–42, Autumn 2000. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Bakker:2000:RDD

[BMOP00] John-Luc Bakker, Judith R.

McGoogan, William F. Opdyke, and Frans Panken. Rapid development and delivery of converged services using APIs. Bell Labs Technical Journal, 5(3):12– 29, Autumn 2000. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Batteram:2006:SMP

[BMvB06] Harold Batteram, Erik Meeuwissen, and Jeroen van Bemmel. SIP message prioritization and its applications. Bell Labs Technical Journal, 11 (1):21–36, Spring 2006. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Batteram:2003:TFC

[BR03] Harold J. Batteram and Willem A. Romijn. A test framework for CORBA component model-based software systems. Bell Labs Technical Journal, 8(3):15– 29, Autumn 2003. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Buvaneswari:2005:NOM

[BRG^{+}05]  Arumugam  Buvaneswari,

Bhanumathi Ravishankar, John M. Graybeal, Mark Haner, and Gee Rittenhouse. New optimization and management services for 3G wireless networks using CEL-NET Xplorer. Bell Labs Technical Journal, 9(4):101– 115, Winter 2005. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Baker:2000:BLI

[BRMS00] William O. Baker, Ian M. Ross, John S. Mayo, and Daniel C. Stanzione. Bell Labs innovations in recent decades. Bell Labs Technical Journal, 5(1):3– 16, Spring 2000. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Brugman:2005:CRA

[Bru05] David Brugman. Call-routing analysis using SS7 data. Bell Labs Technical Journal, 9(4): 133–138, Winter 2005. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Brenner:2007:SOA

[BU07] Michael R. Brenner and Musa R. Unmehopa. Serviceoriented architecture and Web services penetration in next-generation networks. Bell Labs Technical Journal, 12(2):147–159, Summer 2007. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Burns:2008:DHB

[Bur08] Michael J. Burns. Deploying, hardening, and benefiting from next-generation networks. Bell Labs Technical Journal, 12(4):1–3, Winter 2008. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Bennett:2003:PPM

[BUV03] Andrew J. Bennett, Musa R.

Unmehopa, and Kumar V. Vemuri. The Parlay proxy manager — Architecture considerations. Bell Labs Technical Journal, 7(4):91– 104, Winter 2003. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Bi:2007:BME

[BVYY07] Qi Bi, Stan Vitebsky, Yang Yang, and Yifei Yuan.

Broadcast-multicast in 1x EV-do Revision A systems and its application to enhanced mobile services. Bell Labs Technical Journal, 11 (4):237–252, Winter 2007. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Baker:2004:ADR

[BWT^{+}04] Mark C. Baker, Charles A. Witschorik, Jonathan C. Tuch, Waverly Hagey-Espie, and Veena B. Mendiratta. Architectures and disaster recovery strategies for survivable telecommunications services. Bell Labs Technical Journal, 9(2):125–145, Summer 2004. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Bulow:2009:PCM

[BXK^{+}09] Henning Bülow, Chongjin Xie, Axel Klekamp, Xiang Liu, and Bernd Franz. PMD Compensation/mitigation techniques for high-speed optical transport. Bell Labs Technical Journal, 14(1): 105–124, Spring 2009. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Boccardi:2009:RAM

[BYA09] Federico Boccardi, Kai Yu, and Angeliki Alexiou. Relayaided multiple antenna transmissions for wireless backhaul applications. Bell Labs Technical Journal, 13(4): 161–173, Winter 2009. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Chu:2004:IFA

[CC04] Chi-Hung Kelvin Chu and Ming Chu. An integrated framework for the assessment of network operations, reliability, and security. Bell Labs Technical Journal, 8(4): 133–152, Winter 2004. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Chen:2002:WCS

[CCC^{+}02] Yang Chen, Olivier B.

Clarisse, Pascal Collet, Mark A. Hartman, Luciano Rodriguez, Lizette Velazquez, and Bruce A. Westergren. Web communication services and the Packet IN(R) application hosting environment. Bell Labs Technical Journal, 7(1): 25–40, Spring 2002. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Clarisse:2000:PSE

[CCD^{+}00] Olivier B. Clarisse, Pascal Collet, James P. Dunn, Bruce A. Westergren, and Lizette Velazquez. Portal services: An evolution of voice features. Bell Labs Technical Journal, 5(3):203– 215, Autumn 2000. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Chekuri:2006:DTT

[CCE^{+}06] Chandra Chekuri, Paul Claisse, René-Jean Essiambre, Steven Fortune, Daniel C. Kilper, Wonsuck Lee, Nachi K. Nithi, Iraj Saniee, Bruce Shepherd, Christopher A. White, Gordon Wilfong, and Lisa Zhang. Design tools for transparent optical networks. Bell Labs Technical Journal, 11 (2):129–143, Summer 2006. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Colbert:2001:ASC

[CCH^{+}01] Raymond O. Colbert, Diane S. Compton, Randy L. Hackbarth, James D. Herbsleb, Laurie A. Hoadley, and Graham J. Wills. Advanced services: Changing how we communicate. Bell Labs Technical Journal, 6(1): 211–228, Spring 2001. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Chan:2004:RSR

[CCRV04] Chun K. Chan, Uma Chandrashekhar, Steven H. Richman, and S. Rao Vasireddy. The role of SLAs in reducing vulnerabilities and recovering from disasters. Bell Labs Technical Journal, 9(2):

189–203, Summer 2004. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Cheng:2007:CET

[CDL07] Terry S. Cheng, Kenneth W.

Del Signore, and Frances Lin. Cost effective treatment of short message service on the paging channel for the CDMA system. Bell Labs Technical Journal, 12 (2):133–141, Summer 2007. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Chen:2007:QSS

[CDM^{+}07] Pi-Chun Chen, Ren Da, Chris Mooney, Yang Yang, Qinqing Zhang, Lily H. Zhu, and Jialin Zou. Quality of service support in 1x EVdo Revision A systems. Bell Labs Technical Journal, 11 (4):169–184, Winter 2007. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Calme:2006:CSP

[CE06] James A. Calme and Richard P.

Ejzak. A common SIP profile for next-generation networks. Bell Labs Technical Journal, 11(1):107–122, Spring 2006. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Cortes:2004:SP

[CEE04] Mauricio Cortes, J. Robert Ensor, and Jairo O. Esteban. On SIP performance. Bell Labs Technical Journal, 9(3): 155–172, Autumn 2004. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Cheng:2003:NEC

[CEJ^{+}03] Lily Cheng, John Ellson, Admela Jukan, Patrice Lamy, and Eve Varma. Network engineering — Control of dynamic link topology in user networks. Bell Labs Technical Journal, 8(1):207– 218, Spring 2003. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Cortes:2006:DIS

[CEJ06] Mauricio Cortes, Jairo O. Esteban, and Hyewon Jun. Diabelli: An IMS simulation tool. Bell Labs Technical Journal, 10(4):255– 259, Winter 2006. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Cyr:2004:SCC

[CEL^{+}04] Bernard L. Cyr, Richard P. Ejzak, Joseph J. Lichter, John J. MacNamara, Thomas L. McRoberts, and Joseph E. Seitz. Support for CDMA circuit mobiles in a SIP-based packet core network: A step on the path to IMS. Bell Labs Technical Journal, 9(3): 217–236, Autumn 2004. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Choi:2006:TIS

[CFL^{+}06] Dupyo Choi, Christine Fischer, Anne Y. Lee, E-Ling Lou, Chung-Zin Liu, and Hsien-Chuen Yu. Transition to IPv6 and support for IPv4/IPv6 interoperability in IMS. Bell Labs Technical Journal, 10 (4):261–270, Winter 2006. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Chrin:2002:PSP

[CGJ^{+}02] Christopher Chrin, Steven L.

Gay, Mark H. Jones, Jeffrey C. Martin, and Kristin F. Kocan. Performance of soft phones and advances in associated technology. Bell Labs Technical Journal, 7(1): 135–139, Spring 2002. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Cannell:2003:SCN

[CGLU03] Lynell E. Cannell, Michel L. F. Grech, Anne Y. Lee, and Musa R. Unmehopa. Service control for next-generation applications in wireless IP multimedia networks. Bell Labs Technical Journal, 8(1):27– 42, Spring 2003. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Chiang:2002:NTP

[CGR02] Tsun-Chieh Chiang, Vijay K. Gurbani, and John B. Reid. The need for third-party call control. Bell Labs Technical Journal, 7(1):41– 46, Spring 2002. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Chow:2009:ADN

[CGV09] Stanley T. Chow, Christophe Gustave, and Dmitri Vinokurov. Authenticating displayed names in telephony. Bell Labs Technical Journal, 14(1):267–282, Spring 2009. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Choudhary:2004:PBN

[Cho04a] Abdur Rahim Choudhary. Policy-based network management. Bell Labs Technical Journal, 9(1):19–29, Spring 2004. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Choudhary:2004:SIT

[Cho04b] Abdur Rahim Choudhary. Service intelligence through agile information controls. Bell Labs Technical Journal, 8(4):61–70, Winter 2004. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Claussen:2008:OFC

[CHS08] Holger Claussen, Lester T. W. Ho, and Louis G. Samuel. An overview of the femtocell concept. Bell Labs Technical Journal, 13 (1):221–245, Spring 2008. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Chiang:2006:FNW

[CJKS06] Tung Ching Chiang, Andrew P. Johnson, Sarbmeet S. Kanwal, and Farhan Z. Shaikh. Friends night out — A working prototype of a blended lifestyle service enabled through IMS. Bell Labs Technical Journal, 10 (4):17–23, Winter 2006. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Cai:2004:AMM

[CKR^{+}04] Yigang Cai, Jack Kozik, Helmut L. Raether, John B. Reid, Guy H. Starner, Sunil Thadani, and Kumar V. Vemuri. Authorization mechanisms for mobile commerce implementations in enhanced prepaid solutions. Bell Labs Technical Journal, 8(4):121– 131, Winter 2004. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Cheung:2005:SAF

[CKR05] Benjamin Cheung, Gopal Kumar, and Sudarshan A. Rao. Statistical algorithms in fault detection and prediction: Toward a healthier network. Bell Labs Technical Journal, 9(4):171– 185, Winter 2005. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Conte:2008:NHA

[CKT08] Alberto Conte, Sylvaine Kerboeuf, and Laurent Thomas. Network-hosted avatar: Userterminal virtualization in the network. Bell Labs Technical Journal, 13(2):117– 126, Summer 2008. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Claussen:2009:ASD

[Cla09]  Holger Claussen.  Au-

tonomous self-deployment of wireless access networks. Bell Labs Technical Journal, 14 (1):55–72, Spring 2009. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Chen:2007:IMS

[CLC07] Sheng Chen, Xiangyang Li, and Yigang Cai. IP multimedia subsystem converged call control services. Bell Labs Technical Journal, 12 (1):145–160, Spring 2007.

CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Conte:2008:NSI

[CLdJH08] Alberto Conte, Brice Leclerc, and Federico de Juan Huarte. Native support of idle mode in wireless IP networks. Bell Labs Technical Journal, 13 (2):127–133, Summer 2008. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Conrad:2006:CNI

[CLOU06] Stephen H. Conrad, Rene J.

LeClaire, Gerard P. O’Reilly, and Huseyin Uzunalioglu. Critical national infrastructure reliability modeling and analysis. Bell Labs Technical Journal, 11(3):57–71, Autumn 2006. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Cai:2006:IMS

[CLW^{+}06] Yigang Cai, Xian Yang Li, Yile Wang, John B. Reid, and Peng Wang. IP multimedia subsystem online session charging call control. Bell Labs Technical Journal, 10(4):117–132, Winter 2006. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Chan:2005:TEA

[CMC^{+}05] Ho Bun Chan, Zsolt Marcet, Dustin Carr, John Eric Bower, Ray Cirelli, Ed Ferry, Fred P. Klemens, John F. Miner, Chien-Shing Pai, and J. Ashley Taylor. Transmission enhancement in an array of subwavelength slits in aluminum due to surface plasmon resonances. Bell Labs Technical Journal, 10 (3):143–150, Autumn 2005. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Calin:2006:MAS

[CMCP06] Doru Calin, Andrew R. McGee, Uma Chandrashekhar, and Ramjee Prasad. MAG-NET: An approach for secure personal networking in beyond 3g wireless networks. Bell Labs Technical Journal, 11(1):79–98, Spring 2006. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Calin:2008:AJT

[CMMZ08] Doru Calin, Andrew K. Mackay, Thomas B. Morawski, and Hardy Zhang. An approach for just-in-time radio access network capacity planning in CDMA networks. Bell Labs Technical Journal, 12 (4):67–82, Winter 2008. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Chuah:2002:QST

[CMPW02] Mooi Choo Chuah, Kamesh Medepalli, Se-Yong Park, and Jiansong Wang. Quality of service in thirdgeneration IP-based radio access networks. Bell Labs Technical Journal, 7(2):67– 89, Summer 2002. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Cowsar:2007:GRP

[Cow07] Lawrence Cowsar. A global research presence. Bell Labs Technical Journal, 12(2):1– 2, Summer 2007. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Chan:2004:RSM

[CP04] Chun K. Chan and Himanshu Pant. Reliability and security modeling in upgrading wireless backbone networks. Bell Labs Technical Journal, 8(4): 39–53, Winter 2004. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Clark:2007:CSM

[CP07] Andrew C. Clark and Karen E.

Plonty. Common security management for CDMA2000 1x and 1x EV-do networks. Bell Labs Technical Journal, 12(3):187– 203, Autumn 2007. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Claussen:2009:SOF

[CPH09] Holger Claussen, Florian Pivit, and Lester T. W. Ho. Self-optimization of femtocell coverage to minimize the increase in core network mobility signalling. Bell Labs Technical Journal, 14 (2):155–183, Summer 2009. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Chambers:2004:IWP

[CR04] Michael D. Chambers and Douglas H. Riley. Implementing wireless priority service for CDMA networks. Bell Labs Technical Journal, 9(2): 23–36, Summer 2004. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Chandrashekhar:2004:O

[CRT04] Uma Chandrashekhar, Steven H. Richman, and Rati C. Thanawala. Overview. Bell Labs Technical Journal, 8 (4):1–5, Winter 2004. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Chandrashekhar:2002:AEU

[CRVX02] Uma Chandrashekhar, Steven H. Richman, S. Rao Vasireddy, and Chen Xie. Applications exchange: A unified approach to enable services and applications for end users and service providers. Bell Labs Technical Journal, 7(1): 47–62, Spring 2002. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Coss:2004:NPD

[CS04] Michael Coss and Ron Sharp.

The network processor decision. Bell Labs Technical Journal, 9(1):177– 189, Spring 2004. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Coyle:2008:SNC

[CV08] Cheryl L. Coyle and Heather Vaughn. Social networking: Communication revolution or evolution? Bell Labs Technical Journal, 13(2):13– 17, Summer 2008. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Danielsen:2008:AS

[Dan08] Peter J. Danielsen. Absence services. Bell Labs Technical Journal, 13(2):25– 30, Summer 2008. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Das:2000:NEO

[Das00] Saswato R. Das. Noble endeavors: An overview of Nobel Prize-winning research at Bell Labs. Bell Labs Technical Journal, 5(1):95– 106, Spring 2000. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Duran:2009:UFP

[DC09] Alfonso Fernandez Duran and Gloria García Carrasco. UMTS femtocell performance in massive deployments: Capacity and GoS implications. Bell Labs Technical Journal, 14(2):185–202, Summer 2009. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Dianda:2000:PSP

[DCHY00] Janet R. Dianda, Raymond O. Colbert, Philippe J. L. Herve, and Tongzeng Yang. Programmable service platforms for converged voice/data services. Bell Labs Technical Journal, 5(3): 43–58, Autumn 2000. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Dianda:2001:SAT

[DDHS01] Janet R. Dianda, S. Carolyn Darity, Bin-Wen Ho, and Kim J. Scott. Service authoring for thirdparty programmable, servicemediation-enabled feature servers in the multiservice core. Bell Labs Technical Journal, 6(1):192–210, Spring 2001. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Degrande:2008:PIA

[DDL08] Natalie Degrande, Danny De Vleeschauwer, and Koen Laevens. Protecting IPTV against packet loss: Techniques and trade-offs. Bell Labs Technical Journal, 13 (1):35–51, Spring 2008. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

DeVleeschauwer:2009:SNG

[De 09] Danny De Vleeschauwer.

Supporting next-generation multimedia services over a broadband access network. Bell Labs Technical Journal, 14(1):73–88, Spring 2009. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Dedolph:2003:NMA

[Ded03] F. Michael Dedolph. The neglected management activity: Software risk management. Bell Labs Technical Journal, 8(3):91–95, Autumn 2003. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

DeVito:2005:FSS

[DEK^{+}05] Nicholas M. DeVito, Richard T.

Emery, Kristin F. Kocan, William D. Roome, and Byron J. Williams. Functionality and structure of the service broker in advanced service architectures. Bell Labs Technical Journal, 10 (1):17–30, Spring 2005. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Doshi:2003:VNA

[DER^{+}03] Bharat T. Doshi, Dominik Eggenschwiler, Aswath Rao, Behrokh Samadi, Y. T. Wang, and James Wolfson. VoIP network architectures and QoS strategy. Bell Labs Technical Journal, 7(4): 41–59, Winter 2003. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Derman:2004:MLQ

[Der04] Emanuel Derman. My life as a quant: reflections on physics and finance. John Wiley, New York, NY, USA, 2004. ISBN 0-471-39420-3 (hardcover). xi + 292 pp. LCCN HG4621 .D47 2004.

URL http://www.loc.gov/
catdir/description/wiley042/
2004007664.html;   http:
//www.loc.gov/catdir/toc/
ecip0416/2004007664.html;
 http://www.loc.gov/catdir/
enhancements/fy0616/2004007664-
b.html.

Derkits:2006:BMF

[DFR06] Gustav E. Derkits, John P. Franey, and William D. Reents. Biased mixed flowing gas testing for world class reliability. Bell Labs Technical Journal, 11(3):105– 120, Autumn 2006. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Drabeck:2005:NOT

[DFS^{+}05] Lawrence M. Drabeck, Michael J.

Flanagan, Jayanthi Srinivasan, William M. MacDonald, Georg Hampel, and Alvaro Diaz. Network optimization trials of a vendorindependent methodology using the Ocelot(R) tool. Bell Labs Technical Journal, 9(4): 49–66, Winter 2005. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Das:2002:EUT

[DGH^{+}02] Arnab Das, Nandu Gopalakrishnan, Teck Hu, Farooq Khan, Ashok Rudrapatna, Ashwin Sampath, Hsuan-Jung Su, Said Tatesh, and Wenfeng Zhang. Evolution of UMTS toward high-speed downlink packet access. Bell Labs Technical Journal, 7(3): 47–68, Autumn 2002. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Dianda:2002:SIP

[DGJ02] Janet R. Dianda, Vijay K.

Gurbani, and Mark H. Jones. Session initiation protocol services architecture. Bell Labs Technical Journal, 7 (1):3–23, Spring 2002. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Donahue:2007:PEC

[DHJ^{+}07] Chris W. Donahue, G. Jay Hruska, James B. Jones, Joseph J. Lichter, Joseph E. Seitz, Ranga P. Tota, and A. Larry Wierzbicki. Packet evolution of the CDMA2000-1x CDMA system. Bell Labs Technical Journal, 11 (4):75–91, Winter 2007. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Dianda:2000:RCC

[DHK00] Janet R. Dianda, Bin-Wen Ho, and Kristin F. Kocan. Reducing complexity for converged voice/data networks and services architecture. Bell Labs Technical Journal, 5(2):55–71, Summer 2000. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Denison:2003:IMM

[DHM^{+}03] Brett A. Denison, Craig W. Hoopmann, David P. Mongeau, Eric M. Shepherd, and Paul Wu. Intelligent maintenance and management of Service Intelligent^{T M} network architectures. Bell Labs Technical Journal, 7(4): 171–185, Winter 2003. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Deb:2007:AHC

[DJM^{+}07] Supratim Deb, Ankur Jain, Anirban Majumder, K. V. M. Naidu, Jeyashankher Ramamirtham, Rajeev Rastogi, and Anand Srinivasan. Augmenting handset capacity through virtual storage. Bell Labs Technical Journal, 12(2):87–96, Summer 2007. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Dutta:2007:VLC

[DJR07] Partha Dutta, Sharad Jaiswal, and Rajeev Rastogi. VillageNet: A low-cost, IEEE 802.11-based mesh network for connecting rural areas. Bell Labs Technical Journal, 12(2):119–131, Summer 2007. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Dombkowski:2004:FTM

[DK04] Kevin E. Dombkowski and Kristin F. Kocan. FPGA technology to minimize extended life-cycle development. Bell Labs Technical Journal, 9(1):191–195, Spring 2004. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Dowden:2000:FNP

[DKK00] Douglas C. Dowden, Kristin F.

Kocan, and Jack Kozik. The future of network-provided communications services. Bell Labs Technical Journal, 5(3):3–11, Autumn 2000. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Davis:2001:SSF

[DKLS01] R. Drew Davis, Krishnan Kumaran, Gang Liu, and Iraj Saniee. SPIDER: A simple and flexible tool for design and provisioning of protected lightpaths in optical networks. Bell Labs Technical Journal, 6(1):82– 97, Spring 2001. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Das:2006:SAH

[DKR^{+}06] Suman Das, Thierry Klein, Ajay Rajkumar, Sampath Rangarajan, Michael Turner, and Harish Viswanathan. System aspects and handover management for IEEE 802.16e. Bell Labs Technical Journal, 11(1):123– 142, Spring 2006. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Davolos:2002:FPC

[DLM^{+}02] Christopher J. Davolos, Subhasis Laha, Stinson S. Mathai, Albert J. Sawyer, and Jin Wang. The Flexent(R) packet core access platform for CDMA2000 packet data. Bell Labs Technical Journal, 7(2):125– 142, Summer 2002. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Das:2007:EDR

[DLM^{+}07] Suman Das, Shupeng Li, Pantelis Monogioudis, Shirish Nagaraj, Sudhir Ramakrishna, Ashok N. Rudrapatna, Sivarama Venkatesan, Subramanian Vasudevan, Harish Viswanathan, and Jialin Zou. EV-do revision C: Evolution of the CDMA2000 data optimized system to higher spectral efficiencies and enhanced services. Bell Labs Technical Journal, 11 (4):5–24, Winter 2007. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Dunn:2003:EPL

[DM03] Jeffrey H. Dunn and Cynthia E. Martin. The effect of physical and link layers on IP QoS. Bell Labs Technical Journal, 8(2):49– 63, Summer 2003. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Davison:2000:UAE

[DMO00] Joseph W. Davison, Dennis M. Mancl, and William F. Opdyke. Understanding and addressing the essential costs of evolving systems. Bell Labs Technical Journal, 5(2):

44–54, Summer 2000. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Doshi:2001:FWA

[DNPQ01] Bharat T. Doshi, Ramesh Nagarajan, G. N. Srinivasa Prasanna, and M. Akber Qureshi. Future WAN architecture driven by services, traffic volume, and technology trends. Bell Labs Technical Journal, 6(1):13– 32, Spring 2001. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Doshi:2003:SIA

[Dos03] Bharat T. Doshi. Service Intelligent^{T M} architecture. Bell Labs Technical Journal, 7(4):1–2, Winter 2003. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Donadio:2009:SOT

[DPB09] Pasquale Donadio, Andrea Paparella, and Bela Berde. Service-oriented technology for TMN-based network management services. Bell Labs Technical Journal, 14 (1):161–172, Spring 2009. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Dinu:2005:NQD

[DRC^{+}05] Mihaela Dinu, Ronen Rapaport, Gang Chen, Howard R.

Stuart, and Randy Giles. Nanophotonics — quantum dots, photonic crystals, and optical silicon circuits: An excursion into the optical behavior of very small things. Bell Labs Technical Journal, 10(3):215– 234, Autumn 2005. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

DeTura:2004:DRP

[DRNY04] Nick De Tura, Susan Marie Reilly, Srilatha Narasimhan, and Zhenhua Jack Yin. Disaster recovery preparedness through continuous process optimization. Bell Labs Technical Journal, 9(2):147– 162, Summer 2004. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Duan:2009:HPI

[DSA^{+}09] Guang-Hua Duan, Alexandre Shen, Akram Akrout, Frédéric Van Dijk, Francois Lelarge, Frédéric Pommereau, Odile LeGouezigou, Jean-Guy Provost, Harry Gariah, Fabrice Blache, Franck Mallecot, Kamel Merghem, Anthony Martinez, and Abderrahim Ramdane. High performance InPbased quantum dash semiconductor mode-locked lasers for optical communications. Bell Labs Technical Journal, 14(3):63–84, Autumn 2009.

CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Doherty:2008:NGN

[DSW08] Deirdre H. Doherty, Raymond A. Sackett, and Paul Wu. Next-generation networks application traffic modeling. Bell Labs Technical Journal, 13(2):209– 222, Summer 2008. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Epstein:2000:HNM

[EACM00] Harvey I. Epstein, Abhaya Asthana, Stephen A. Corum, and Leen Mak. Hybrid network management. Bell Labs Technical Journal, 5(4): 63–79, Winter 2000. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Epstein:2001:MON

[ECGL01] Harvey Epstein, Steve Corum, Harald Graber, and Patrice N. Lamy. Managing optical networks. Bell Labs Technical Journal, 6(1): 50–66, Spring 2001. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Ejzak:2004:NOC

[EFH04] Richard P. Ejzak, Cynthia K. Florkey, and Richard W. Hemmeter. Network overload and congestion: A comparison of ISUP and SIP. Bell Labs Technical Journal, 9(3): 173–182, Autumn 2004. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Eggert:2000:FTE

[EJV00] Marvin A. Eggert, Richard A.

Johnston, and George W. Vaughan. The Flexent^{T M} Element Management System — Using web and object technologies to manage a wireless network. Bell Labs Technical Journal, 5(4): 113–125, Winter 2000. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Erfani:2000:MPS

[ELM00] Shervin Erfani, Victor B.

Lawrence, and Manu Malek. The management paradigm shift: Challenges from element management to service management. Bell Labs Technical Journal, 5(4):3– 20, Winter 2000. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Ejzak:2002:FIC

[ELOW02] Richard P. Ejzak, Harold A.

Lassers, Scott D. Olmstead, and Randall J. Wilson. Flexent(R) IMS — the convergence of circuit and packet core networks. Bell Labs Technical Journal, 7(2):105– 124, Summer 2002. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Erman:2008:ARI

[EM08] Bilgehan Erman and Elissa P. Matthews. Analysis and realization of IPTV service quality. Bell Labs Technical Journal, 12(4):195– 212, Winter 2008. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Elko:2005:CMM

[EPL^{+}05] Gary W. Elko, Flavio Pardo, Daniel López, David Bishop, and Peter Gammel. Capacitive MEMS microphones. Bell Labs Technical Journal, 10(3):187–198, Autumn 2005. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

El-Sayed:2006:CTN

[ESHKW06] Mohamed L. El-Sayed, Ying Hu, Samrat Kulkarni, and Newman Wilson. Comparison of transport network technologies for IPTV distribution. Bell Labs Technical Journal, 11(2):215– 240, Summer 2006. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Fishman:2006:RON

[FCG^{+}06] Daniel A. Fishman, Diego L. Correa, Edward H. Goode, Terry L. Downs, Augustine Y. Ho, Alan Hale, Peter Hofmann, Bert Basch, and Steven Gringeri. The rollout of optical networking: LambdaXtreme(R) national network deployment. Bell Labs Technical Journal, 11(2):55–63, Summer 2006. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Flanagan:2007:WNA

[FDC^{+}07] Michael J. Flanagan, Lawrence M.

Drabeck, Larry A. Cohen, Alvaro H. Diaz, and Jayanthi Srinivasan. Wireless network analysis using per call measurement data. Bell Labs Technical Journal, 11 (4):307–313, Winter 2007. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Fischer:2002:RMS

[FEK02] Georg Fischer, Wolfgang Eckl, and Gerhard Kaminski. RF-MEMS and SiC/GaN as enabling technologies for a reconfigurable multi-band/ multi-standard radio. Bell Labs Technical Journal, 7(3): 169–189, Autumn 2002. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Fisher:2007:EMGa

[FFG^{+}07a] Eric S. Fisher, Steven Fortune, Martin K. Gladstein, Suresh Goyal, William B.

Lyons, James H. Mosher Jr., and Gordon Wilfong. Economic modeling of global test strategy I: Mathematical models. Bell Labs Technical Journal, 12(1):161– 173, Spring 2007. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Fisher:2007:EMGb

[FFG^{+}07b] Eric S. Fisher, Steven Fortune, Martin K. Gladstein, Suresh Goyal, William B. Lyons, James H. Mosher Jr., and Gordon Wilfong. Economic modeling of global test strategy II: Software system and examples. Bell Labs Technical Journal, 12 (1):175–186, Spring 2007. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Foschini:2005:PLD

[FHM^{+}05] Gerard J. Foschini, Howard C. Huang, Sape J. Mullender, Sivarama Venkatesan, and Harish Viswanathan. Physical-layer design for next-generation cellular wireless systems. Bell Labs Technical Journal, 10(2):157– 172, Summer 2005. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Fithen:2004:FMV

[FHOS04] William L. Fithen, Shawn V. Hernan, Paul F. O’Rourke, and David A. Shinberg. Formal modeling of vulnerability. Bell Labs Technical Journal, 8(4):173–186, Winter 2004. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Fischer:2001:ASU

[FHW01] Christine A. Fischer, Keli Han, and Zhibi Wang. Accounting solutions in the UMTS core network. Bell Labs Technical Journal, 6(2): 65–73, Summer 2001. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Fischer:2007:NGB

[Fis07]  Georg Fischer.  Next-

generation base station radio frequency architecture. Bell Labs Technical Journal, 12 (2):3–18, Summer 2007. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Farhang:2004:PBQ

[FK04] Bijan Farhang and Roy Kopeikin. Policy-based quality of service in 3G networks. Bell Labs Technical Journal, 9(1):31–40, Spring 2004. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Friedlander:2006:YTR

[FLM^{+}06] Jay Friedlander, Karthic Loganathan, Ransom Murphy, Ramesh V. Pattabhiraman, and Kumar V. Vemuri. Are you there? Reflections on presence server architectures. Bell Labs Technical Journal, 10(4):77–82, Winter 2006. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Floyd:2006:MAS

[Flo06] David Floyd. Mobile application security system (MASS). Bell Labs Technical Journal, 11(3):191– 198, Autumn 2006. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Fogliata:2008:IRN

[FM08] Paolo Fogliata and Marco Torquato Mussini. Intelligence-ready network infrastructure: An ecosystem to control thirdparty intelligence distribution close to nomadic users. Bell Labs Technical Journal, 13(2):105–116, Summer 2008. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Freund:2006:CAS

[Fre06] Jack F. Freund. Catastrophe alerting system: Cellular systems as scale-free networks. Bell Labs Technical Journal, 11(1):191–196, Spring 2006. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Franey:2006:SIT

[FS06] John P. Franey and Dawn-Marie Sutton. Static intercept technology: A new packaging platform for corrosion and ESD protection. Bell Labs Technical Journal, 11 (3):137–146, Autumn 2006. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Fishman:2006:LTS

[FTV06] Daniel A. Fishman, William A.

Thompson, and Lee Vallone. LambdaXtreme(R) transport system: R&D of a high capacity system for low cost, ultra long haul DWDM transport. Bell Labs Technical Journal, 11(2):27– 53, Summer 2006. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Feng:2009:DSD

[FYJB09] Yang Feng, Song Yang, Qiu Jigang, and Xu Binyang. Distributed spatial division multiple access technique for spectrum sharing systems. Bell Labs Technical Journal, 13(4):119–128, Winter 2009. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Frey:2004:RSM

[FZ04] Alan E. Frey and Guy J. Zenner. The role of SIP in the migration of service provider networks to VoIP. Bell Labs Technical Journal, 9(3):199– 216, Autumn 2004. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Fortune:2009:EST

[FZ09] Steven Fortune and Martin Zirngibl. Enabling science and technology. Bell Labs Technical Journal, 14 (3):1–5, Autumn 2009. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Gundanna:2000:DWB

[GA00] Veda Gundanna and Leo J. Agrillo. DEFINITY wireless business system — Health care application: nurse call system integration. Bell Labs Technical Journal, 5(4):171– 184, Winter 2000. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Goers:2000:IMS

[GB00] William C. Goers and Michael R. Brenner. Implementing a management system architecture framework. Bell Labs Technical Journal, 5(4):31–43, Winter 2000. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Gurbani:2001:SRP

[GCK01] Vijay K. Gurbani, Tsun-Chieh Chiang, and Suresh Kumar. SIP: A routing protocol. Bell Labs Technical Journal, 6(2):136– 152, Summer 2001. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Gupta:2007:BSP

[GCSB07] Ashok K. Gupta, Uma Chandrashekhar, Suhasini V. Sabnis, and Frank A. Bastry. Building secure products and solutions. Bell Labs Technical Journal, 12(3):21– 38, Autumn 2007. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Glass:2000:AFO

[GDS^{+}00] Alastair M. Glass, David J.

DiGiovanni, Thomas A. Strasser, Andrew J. Stentz, Richart E. Slusher, Alice E. White, A. Refik Kortan, and Benjamin J. Eggleton. Advances in fiber optics. Bell Labs Technical Journal, 5(1): 168–187, Spring 2000. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Gehani:2003:BLL

[Geh03] Narain Gehani. Bell Labs:

life in the crown jewel. Silicon Press, 25 Beverly Road, Summit, NJ 07901, USA, 2003. ISBN 0-929306-27-9. xi + 258 pp. LCCN TK415.B45 G44 2003. URL http:/

                                        /www.loc.gov/catdir/toc/
                                        fy043/2002012730.html.

Goldey:2004:URS

[GEOT04] Charles L. Goldey, Lisa M. Enaire, Donald Olson, and Stacy L. Trottier. Using remanufactured switching equipment to reduce restoration time in case of a disaster. Bell Labs Technical Journal, 9(2):181–188, Summer 2004. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Gammel:2005:RMN

[GFB05] Peter Gammel, Georg Fischer, and Jérémie Bouchaud. RF MEMS and NEMS technology, devices, and applications. Bell Labs Technical Journal, 10(3):29–59, Autumn 2005. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Ghosale:2003:IHP

[GGHH^{+}03] Nandan Ghosale, Janel A. Green, Jorge Hernandez-Herrero, Gang G. Huang, and Prashant S. Parikh. On implementing a highperformance open API with Java. Bell Labs Technical Journal, 7(4):123–138, Winter 2003. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Graber:2006:OPT

[GHVH06] Harald Graber, Enrique Hernandez-Valencia, and Bernd Heinmueller. Optical/packet transport convergence. Bell Labs Technical Journal, 11(1):143– 158, Spring 2006. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Gehrke:2008:QEC

[GHW08] Gerhard Gehrke, Ralf Hocke, and Georg Wenzel. QoE estimation of compound services in significance-aware packet networks. Bell Labs Technical Journal, 13(3):95– 109, Autumn 2008. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Gurbani:2004:CSO

[GJ04a] Vijay K. Gurbani and Rajnish Jain. Contemplating some open challenges in SIP. Bell Labs Technical Journal, 9(3):255–269, Autumn 2004. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Gurbani:2004:TPC

[GJ04b] Vijay K. Gurbani and Rajnish Jain. Transport protocol considerations for session initiation protocol networks. Bell Labs Technical Journal, 9(1):83–97, Spring 2004. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Gopalakrishnan:2002:ERL

[GKP^{+}02] Nandu Gopalakrishnan, Srinivas R. Kadaba, Eshwar Pittampalli, Sudhir Ramakrishna, Ashok N. Rudrapatna, Ganapathy S. Sundaram, Subramanian Vasudevan, and Yunsong Yang. Evolution of the reverse link of CDMA-based systems to support high-speed data. Bell Labs Technical Journal, 7(3): 69–82, Autumn 2002. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Gurbani:2003:SIP

[GL03] Vijay K. Gurbani and Kim Q. Liu. Session initiation protocol: Service residency and resiliency. Bell Labs Technical Journal, 8(1):83– 94, Spring 2003. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Gaglianello:2006:ISS

[GL06] Robert Gaglianello and Dong Liu. IMS shared streaming video. Bell Labs Technical Journal, 10(4):71–75, Winter 2006. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Gordon:2002:SPM

[GLM^{+}02] Minerva E. Gordon, Carolyn C. Leech, Greg A. Martin, Rick R. McTee, Reinhard Metz, and Douglas A. Reneker. Success in processor multiuse. Bell Labs Technical Journal, 7(1):215– 226, Spring 2002. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Goyal:2006:ITP

[GM06] Suresh Goyal and James H.

Mosher Jr. An improved test process model for cost reduction. Bell Labs Technical Journal, 11(1):173– 190, Spring 2006. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Gurbani:2007:EAB

[GM07] Vijay K. Gurbani and Andrew R. McGee. An early application of the Bell Labs Security framework to analyze vulnerabilities in the Internet telephony domain. Bell Labs Technical Journal, 12 (3):7–19, Autumn 2007. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Grech:2000:DSS

[GMS^{+}00] Michel L. F. Grech, Robert D.

McKinney, Sharad Sharma, John J. Stanaway Jr., Douglas W. Varney, and Kumar V. Vemuri. Delivering seamless services in open networks using intelligent service mediation. Bell Labs Technical Journal, 5(3):186– 202, Autumn 2000. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Gayde:2007:ANS

[GMW07] Ruth S. Gayde, Stinson S. Mathai, and Kevin Wang. Aspects of network security for VoIP solutions using IMS core network and Wi-Fi access. Bell Labs Technical Journal, 12(3):167– 185, Autumn 2007. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Goldstein:2005:SSR

[Gol05] Robert E. Goldstein. Services to support regulatory action and requirements. Bell Labs Technical Journal, 9(4): 31–34, Winter 2005. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Golash:2006:REN

[Gol06] Mukul Golash. Reliability in Ethernet networks: A survey of various approaches. Bell Labs Technical Journal, 11 (3):161–171, Autumn 2006. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Graber:2003:MSS

[GR03] Harald Graber and Humberto J. La Roche. Multiservice switches and the Service Intelligent^{T M} optical architecture for SONET/SDH metro networks. Bell Labs Technical Journal, 8(2):111– 127, Summer 2003. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Gollamudi:2009:BMT

[GSCF09] Sridhar Gollamudi, Robert Soni, and Monica Casado-Fernandez. Broadcast and multicast transmission over HSDPA. Bell Labs Technical Journal, 13(4):195– 205, Winter 2009. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Greenwald:2007:UPN

[GT07] Lloyd G. Greenwald and Tavaris J. Thomas. Understanding and preventing network device fingerprinting. Bell Labs Technical Journal, 12(3):149– 166, Autumn 2007. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Heng:2005:BGC

[HAH^{+}05] Jiunn Benjamin Heng, Aleksei Aksimentiev, Chuen Ho, Valentin Dimitrov, Thomas W. Sorsch, John F. Miner, William M. Mansfield, Klaus Schulten, and Gregory Timp. Beyond the gene chip. Bell Labs Technical Journal, 10 (3):5–22, Autumn 2005. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Hao:2008:TUS

[Hao08] Qingwang Michael Hao. Toward a unified service delivery process for nextgeneration services. Bell Labs Technical Journal, 12 (4):5–20, Winter 2008. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Hartley:2005:DES

[Har05] Karen L. Hartley. Defining effective service level agreements for network operation and maintenance. Bell Labs Technical Journal, 9(4):139– 143, Winter 2005. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Ho:2009:OAI

[HB09] Tin Kam Ho and Thomas Bengtsson. Opportunities for automated inference with data in tomorrow’s communication networks and services. Bell Labs Technical Journal, 14(3):193– 206, Autumn 2009. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Hobby:2009:DOF

[HC09] John D. Hobby and Holger Claussen. Deployment options for femtocells and their impact on existing macrocellular networks. Bell Labs Technical Journal, 13(4): 145–160, Winter 2009. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Hilt:2004:AIS

[HH04] Volker Hilt and Markus Hofmann. Approaches to implementing services in SIP networks. Bell Labs Technical Journal, 9(3):39–44, Autumn 2004. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Hari:2005:IMG

[HHH05] Adiseshu Hari, Volker Hilt, and Markus Hofmann. Intelligent media gateway selection in a VoIP network. Bell Labs Technical Journal, 10 (1):47–57, Spring 2005. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Hari:2006:IMG

[HHH^{+}06] Adiseshu Hari, Volker Hilt, Markus Hofmann, Debasis Mitra, Iraj Saniee, Anwar Walid, and Indra Widjaja. Intelligent media gateway selection in carrier-grade VoIP networks. Bell Labs Technical Journal, 10(4):133– 150, Winter 2006. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Hernandez-Herrero:2007:NMP

[HHS07] Jorge Hernandez-Herrero and Jon A. Solworth. The need for a multi-perspective approach to solve the DDos problem. Bell Labs Technical Journal, 12(3):121– 130, Autumn 2007. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Hong:2008:ECM

[HHS08] Se Gi Hong, Volker Hilt, and Henning Schulzrinne. Evaluation of control message overhead of a DHT-based P2P system. Bell Labs Technical Journal, 13(3):79– 86, Autumn 2008. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Hermsmeyer:2007:EAC

[HHVST07] Christian Hermsmeyer, Enrique Hernandez-Valencia, Dieter Stoll, and Oliver Tamm. Ethernet aggregation and core network models for efficient and reliable IPTV services. Bell Labs Technical Journal, 12 (1):57–76, Spring 2007. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Houck:2004:WSP

[HJM^{+}04] David J. Houck, Jean-Philippe Joseph, Frank Magee, Amit Mukhopadhyay, and Benjamin Tang. Wireless service provider backbone network evolution with packet technologies. Bell Labs Technical Journal, 9(1): 41–64, Spring 2004. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Hajjaoui:2006:NAU

[HJvO^{+}06] Abdelkader Hajjaoui, Hanno Juechter, Gijs van Ooijen, Govinda Rajan, Walter Rothkegel, and Willem van Willigenburg. Network analysis using the ASON/ GMPLS emulator eGEM. Bell Labs Technical Journal, 11(2):157–169, Summer 2006. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Hanmer:2004:DAP

[HK04] Robert S. Hanmer and Kristin F. Kocan. Documenting architectures with patterns. Bell Labs Technical Journal, 9(1):143– 163, Spring 2004. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Hinton:2005:APF

[HKH05] Danielle Hinton, Thierry E.

Klein, and Mark Haner. An architectural proposal for future wireless emergency response networks with broadband services. Bell Labs Technical Journal, 10(2): 121–138, Summer 2005. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Houck:2004:NSM

[HKO^{+}04] David J. Houck, Eunyoung Kim, Gerard P. O’Reilly, David D. Picklesimer, and Huseyin Uzunalioglu. A network survivability model for critical national infrastructures. Bell Labs Technical Journal, 8(4):153–172, Winter 2004. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Hull:2004:PES

[HKQ^{+}04] Richard B. Hull, Bharat B. Kumar, S. Shehryar Qutub, Musa R. Unmehopa, and Douglas W. Varney. Policy enabling the services layer. Bell Labs Technical Journal, 9(1):5–18, Spring 2004. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Houck:2003:MBA

[HKUW03] David J. Houck, Eunyoung Kim, Huseyin Uzunalioglu, and Larry A. Wehr. A measurement-based admission control algorithm for VoIP. Bell Labs Technical Journal, 8(2):97–110, Summer 2003. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Hanmer:2003:BPP

[HL03] Robert S. Hanmer and John P. Letourneau. A best practice for performance engineering. Bell Labs Technical Journal, 8(3):75– 89, Autumn 2003. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Haworth:2006:RNM

[HL06] James W. Haworth and John A. Liffrig. Remote network management for optical networks. Bell Labs Technical Journal, 11(2): 191–201, Summer 2006. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Huang:2006:MLB

[HLS^{+}06] Dawei Huang, Fang Liu, Xiangdong Shi, Guangxin Yang, Ludi Zheng, and Zhiyu Zhou. MapWeb: A locationbased converged communications platform. Bell Labs Technical Journal, 11(1): 159–171, Spring 2006. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Hofmann:2009:NAC

[HLS09] Stefan Hofmann, Mohamed Louizi, and Dieter Stoll. A novel approach to counter denial of service attacks against transport network resources. Bell Labs Technical Journal, 14(1):219–242, Spring 2009. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Hilt:2004:FSS

[HMH04] Volker Hilt, Allison Mankin, and Markus Hofmann. A framework for SIP session policies. Bell Labs Technical Journal, 9(3):45–56, Autumn 2004. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Hanmer:2007:CRS

[HMM07] Robert S. Hanmer, Donald T. McBride, and Veena B. Mendiratta. Comparing reliability and security: Concepts, requirements, and techniques. Bell Labs Technical Journal, 12(3):65–78, Autumn 2007. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Hlavacek:2004:VSP

[HMR04] Donna M. Hlavacek, Karl A. Madsen, and Robert M. Reimer. A vendor and service provider partnership for preparing to manage disaster recovery. Bell Labs Technical Journal, 9(2):173– 180, Summer 2004. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Hu:2007:ISE

[HMS^{+}07] Ying Hu, Amit Mukhopadhyay, Zulfiquar Sayeed, Carlos Urrutia-Valdés, and Mohamed El-Sayed. IMS service enhancement layer: A quantitative value proposition. Bell Labs Technical Journal, 12(1):95–114, Spring 2007. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Holzmann:2000:SVB

[Hol00] Gerard J. Holzmann. Software verification at Bell Labs: One line of development. Bell Labs Technical Journal, 5(1):35–45, Spring 2000. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Hehmann:2008:NMC

[HP08] Jörg Hehmann and Thomas Pfeiffer. New monitoring concepts for optical access networks. Bell Labs Technical Journal, 13(1):183– 198, Spring 2008. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Heron:2008:TIA

[HPvV^{+}08] Ronald W. Heron, Thomas Pfeiffer, Doutje T. van Veen, Joe Smith, and Sanjay S. Patel. Technology innovations and architecture solutions for the next-generation optical access network. Bell Labs Technical Journal, 13 (1):163–181, Spring 2008. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Hoekstra:2002:INM

[HRB^{+}02] Geert Jan Hoekstra, Willem A. Romijn, Harold C. H. Balemans, Abdelkader Hajjaoui, and Gijs G. van Ooijen. An integrated network management solution for multitechnology domain networks. Bell Labs Technical Journal, 7(1):115–120, Spring 2002. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Holzmann:2000:ASF

[HS00] Gerard J. Holzmann and Margaret H. Smith. Automating software feature verification. Bell Labs Technical Journal, 5(2):72– 87, Summer 2000. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Ho:2005:SPE

[HSF^{+}05] Tin Kam Ho, Todd Salamon, Roland W. Freund, Christopher A. White, Bruce K. Hillyer, Lawrence C. Cowsar, Carl J. Nuzman, and Daniel C. Kilper. Simulation of power evolution and control dynamics in optical transport systems. Bell Labs Technical Journal, 10(1):119– 137, Spring 2005. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Hernon:2009:TME

[HSK^{+}09] Domhnaill Hernon, Todd Salamon, Roger Kempers, Shankar Krishnan, Alan Lyons, Marc Hodes, Paul Kolodner, John Mullins, and Liam McGarry. Thermal management: Enabling enhanced functionality and reduced carbon footprint. Bell Labs Technical Journal, 14 (3):7–19, Autumn 2009. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Ho:2003:AES

[HSP03] Lester T. W. Ho, Louis G.

Samuel, and Jonathan M. Pitts. Applying emergent self-organizing behavior for the coordination of 4G networks using complexity metrics. Bell Labs Technical Journal, 8(1):5– 25, Spring 2003. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Hermsmeyer:2009:TPP

[HSS^{+}09] Christian Hermsmeyer, Haoyu Song, Ralph Schlenk, Riccardo Gemelli, and Stephan Bunse. Towards 100G packet processing: Challenges and technologies. Bell Labs Technical Journal, 14(2):57– 79, Summer 2009. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Hofmann:2008:O

[HT08] Markus Hofmann and Marco Tomsu. Overview. Bell Labs Technical Journal, 13 (3):1–4, Autumn 2008. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Heck:2006:BIM

[HV06] John Heck and Greg Vaudreuil. Blended IMS messaging applications. Bell Labs Technical Journal, 10 (4):39–52, Winter 2006. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Hernandez-Valencia:2003:MVP

[HVKL03] Enrique J. Hernandez-Valencia, Pramod Koppol, and Wing Cheong Lau. Managed virtual private LAN services. Bell Labs Technical Journal, 7(4): 61–76, Winter 2003. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Isukapalli:2002:GRP

[IAM02] Ramana Isukapalli, Triantafyllos Alexiou, and Kazutaka Murakami. Global roaming and personal mobility with COPS architecture in SuperDHLR. Bell Labs Technical Journal, 7(2): 3–18, Summer 2002. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Islam:2003:CGD

[IB03] Munirul Islam and Eugene P.

Bordelon. CCMS: A geographically distributed concurrent change and configuration management system. Bell Labs Technical Journal, 8(3):111–133, Autumn 2003. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Isukapalli:2008:AIC

[IBPF08] Ramana Isukapalli, Steven Benno, Candace Park, and Peretz M. Feder. Advanced IMS client supporting secure signaling. Bell Labs Technical Journal, 12(4):49– 65, Winter 2008. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Ionescu-Graff:2005:QVP

[IGNC^{+}05] Alina M. Ionescu-Graff, Cheryl F. Newman, Chi-Hung Kelvin Chu, Bhadrayu J. Trivedi, Benjamin Tang, and Alexander Y. Zhu. Quantifying the value proposition of advanced fault management systems in MPLS core networks. Bell Labs Technical Journal, 10(1): 157–167, Spring 2005. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Jain:2002:IMN

[Jai02] Anant Kumar Jain. Intelligent multiservice networks. Bell Labs Technical Journal, 7(1):81–97, Spring 2002. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Jones:2009:OUI

[JBR^{+}09] Christopher D. W. Jones, Cristian A. Bolle, Roland Ryf, Maria Elina Simon, Flavio Pardo, Nagesh Basavanhally, and Arthur P. Ramirez. Opportunities in uncooled infrared imaging: A MEMS perspective. Bell Labs Technical Journal, 14(3):85– 98, Autumn 2009. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Jiao:2007:PEE

[JCL07] Wenhua Jiao, Jianfeng Chen, and Fang Liu. Provisioning end-to-end QoS under IMS over a WiMAX architecture. Bell Labs Technical Journal, 12(1):115–121, Spring 2007. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Jayadevan:2004:ASA

[JGA04] Ganesh Jayadevan, Vijay K.

Gurbani, and Robert M. Arlein. Adapting SIP for application server requirements in 3G networks. Bell Labs Technical Journal, 9(3):57– 71, Autumn 2004. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Joseph:2005:CWW

[JJM^{+}05] Jean-Philippe Joseph, Paul Justl, Francis R. Magee, Jr., Amit Mukhopadhyay, and Dong Sun. Converged wireline-wireless network evolution: Opportunities and challenges. Bell Labs Technical Journal, 10(2):57– 80, Summer 2005. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Jackson:2002:CSA

[JK02] Donald S. Jackson and Frederick F. Kunzinger. Calculation of system availability using traffic statistics. Bell Labs Technical Journal, 7(3): 139–150, Autumn 2002. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Joyce:2006:PAM

[JLD^{+}06] Toby Joyce, Edward J. Lisay, Jr., David E. Dalton, Jeff M. Punch, Michael S. Shellmer, Shirish N. Kher, and Suresh Goyal. Piecewise analysis and modeling of circuit pack temperature cycling data. Bell Labs Technical Journal, 11(3):21–37, Autumn 2006. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Joyce:2003:EVA

[JLN03] Toby Joyce, Edward J. Lisay, and Robert L. Nickerson. An excel VBA application of Hoadley’s QMP algorithm. Bell Labs Technical Journal, 8(1):233–237, Spring 2003. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Jrad:2004:MQB

[JMS04] Ahmad Jrad, Thomas Morawski, and Louise Spergel. A model for quantifying business continuity preparedness risks for telecommunications networks. Bell Labs Technical Journal, 9(2):107– 123, Summer 2004. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Jai:2000:ESI

[JOR00] Benchiao Jai, Michael Ogg, and Aleta Ricciardi. Effortless software interoperability with Jini connection technology. Bell Labs Technical Journal, 5(2):88– 101, Summer 2000. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Joseph:2008:INS

[Jos08] Jean-Philippe Joseph. IMS network signaling peering: Challenges and proposal. Bell Labs Technical Journal, 12 (4):33–48, Winter 2008. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Kowtarapu:2009:NSI

[KAGS09] Chakravarthy Kowtarapu, Chetan Anand, Guruprasad K. G., and Shishir Sharma. Network separation and IPsec CA certificates-based security management for 4G networks. Bell Labs Technical Journal, 13(4):245– 255, Winter 2009. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Kebir:2000:MSS

[Keb00] Youcef Kebir. A model for the sizing of software updates. Bell Labs Technical Journal, 5(2):102–120, Summer 2000. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Kozik:2000:OPE

[KFL00] Jack Kozik, Igor Faynberg, and Hui-Lan Lu. On opening PSTN to enhanced voice/ data services — The PINT Protocol Solution. Bell Labs Technical Journal, 5(3):153– 165, Autumn 2000. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Kraeutler:2006:NGT

[KGT06] James A. Kraeutler, Richard J. Goode, and Jeffrey R. Towne. Next-generation transport systems: 4th generation MSPPS. Bell Labs Technical Journal, 11(2):13–25, Summer 2006. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Kumar:2002:QSU

[KH02] Suresh Kumar and Min Huang. Quality of service in UMTS wireless networks. Bell Labs Technical Journal, 7(2):49–66, Summer 2002. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Kirova:2008:ERT

[KKKC08] Vassilka Kirova, Neil Kirby, Darshak Kothari, and Glenda Childress. Effective requirements traceability: Models, tools, and practices. Bell Labs Technical Journal, 12 (4):143–157, Winter 2008. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Kermalli:2002:NBS

[KKL^{+}02] Munawar Kermalli, Sanjay Kumar, Michael LaSpisa, Zhengxiang Ma, Samuel Martin, and Krishnamurthy Sreenath. Novel base station radio technologies for third-generation wireless applications. Bell Labs Technical Journal, 7(3):151–167, Autumn 2002. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Kozik:2006:O

[KL06] Jack Kozik and Anne Y.

Lee. Overview. Bell Labs Technical Journal, 10(4): 1–5, Winter 2006. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Kuruganti:2003:BAR

[KM03] Indira Kuruganti and Dennis M. Mancl. Building and applying requirements models. Bell Labs Technical Journal, 8(3):67– 73, Autumn 2003. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Kocan:2002:SCN

[KMS^{+}02] Kristin F. Kocan, Warren A.

Montgomery, Steven A. Siegel, Robert J. Thornberry Jr., and Guy J. Zenner. Service creation for nextgeneration networks. Bell Labs Technical Journal, 7(1): 63–79, Spring 2002. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Kamel:2002:DEP

[KMSW02] Raafat E. Kamel, Martin H. Meyers, Susan Wu Sanders, and Carl F. Weaver. Dynamic effects of power control on third-generation system capacity. Bell Labs Technical Journal, 7(3):127– 137, Autumn 2002. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Kocan:2002:ASN

[Koc02] Kristin F. Kocan. Advances in services, networks, and technology addressing service providers’ future needs. Bell Labs Technical Journal, 7 (1):1–2, Spring 2002. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Kocan:2003:CGP

[Koc03] Kristin F. Kocan. Capitalizing on the growing presence of IP networks in the telecom space. Bell Labs Technical Journal, 8 (1):1–3, Spring 2003. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Kocan:2004:AIC

[Koc04] Kristin F. Kocan. Areas of innovation and creativity. Bell Labs Technical Journal, 9(1):1–3, Spring 2004. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Kocan:2005:MNG

[Koc05] Kristin F. Kocan. Moving to next-generation communication networks. Bell Labs Technical Journal, 10 (1):1–3, Spring 2005. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Kocan:2007:ATN

[Koc07] Kristin F. Kocan. Achieving the transformation to nextgeneration networks. Bell Labs Technical Journal, 12 (1):1–4, Spring 2007. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Kolesnikov:2009:AIS

[Kol09] Vladimir Kolesnikov. Advances and impact of secure function evaluation. Bell Labs Technical Journal, 14 (3):187–192, Autumn 2009. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Korotky:2006:I

[Kor06] Steven K. Korotky. Introduction. Bell Labs Technical Journal, 11(2):1–2, Summer 2006. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Korotky:2009:CAN

[KP09] Steven K. Korotky and Thomas Pfeiffer. Continuing advances in next-generation communication technologies, services, and networks. Bell Labs Technical Journal, 14 (1):1–5, Spring 2009. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Kocan:2006:NSA

[KRA06a] Kristin F. Kocan, William D. Roome, and Vinod Anupam. A novel software approach for service brokering in advanced service architectures. Bell Labs Technical Journal, 11(1):5–20, Spring 2006. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Kocan:2006:SCI

[KRA06b] Kristin F. Kocan, William D. Roome, and Vinod Anupam. Service capability interaction management in IMS using the Lucent Service Broker^{T M} product. Bell Labs Technical Journal, 10 (4):217–232, Winter 2006. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Kessler:2008:SOI

[KRLV08] Marcus Kessler, Andreas Reifert, Dominik Lamp, and Thomas Voith. A serviceoriented infrastructure for providing virtualized networks. Bell Labs Technical Journal, 13(3):111– 127, Autumn 2008. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Krock:2004:EQC

[Kro04] Richard E. Krock. Effective quality control during disaster recovery. Bell Labs Technical Journal, 9(2):163– 171, Summer 2004. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Kocan:2006:NFB

[KS06] Kristin F. Kocan and Tippure S. Sundresh. Networks of the future become reality. Bell Labs Technical Journal, 11(1):1–4, Spring 2006. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Krogfoss:2008:CAO

[KSA08] Bill Krogfoss, Lev Sofman, and Anshul Agrawal. Caching architectures and optimization strategies for IPTV networks. Bell Labs Technical Journal, 13(3):13– 28, Autumn 2008. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Kuo:2003:ANL

[KSO03] Hong-Kwang J. Kuo, Olivier Siohan, and Joseph P. Olive.

Advances in natural language call routing. Bell Labs Technical Journal, 7(4):155– 170, Winter 2003. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Kim:2007:ANA

[KT07] Doh-Suk Kim and Ahmed Tarraf. ANIQUE+: A new American national standard for non-intrusive estimation of narrowband speech quality. Bell Labs Technical Journal, 12(1):221– 236, Spring 2007. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Krupenkin:2005:ETS

[KTKH05] Tom Krupenkin, J. Ashley Taylor, Paul Kolodner, and Marc Hodes. Electrically tunable superhydrophobic nanostructured surfaces. Bell Labs Technical Journal, 10 (3):161–170, Autumn 2005. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Kelly:2000:MDB

[KTW00] Van E. Kelly, Catharine P. Thomas, and Huiyu Wang. Managing data-based systems across releases using historical data dictionaries. Bell Labs Technical Journal, 5(2):121–133, Summer 2000. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Kozik:2003:PSB

[KUV03] Jack Kozik, Musa R. Unmehopa, and Kumar V. Vemuri. A Parlay and SPIRITS-based architecture for service mediation. Bell Labs Technical Journal, 7(4): 105–122, Winter 2003. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Kurien:2003:O

[KW03] Thomas V. Kurien and Gary N. Weber. Overview. Bell Labs Technical Journal, 8(2):1–2, Summer 2003. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Kimber:2006:MPD

[KZFB06] Douglas A. Kimber, Xuemei Zhang, Paul H. Franklin, and Eric J. Bauer. Modeling planned downtime. Bell Labs Technical Journal, 11 (3):7–19, Autumn 2006. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Labrogere:2008:CPT

[Lab08] Paul Labrogere. Com 2.0: A path towards web communicating applications. Bell Labs Technical Journal, 13(2):19– 24, Summer 2008. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Larvet:2008:SAD

[Lar08] Philippe Larvet. Semantic application design. Bell Labs Technical Journal, 13(2):75– 91, Summer 2008. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Lau:2000:VRA

[Lau00] Vincent K. N. Lau. Variablerate adaptive channel coding for CDMA — Reverse Link. Bell Labs Technical Journal, 5(4):138–156, Winter 2000. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Luettmann:2007:MMA

[LB07] Bjoern M. Luettmann and Adam C. Bender. Man-in-the-middle attacks on autoupdating software. Bell Labs Technical Journal, 12 (3):131–138, Autumn 2007. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Liu:2004:SBS

[LBF^{+}04] Chung-Zin Liu, Bill Bushnell, Jim Freeburg, Tod Sizer, Michael Recchione, and Tom Anderson. SIP-based services architecture across wireless and wireline access. Bell Labs Technical Journal, 9(3): 5–13, Autumn 2004. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Li:2008:CNC

[LC08] Xiang Yang Li and Yigang Cai. Converged network common charging controller function. Bell Labs Technical Journal, 13(2):161– 183, Summer 2008. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Li:2009:DSU

[LCBM09] Li Erran Li, Aiyou Chen, Tian Bu, and Scott Miller. Detecting subscribers using NAT devices in wireless data networks. Bell Labs Technical Journal, 14(2):223– 233, Summer 2009. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Liu:2002:IDB

[LCM^{+}02] Youjian Liu, Tai-Ann Chen, Martin H. Meyers, Vincent K. N. Lau, and Jung-Tao Liu. Iterative detection for BLAST systems with an arbitrary number of receive antennas. Bell Labs Technical Journal, 7(3):5– 25, Autumn 2002. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Lopez:2005:MBF

[LDFK05] Daniel López, Ricardo S.

Decca, Ephraim Fischbach, and Dennis E. Krause. MEMS-based force sensor: Design and applications. Bell Labs Technical Journal, 10 (3):61–80, Autumn 2005. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Lautenschlaeger:2009:DCG

[LE09] Wolfram Lautenschlaeger and Gert J. Eilenberger. Dimensioning of carrier grade packet transport for simple network operation. Bell Labs Technical Journal, 14 (1):147–160, Spring 2009. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Liu:2006:OTT

[LGC06] Xiang Liu, Douglas M. Gill, and Sethumadhavan Chandrasekhar. Optical technologies and techniques for high bit rate fiber transmission. Bell Labs Technical Journal, 11(2):83–104, Summer 2006. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Li:2007:GPD

[LGH07] Dalton Li, Andrew Guo, and Brian Huang. Generic policy decision function framework. Bell Labs Technical Journal, 12(1):123–129, Spring 2007. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Lanzerotti:2002:SRB

[LGTM02] Louis J. Lanzerotti, Dale E.

Gary, David J. Thomson, and Carol G. Maclennan. Solar radio burst event (6 April 2001) and noise in wireless communications systems. Bell Labs Technical Journal, 7(1):159–163, Spring 2002. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Leng:2005:MDC

[LHMS05] Bing Leng, Randall K.

Hands, Manoj K. Mahapatra, and James R. Stuhlmacher. Maintaining data consistency across distributed elements using XML/XSLT. Bell Labs Technical Journal, 9(4):117– 131, Winter 2005. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Liu:2008:GSP

[Liu08] Hans Peng Liu. Generic service provisioning and invocation in IMS. Bell Labs Technical Journal, 12 (4):173–178, Winter 2008. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Lui:2000:ESM

[LKL00] Anthony Y. Lui, Shanika A.

Karunasekera, and Chokchai Leangsuksun. The Enhanced Service Manager: A service management system for nextgeneration networks. Bell Labs Technical Journal, 5(3): 130–144, Autumn 2000. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Lau:2002:OST

[LLCM02] Vincent K. N. Lau, Youjian Liu, Tai-Ann Chen, and Martin H. Meyers. Optimal space-time scheduling for block fading channels with partial power feedbacks. Bell Labs Technical Journal, 7(3): 27–46, Autumn 2002. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Li:2009:DCP

[LLZH09] Dalton Li, David Lin, Grace Zhao, and Brian Huang. Design and correctness proof of a security protocol for mobile banking. Bell Labs Technical Journal, 14(1): 259–265, Spring 2009. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Lamy:2006:SPO

[LMMV06] Patrice Lamy, Hector Menendez, Thomas Müller, and Eve L. Varma. Status and perspectives for the optical control plane. Bell Labs Technical Journal, 11 (2):171–189, Summer 2006. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Lieuwen:2006:SDM

[LMR^{+}06] Daniel F. Lieuwen, Todd C.

Morgan, Helmut L. Raether, Satish K. Ramamoorthy, Ming Xiong, and Richard B. Hull. Subscriber data management in IMS networks. Bell Labs Technical Journal, 10(4):197–215, Winter 2006. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Lee:2007:EWV

[LMR07] Minkyu Lee, James W. Mc-Gowan, and Michael C. Recchione. Enabling wireless VoIP. Bell Labs Technical Journal, 11(4):201– 215, Winter 2007. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Ling:2009:TDC

[LMR09] Wei Ling, Salvatore J. Messana, and Paul M. Rominski. Thermal design of central office telecommunication equipment. Bell Labs Technical Journal, 14(3):21– 41, Autumn 2009. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Lui:2001:ERS

[LPST01] Anthony Lui, Ramesh Pattabhiraman, Srinath Subramanian, and Prasad Tadicherla. eSAE: A rapid service creation environment for nextgeneration services. Bell Labs Technical Journal, 6(2): 18–29, Summer 2001. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Lagerberg:2002:WST

[LPW02] Ko Lagerberg, Dirk-Jaap Plas, and Maarten Wegdam. Web services in thirdgeneration service platforms. Bell Labs Technical Journal, 7(2):167–183, Summer 2002. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Lennert:2004:ABS

[LRR^{+}04] Joseph F. Lennert, William Retzner, Monica G. Rodgers, Bernard G. Ruel, Srini Sundararajan, and Paul D. Wolfson. The automated backup solution — safeguarding the communications network infrastructure. Bell Labs Technical Journal, 9(2):59– 84, Summer 2004. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Lawrence:2001:ISD

[LS01] Victor B. Lawrence and Kazem Sohraby. Internet service delivery, application platforms, and product evolution. Bell Labs Technical Journal, 6(2):1– 5, Summer 2001. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Lemke:2007:ABT

[LS07] Michael J. Lemke and Albert J. Sawyer. Alternative backhaul transport techniques for VoIP and PTT in 1x Ev-Do. Bell Labs Technical Journal, 11(4): 273–289, Winter 2007. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Lifton:2005:RBA

[LSF05] Victor A. Lifton, Steve Simon, and Robert E. Frahm. Reserve battery architecture based on superhydrophobic nanostructured surfaces. Bell Labs Technical Journal, 10 (3):81–85, Autumn 2005. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Lubachevsky:2000:FSM

[Lub00] Boris D. Lubachevsky. Fast simulation of multicomponent dynamic systems. Bell Labs Technical Journal, 5(2): 134–156, Summer 2000. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Lee:2008:ACE

[LUV08] Anne Lee and Carlos Urrutia-Valdés. An analysis of capital expenses and performance trade-offs among IMS CSCF deployment options. Bell Labs Technical Journal, 12 (4):21–31, Winter 2008. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Leng:2008:ANC

[LWM08] Bing Leng, Dave P. Weeks, and Manoj K. Mahapatra. Automating network configuration tasks using multilevel modeling. Bell Labs Technical Journal, 12(4):83– 101, Winter 2008. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Malone:2006:ANA

[MA06] Bernard L. Malone III and Abhaya Asthana. Analyzing network availability of a mobile data network: A case study. Bell Labs Technical Journal, 11(3):47– 56, Autumn 2006. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Ma:2007:DBP

[Ma07] David C. Ma. Deterministicbased performance modeling of a cluster of nodes handling subscriber profile query and update in CDMA mobile switching center. Bell Labs Technical Journal, 12 (1):247–261, Spring 2007. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Macwan:2004:AIA

[Mac04] Anil Macwan. Approach for identification and analysis of human vulnerabilities in protecting telecommunications infrastructure. Bell Labs Technical Journal, 9(2): 85–89, Summer 2004. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Malone:2004:WSR

[Mal04] Bernard L. Malone III. Wireless search and rescue: Concepts for improved capabilities. Bell Labs Technical Journal, 9(2):37–49, Summer 2004. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Mankiewich:2002:Oa

[Man02a] Paul M. Mankiewich. Overview.

Bell Labs Technical Journal, 7(2):1–2, Summer 2002. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Mankiewich:2002:Ob

[Man02b] Paul M. Mankiewich. Overview.

Bell Labs Technical Journal, 7(3):1–3, Autumn 2002. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Mazzarella:2006:AHI

[Maz06] Nick J. Mazzarella. Advantages of harmonized IMSbased charging architecture in different access technologies. Bell Labs Technical Journal, 10(4):109– 115, Winter 2006. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

McKiou:2003:DDF

[MB03] Kevin W. McKiou and Frank R. Buckley Jr. Datadriven fault management within a distributed objectoriented OAM&P framework. Bell Labs Technical Journal, 8(1):157–179, Spring 2003. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Meyer:2005:O

[MB05] John Meyer and Alvin Barshefsky. Overview. Bell Labs Technical Journal, 9 (4):1–4, Winter 2005. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Meeuwissen:2000:FPS

[MBB00] Hendrik B. Meeuwissen, Harold J. Batteram, and John-Luc Bakker. The FRIENDS platform — A software platform for advanced services and applications. Bell Labs Technical Journal, 5(3):59–75, Autumn 2000. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

McGee:2007:UBL

[MBC^{+}07] Andrew R. McGee, Frank A.

Bastry, Uma Chandrashekhar, S. Rao Vasireddy, and Lori A. Flynn. Using the Bell Labs security framework to enhance the ISO 17799/27001 Information Security Management System. Bell Labs Technical Journal, 12(3):39– 54, Autumn 2007. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Mezhoudi:2006:IOT

[MC06] Mohcene Mezhoudi and Chi-Hung Kelvin Chu. Integrating optical transport quality, availability, and cost through reliability-based optical network design. Bell Labs Technical Journal, 11(3):91– 104, Autumn 2006. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

McGee:2003:BFB

[MCP^{+}03] Andrew R. McGee, Chun K.

Chan, David D. Picklesimer, Uma Chandrashekhar, Steven H. Richman, and Edwin O. Lambert. A brink of failure and breach of security detection and recovery system. Bell Labs Technical Journal, 8(1):113– 130, Spring 2003. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

McPhail:2006:RRR

[McP06] Charles Eugene McPhail. Respond, restore, resolve: Achieving 7-nines availability telecommunications systems in the field. Bell Labs Technical Journal, 11 (3):173–189, Autumn 2006. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Martin:2002:TDT

[MD02] Cynthia E. Martin and Jeffrey H. Dunn. Test development and test automation for the PacketStar(R) PSAX product. Bell Labs Technical Journal, 7(1):183– 195, Spring 2002. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Martin:2004:AMC

[MD04] Cynthia E. Martin and Jeffrey H. Dunn. Authentication mechanisms for call control message integrity and origin verification. Bell Labs Technical Journal, 8(4):71– 91, Winter 2004. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

McBride:2003:ICT

[MDMF03] Alan J. McBride, Mario Lobo Del Olmo, Brian T. Martin, and Francis J. Finnegan. Increased competitiveness through component technology for element management systems. Bell Labs Technical Journal, 8(3):31– 49, Autumn 2003. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

McKiou:2000:OMT

[ME00] Kevin W. McKiou and Daniel Esposito. OneLink Manager^{T M} EMS for the 7R/E^{T M} Switch. Bell Labs Technical Journal, 5(4):80– 96, Winter 2000. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Mezhoudi:2006:NSP

[MESG06] Mohcene Mezhoudi, Mohamed L. El-Sayed, and Paul F. Gagen. Network service providers: Practical considerations for optical network planning. Bell Labs Technical Journal, 11 (2):145–155, Summer 2006. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Maknavicius:2008:CSC

[MGKO08] Linas Maknavicius, Arnaud Gonguet, Gunn-Kristin Klungsøyr, and Terje Ormhaug. Converged service continuity: A challenge for the merchant and the architect. Bell Labs Technical Journal, 13 (2):185–192, Summer 2008. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Malis:2005:QCL

[MGS^{+}05] Oana Malis, Claire Gmachl, Deborah L. Sivco, Loren N. Pfeiffer, A. Michael Sergent, and Kenneth W. West. The quantum cascade laser: A versatile high-power semiconductor laser for mid-infrared applications. Bell Labs Technical Journal, 10(3):199– 214, Autumn 2005. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

McMahon:2005:OVA

[MHH05] Thomas J. McMahon, Markus Hofmann, and Volker Hilt. Opportunities for VoIP application integration services. Bell Labs Technical Journal, 9(4):25–30, Winter 2005. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Macwan:2006:RCN

[MHM06] Anil P. Macwan, Robert S. Hanmer, and Kailash K. Mutha. Reliability and communications networks. Bell Labs Technical Journal, 11 (3):1–6, Autumn 2006. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Milonas:2000:ENN

[Mil00] Argyrios C. Milonas. Enterprise networking for the new millennium. Bell Labs Technical Journal, 5(1):73– 94, Spring 2000. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Moganti:2001:FUN

[MK01] Madhav Moganti and Steven F.

Knittel. A framework for understanding network intelligence. Bell Labs Technical Journal, 6(1):163– 191, Spring 2001. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Mukherjee:2005:WER

[MKL^{+}05] Sarit Mukherjee, Anand Kagalkar, John Lin, Sarang Gadgil, and Sanjoy Paul. The wireless edge router: A network processor-based packet data serving node for a CDMA2000 network. Bell Labs Technical Journal, 10 (1):83–104, Spring 2005. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Masood:2008:SXR

[MMBDK08] Khurram Masood, Ashraf Matrawy, Bashar Bou-Diab, and Ross Kouhi. Scalable XML routed network topology generation. Bell Labs Technical Journal, 13(3):87– 93, Autumn 2008. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

McGoogan:2000:ESA

[MMD00] Judith R. McGoogan, Joseph E. Merritt, and Yogesh J. Dave. Evolution of switching architecture to support voice telephony over ATM. Bell Labs Technical Journal, 5(2): 157–168, Summer 2000. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Moongilan:2004:DRI

[Moo04] Dheena D. Moongilan. Distributed RF immunity hardening design approach for telecommunication hardware and installation sites. Bell Labs Technical Journal, 9(2): 51–58, Summer 2004. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Maes:2008:MDS

[MPGS08] Jochen Maes, Michael Peeters, Mamoun Guenach, and Charles Storry. Maximizing digital subscriber line performance. Bell Labs Technical Journal, 13(1):105– 115, Spring 2008. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Mitra:2001:TTE

[MR01] Debasis Mitra and K. G. Ramakrishnan. Techniques for traffic engineering of multiservice, multipriority networks. Bell Labs Technical Journal, 6(1):139– 151, Spring 2001. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Maltzman:2005:DNU

[MRD^{+}05] Richard Maltzman, Kevin M.

Rembis, Marirosa Donisi, Matthew Farley, Roberto C. Sanchez, and Augustine Y. Ho. Design for networks — The ultimate design for X. Bell Labs Technical Journal, 9(4):5–23, Winter 2005. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Meeuwissen:2007:IPC

[MRL07] Erik Meeuwissen, Paul Reinold, and Cynthia Liem. Inferring and predicting context of mobile users. Bell Labs Technical Journal, 12(2):79– 86, Summer 2007. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Mitra:2000:NDS

[MSSS00] Debasis Mitra, Kenan E.

Sahin, Ravi Sethi, and Avi Silberschatz. New directions in services management. Bell Labs Technical Journal, 5(1): 17–34, Spring 2000. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Malhotra:2003:MSL

[MvHdMvE03] Richa Malhotra, Ronald van Haalen, Ronald de Man, and Michiel van Everdingen. Managing service-level agreements in metro Ethernet networks using backpressure. Bell Labs Technical Journal, 8(2):83–95, Summer 2003. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

McGee:2001:DVP

[MVJ^{+}01] Andrew R. McGee, S. Rao Vasireddy, K. Jeffrey Johnson, Uma Chandrashekhar, Steven H. Richman, and Mohamed El-Sayed. Dynamic virtual private networks. Bell Labs Technical Journal, 6(2): 116–135, Summer 2001. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

McGee:2004:FEN

[MVX^{+}04] Andrew R. McGee, S. Rao Vasireddy, Chen Xie, David D. Picklesimer, Uma Chandrashekhar, and Steven H. Richman. A framework for ensuring network security. Bell Labs Technical Journal, 8(4):7–27, Winter 2004. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Mockus:2000:PRS

[MW00] Audris Mockus and David M. Weiss. Predicting risk of software changes. Bell Labs Technical Journal, 5(2):169– 180, Summer 2000. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Makowski:2006:PPA

[MWW06] Steven Makowski, Anwar Walid, and Indra Widjaja. Packet path assurance for IP networks. Bell Labs Technical Journal, 11(1):99– 106, Spring 2006. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Mizikovsky:2007:CED

[MWZ07] Semyon Mizikovsky, Zhibi Wang, and Hongru Zhu. CDMA 1x Ev-DO security. Bell Labs Technical Journal, 11(4):291–305, Winter 2007. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Ma:2009:RPW

[MZP^{+}09] Zhengxiang Ma, Michael Zierdt, John Pastalan, Arnold Siegel, Tod Sizer, Adriaan J. de Lind van Wijngaarden, Praveen R. Kasireddy, and Dragan M. Samardzija. RADIOSTAR: Providing wireless coverage over Gigabit Ethernet. Bell Labs Technical Journal, 14(1): 7–24, Spring 2009. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Nakata:2003:MEA

[Nak03]  Tsuneo Nakata.  Multi-

event algorithms and protocols for fast and robust distributed mesh provisioning and restoration. Bell Labs Technical Journal, 7(4): 23–39, Winter 2003. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Noirie:2009:SNF

[NDC^{+}09] Ludovic Noirie, Emmanuel Dotaro, Giovanna Carofiglio, Arnaud Dupas, Pascal Pecci, Daniel Popa, and Georg Post. Semantic networking: Flowbased, traffic-aware, and selfmanaged networking. Bell Labs Technical Journal, 14 (2):23–38, Summer 2009. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Neilson:2006:WSS

[NDM^{+}06] David T. Neilson, Christopher R. Doerr, Dan M. Marom, Roland Ryf, and Mark P. Earnshaw. Wavelength selective switching for optical bandwidth management. Bell Labs Technical Journal, 11(2):105– 128, Summer 2006. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Netravali:2000:WNB

[Net00] Arun N. Netravali. When networking becomes second nature: The next 25 years … and beyond. Bell Labs Technical Journal, 5(1):203– 214, Spring 2000. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Nithi:2005:NCR

[NNT05] Nachi K. Nithi, Carl J. Nuzman, and Benjamin Y. C. Tang. Network capacity recovery and efficient capacity deployment in switching centers. Bell Labs Technical Journal, 9(4):83– 100, Winter 2005. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Nagarajan:2008:NGA

[NO08] Ramesh Nagarajan and Sven Ooghe. Next-generation access network architectures for video, voice, interactive gaming, and other emerging applications: challenges and directions. Bell Labs Technical Journal, 13(1):69– 86, Spring 2008. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Nagarajan:2001:FTM

[NQR^{+}01] Ramesh Nagarajan, M. Akber Qureshi, Bharathkumar Ramanna, Christopher J. Hunt, Rakesh Chandwani, Basudeb Das, and Abha Shrotriya. A fault-tolerant MPLS-based control and communication network for the Lucent LambdaRouter. Bell Labs Technical Journal, 6(2):153–169, Summer 2001. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Nguyen:2003:RAS

[NS03] Ai G. Nguyen and Richard L. Sohn. Refactoring administrative services module platform software. Bell Labs Technical Journal, 8(3):105– 110, Autumn 2003. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Nguyen:2009:EDM

[NT09] Van Minh Nguyen and Laurent Thomas. Efficient dynamic multi-step paging for cellular wireless networks. Bell Labs Technical Journal, 14(2):203–221, Summer 2009. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

OReilly:2008:ODP

[OC08] Gerard P. O’Reilly and Chi-Hung Kelvin Chu. Optimal deployment of power reserves across telecom critical infrastructures. Bell Labs Technical Journal, 12 (4):127–141, Winter 2008. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

OShea:2004:LPB

[OJ04] William T. O’Shea and Jeffrey M. Jaffe. Letter from President, Bell Labs, and President, Bell Labs Advanced Technologies. Bell Labs Technical Journal, 9 (2):fmv–fmvi, Summer 2004. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Ooghe:2008:RAC

[OVN08] Sven Ooghe, Adrianus Van Ewijk, and Ramesh Nagarajan. Resource admission control in access networks. Bell Labs Technical Journal, 13 (1):87–104, Spring 2008. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Panchmatia:2006:III

[Pan06] Manish V. Panchmatia. Intelligent integrated insight into wireless networks. Bell Labs Technical Journal, 11 (1):197–202, Spring 2006. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Pallec:2009:TFD

[PBDS09] Michel Le Pallec, Dinh Thai Bui, François Dorgeuille, and Nicolas Le Sauze. Time and frequency distribution over packet switched networks. Bell Labs Technical Journal, 14(2):131–153, Summer 2009. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Pant:2008:RNG

[PCR^{+}08] Himanshu Pant, Chi-Hung Kelvin Chu, Steven H. Richman, Ahmad Jrad, and Gerard P. O’Reilly. Reliability of nextgeneration networks with a focus on IMS architecture. Bell Labs Technical Journal, 12(4):109–125, Winter 2008. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Pouyllau:2009:ETP

[PDDS09] Hélia Pouyllau, Richard Douville, Nabil Bachir Djarallah, and Nicolas Le Sauze. Economic and technical propositions for inter-domain services. Bell Labs Technical Journal, 14(1):185– 202, Spring 2009. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Pant:2006:OAS

[PMCR06] Himanshu Pant, Andrew R. McGee, Uma Chandrashekhar, and Steven H. Richman. Optimal availability and security for IMS-based VoIP networks. Bell Labs Technical Journal, 11(3):211– 223, Autumn 2006. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Powell:2004:LCF

[Pow04] Michael K. Powell. Letter from Chairman, Federal Communications Commission. Bell Labs Technical Journal, 9(2):fmi, Summer 2004. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Popa:2009:FHA

[PPN09] Daniel Popa, Georg Post, and Ludovic Noirie. A fast hierarchical arbitration scheme for multi-Tb/s packet switches with shared memory switching. Bell Labs Technical Journal, 14(2):81– 95, Summer 2009. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Patel:2001:SWI

[PRS01] Sarvar Patel, Zulfikar Ramzan, and Ganapathy S. Sundaram. Security for wireless Internet access. Bell Labs Technical Journal, 6(2): 74–83, Summer 2001. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Prabhakar:2005:OAR

[PRT05] Gokul Prabhakar, Rajeev Rastogi, and Marina Thottan. OSS architecture and requirements for VoIP networks. Bell Labs Technical Journal, 10(1):31–45, Spring 2005. CODEN BLTJFD.

ISSN 1089-7089 (print), 1538-7305 (electronic).

Park:2007:IBR

[PS07] Seyong Park and Alexandro Salvarani. IP-based radio access network implementation cost vs. network security trade-off. Bell Labs Technical Journal, 12(3):95– 99, Autumn 2007. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Prakash:2007:EAC

[PSST07] Subramanian Prakash, Abdol Saleh, Salman Shaikh, and David Tsay. Economic analysis of class 5 migration to IMS. Bell Labs Technical Journal, 12(1):77– 93, Spring 2007. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Peterson:2007:WTI

[PT07] James S. Peterson and Joseph A. Tarallo. Wireless technology issue overview. Bell Labs Technical Journal, 11(4):1–4, Winter 2007. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Pattabhiraman:2007:EAP

[PUV07] Ramesh Pattabhiraman, Musa Unmehopa, and Kumar Vemuri. Enhanced active phone book services: Blended lifestyle services made real!

Bell Labs Technical Journal, 11(4):315–326, Winter 2007. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Patel:2000:RDC

[PWHH00] Vipin N. Patel, Christopher K. Wiese, Fedde M. Hiemstra, and Steven C. Himes. Rapid development and commercialization of products — A business imperative in the global telecommunication landscape. Bell Labs Technical Journal, 5(4):157–170, Winter 2000. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Pfeiffer:2005:NGF

[PWW^{+}05] Loren N. Pfeiffer, Kenneth W. West, Robert L. Willett, Hidefumi Akiyama, and Leonid P. Rokhinson. Nanostructures in GaAs fabricated by molecular beam epitaxy. Bell Labs Technical Journal, 10(3):151– 159, Autumn 2005. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Pyritz:2003:CVE

[Pyr03a] Bill Pyritz. Craftsmanship versus engineering: Computer programming — An art or a science? Bell Labs Technical Journal, 8(3):101– 104, Autumn 2003. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Pyritz:2003:EPT

[Pyr03b] Bill Pyritz. Extreme programming in the telecommunications domain. Bell Labs Technical Journal, 8(3):97– 100, Autumn 2003. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Peelen:2003:SQB

[PZBT03] Bastien Peelen, Miroslav Zivkovic, Dennis Bijwaard, and Harold Teunissen. Supporting QoS in broadband wireless and wired access. Bell Labs Technical Journal, 8(2):65–81, Summer 2003. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Qutub:2005:APM

[QW05] S. Shehryar Qutub and Indaka N. Weerasekera. Advances in policy management standardization in mobility networks. Bell Labs Technical Journal, 10(1):77– 82, Spring 2005. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Radhakrishnan:2004:AAC

[Rad04]  Ganesan  Radhakrishnan.

Adaptive application caching. Bell Labs Technical Journal, 9(1):165–175, Spring 2004.

CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Radhakrishnan:2007:PCN

[Rad07]  Ganesan  Radhakrishnan.

Provisioning in complex networks. Bell Labs Technical Journal, 12(1):41–55, Spring 2007. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Ramakrishnan:2004:SRM

[Ram04] Murali Ramakrishnan. Software release management. Bell Labs Technical Journal, 9(1):205–210, Spring 2004. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Ramirez:2005:CNS

[Ram05] Arthur P. Ramirez. Carbon nanotubes for science and technology. Bell Labs Technical Journal, 10(3): 171–185, Autumn 2005. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Ramirez:2007:CSI

[Ram07]  David Ramirez.  Case

study: ITU-T recommendation X.805 applied to an enterprise environment — banking. Bell Labs Technical Journal, 12(3):55– 64, Autumn 2007. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Rao:2006:AOF

[Rao06] Venkatesh S. Rao. Analysis of optimal failure group distribution in a hierarchical aggregated architecture. Bell Labs Technical Journal, 11(3):83–89, Autumn 2006. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Rauscher:2004:PCI

[Rau04] Karl F. Rauscher. Protecting communications infrastructure. Bell Labs Technical Journal, 9(2):1– 4, Summer 2004. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Rimac:2008:PAC

[RBW08] Ivica Rimac, Sem Borst, and Anwar Walid. Peerassisted content distribution networks: performance gains and server capacity savings. Bell Labs Technical Journal, 13(3):59–69, Autumn 2008. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Ryf:2009:ALM

[RCB^{+}09] Roland Ryf, Gang Chen, Nagesh Basavanhally, Mihaela Dinu, Alex Duque, Yee L. Low, Jay M. Wiesenfeld, Yulia Shapiro, and Randy Giles. The Alcatel– Lucent microprojector: What every cell phone needs. Bell Labs Technical Journal, 14 (3):99–112, Autumn 2009. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Rajkumar:2004:SSB

[RFBJ04] Ajay Rajkumar, Peretz Feder, Steven Benno, and Tom Janiszewski. Seamless SIP-based VoIP in disparate wireless systems and networks. Bell Labs Technical Journal, 9(1):65–82, Spring 2004. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Raghavan:2001:VPN

[RGAK01] Narasimhan Raghavan, Rajat Gopal, Sashidhar Annaluru, and Shasidhar Kura. Virtual private networks and their role in e-business. Bell Labs Technical Journal, 6(2): 99–115, Summer 2001. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Romijn:2003:ENM

[RHSG03] Willem A. Romijn, Geert Jan Hoekstra, Joan Serrat, and Eduardo Grampin. Enhancing network management by applying policy management principles. Bell Labs Technical Journal, 8(1):151– 156, Spring 2003. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Reichmanis:2005:PED

[RKKM05] Elsa Reichmanis, Howard Katz, Christian Kloc, and Ashok Maliakal. Plastic electronic devices: From materials design to device applications. Bell Labs Technical Journal, 10(3):87– 105, Autumn 2005. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Rauscher:2006:EIC

[RKR06] Karl F. Rauscher, Richard E. Krock, and James P. Runyon. Eight ingredients of communications infrastructure: A systematic and comprehensive framework for enhancing network reliability and security. Bell Labs Technical Journal, 11(3):73– 81, Autumn 2006. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Rangarajan:2005:CWH

[RLP05] Sampath Rangarajan, John Lin, and Sanjoy Paul. A campus-wide hybrid 802.11/ 3G network and its implementation. Bell Labs Technical Journal, 10(2):21– 38, Summer 2005. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Royer:2008:MIS

[RNMP08] Julien Royer, Hang Nguyen, Olivier Martinot, and Françoise Prêteux. Multimedia interactive services automation based on content indexing. Bell Labs Technical Journal, 13(2):147–154, Summer 2008. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Richman:2008:RCN

[RP08] Steven H. Richman and Himanshu Pant. Reliability concerns for next-generation networks. Bell Labs Technical Journal, 12(4):103– 108, Winter 2008. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Romijn:2004:MMS

[RPB^{+}04] Willem A. Romijn, Dirk-Jaap Plas, Dennis Bijwaard, Erik Meeuwissen, and Gijs van Ooijen. Mobility management for SIP sessions in a heterogeneous network environment. Bell Labs Technical Journal, 9(3):237– 253, Autumn 2004. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Ramakrishnan:2001:ORS

[RR01] K. G. Ramakrishnan and Manoel A. Rodrigues. Optimal routing in shortest-path data networks. Bell Labs Technical Journal, 6(1):117– 138, Spring 2001. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Rissen:2009:ES

[RS09] Jean-Paul Rissen and Robert Soni. The evolution to 4G systems. Bell Labs Technical Journal, 13(4): 1–5, Winter 2009. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Rangarajan:2004:AVT

[RTM^{+}04] Sampath Rangarajan, Anil Takkallapalli, Sarit Mukherjee, Sanjoy Paul, and Scott Miller. Adaptive VPN: Tradeoff between security levels and value-added services in virtual private networks. Bell Labs Technical Journal, 8(4):93–113, Winter 2004. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Rubio:2009:PGD

[Rub09] Alberto E. Rubio. Practical guide to designing a DVB-SH network for RF engineers. Bell Labs Technical Journal, 13(4):207–225, Winter 2009. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Russo:2004:LCC

[Rus04] Patricia F. Russo. Letter from Chairman and CEO, Lucent Technologies. Bell Labs Technical Journal, 9(2): fmiii, Summer 2004. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Rao:2009:LHR

[RWGS09] Anil M. Rao, Andreas Weber, Sridhar Gollamudi, and Robert Soni. LTE and HSPA+: Revolutionary and evolutionary solutions for global mobile broadband. Bell Labs Technical Journal, 13(4):7–34, Winter 2009. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Sahin:2000:O

[Sah00] Kenan E. Sahin. Overview.

Bell Labs Technical Journal, 5(2):3–6, Summer 2000. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Samuel:2003:LBP

[SAN03] Isaac Samuel, Kishore Arora, and Bhuvarahamurthy Narasimhan. Location-based performancemeasuring techniques in UMTS. Bell Labs Technical Journal, 8(2):15–32, Summer 2003. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Sharp:2001:SBN

[SBC^{+}01] Ronald L. Sharp, Michaela Blott, Michael J. Coss, Bruce K. Ellis, David L. Majette, and Vinay Purohit.

Starburst: Building nextgeneration Internet devices. Bell Labs Technical Journal, 6(2):6–17, Summer 2001. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Sigle:2009:MRI

[SBEW09] Rolf Sigle, Oliver Blume, Lutz Ewe, and Wieslawa Wajda. Multi-radio infrastructure for 4G. Bell Labs Technical Journal, 13(4): 257–276, Winter 2009. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Sabnani:2009:O

[SBG09] Krishan K. Sabnani, Milind M. Buddhikot, and Katherine H. Guo. Overview. Bell Labs Technical Journal, 14(2):1– 4, Summer 2009. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Skraba:2008:PIM

[SBLD08] Ryan Skraba, Olivier Le Berre, Patrick Legrand, and Johann Daigremont. Promoting innovative mobile applications via a third party IMSenabled ecosystem. Bell Labs Technical Journal, 13(2):57– 73, Summer 2008. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Silverman:2000:TVN

[SBS00] Kurt S. Silverman, Michael R.

Brenner, and Gregory E. Shannon. Toward a vision for network and service management. Bell Labs Technical Journal, 5(4):21– 30, Winter 2000. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Sijben:2001:ALC

[SBSS01] Paul G. A. Sijben, Mike Buckley, John P. L. Segers, and Louise F. A. Spergel. Application-level control of IP networks: IP beyond the Internet. Bell Labs Technical Journal, 6(1):98– 116, Spring 2001. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Schlaerth:2000:SNM

[Sch00] Joseph P. Schlaerth. Service and network management strategies for ATM in wireless networks. Bell Labs Technical Journal, 5(4):126– 137, Winter 2000. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Sridhar:2008:EED

[SDC08] Kamakshi Sridhar, Gérard Damm, and Hakki C. Cankaya. End-to-end diagnostics in IPTV architectures. Bell Labs Technical Journal, 13 (1):29–34, Spring 2008. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Stiliadis:2007:ESD

[SFK^{+}07] Dimitrios Stiliadis, Andrea Francini, Sanjay Kamat, Mansoor Alicherry, Adiseshu Hari, Pramod V. Koppol, Ashok K. Gupta, and Dor Skuler. Evros: A servicedelivery platform for extending security coverage and IT reach. Bell Labs Technical Journal, 12(3):101– 119, Autumn 2007. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Soljanin:2009:NCE

[SGK09] Emina Soljanin, Piyush Gupta, and Gerhard Kramer. Network coding for efficient network multicast. Bell Labs Technical Journal, 14 (3):157–166, Autumn 2009. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Siddiqui:2004:EEQ

[SGRP04] M. Ali Siddiqui, Katherine Guo, Sampath Rangarajan, and Sanjoy Paul. End-to-end QoS support for SIP sessions in CDMA2000 networks. Bell Labs Technical Journal, 9(3): 135–153, Autumn 2004. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Sharma:2004:IDI

[Sha04] Vivek Sharma. Intrusion detection in infrastructure wireless LANs. Bell Labs Technical Journal, 8(4):115– 119, Winter 2004. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Shaw:2007:UCC

[Sha07]  Bradley D. Shaw.  The

use of conduction to cool small form-factor pluggable optical transceivers. Bell Labs Technical Journal, 12 (2):143–146, Summer 2007. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Sheehan:2007:SPF

[She07] Michael J. Sheehan. Secure parallel file distribution through a streaming worm network. Bell Labs Technical Journal, 12(1): 237–246, Spring 2007. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Siegel:2001:NSA

[Sie01] Steven A. Siegel. Networksupported applications. Bell Labs Technical Journal, 6(2): 30–46, Summer 2001. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Streilein:2001:ERG

[SJ01] Patricia Streilein and Joe John. Enabling revenuegenerating services — The evolution of next-generation networks. Bell Labs Technical Journal, 6(1):3–12, Spring 2001. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Sun:2004:SEA

[SJM^{+}04] Dong Sun, Jean-Philippe Joseph, Francis R. Magee, Jr., Amit Mukhopadhyay, and Benjamin Tang. A SIP-enabled all-IP architecture for converged nextgeneration networks. Bell Labs Technical Journal, 9(3): 15–37, Autumn 2004. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Schuh:2009:TTG

[SLJ^{+}09] Karsten Schuh, Eugen Lach, Bernhard Junginger, Gustav Veith, Jeremie Renaudier, Gabriel Charlet, and Patrice Tran. 8 Tb/s (80 × 107 Gb/s) DWDM NRZ-VSB transmission over 510 km NZDSF with 1 bit/s/Hz spectral efficiency. Bell Labs Technical Journal, 14(1):89– 104, Spring 2009. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Sayadi:2009:MIE

[SLK^{+}09] Bessem Sayadi, Yann Leprovost, Sylvaine Kerboeuf, Marie Line Alberi-Morel, and Laurent Roullet. MPE-IFEC: An enhanced burst error protection for DVB-SH systems. Bell Labs Technical Journal, 14(1):25– 40, Spring 2009. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Sayadi:2009:ERM

[SLKAM09] Bessem Sayadi, Yann Leprovost, Sylvaine Kerboeuf, and Marie Line Alberi-Morel. Efficient repair mechanism of real-time broadcast services in hybrid DVB-SH and cellular systems. Bell Labs Technical Journal, 14(1):41– 54, Spring 2009. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Sohraby:2002:NGS

[SLM02] Kazem Sohraby, Victor Lawrence, and Sukant Mohapatra. The Next-Generation Signaling System for switching and control in integrated optical networks. Bell Labs Technical Journal, 7(1):121– 133, Spring 2002. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Shen:2009:MHR

[SLW^{+}09] Gang Shen, Jimin Liu, Dongyao Wang, Jikang Wang, and Shan Jin. Multihop relay for next-generation wireless access networks. Bell Labs Technical Journal, 13 (4):175–193, Winter 2009. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Shah:2005:IMC

[SM05] Nikhil B. Shah and Sukant K. Mohapatra. Integrating and managing converged multiservice networks. Bell Labs Technical Journal, 10(1): 139–156, Spring 2005. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Sienel:2009:OTD

[SMZ^{+}09] Jürgen Sienel, Alberto León Martín, Carlos Baladrón Zorita, Laurent-Walter Goix, Álvaro Martínez Reol, and Belén Carro Martínez. OP-UCE: A telco-driven service mash-up approach. Bell Labs Technical Journal, 14 (1):203–218, Spring 2009. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Solari:2007:DS

[Sol07]  Carlos Solari.  Designing

for security. Bell Labs Technical Journal, 12(3):1– 6, Autumn 2007. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Sripad:2003:SCN

[SPNR03] Anekal B. Sripad, Stanislas Peijffers, Ramesh Nagarajan, and Jaap W. Rodenburg. Signaling communications network architectures for Service Intelligent^{T M} optical transport networks. Bell Labs Technical Journal, 7(4): 3–22, Winter 2003. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Schaepperle:2009:ETF

[SR09] Joerg Schaepperle and Andreas Rüegg. Enhancement of throughput and fairness in 4G wireless access systems by non-orthogonal signaling. Bell Labs Technical Journal, 13(4):59–77, Winter 2009. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Skraba:2008:ESI

[SRR08] Ryan Skraba, David Roxburgh, and Simon Ringland. Enabling structured interactions between third party mobile applications and the user terminal. Bell Labs Technical Journal, 13(2):49– 55, Summer 2008. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Strauss:2008:MTN

[SSC^{+}08] Thomas Strauss, Yuzhong Shen, Duan Chen, Juan Wu, and Hong Tang. Media telecom network: A Telco approach to peer-to-peer delivery of streaming content. Bell Labs Technical Journal, 13(3):71–77, Autumn 2008. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Shah:2000:WDN

[SSM^{+}00] Nitin J. Shah, Anil S. Sawkar, John A. Marinho, Krishan K. Sabnani, Thomas F. La Porta, and Tung Ching Chiang. Wireless data networking, standards, and applications. Bell Labs Technical Journal, 5(1):130– 149, Spring 2000. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Sijben:2000:BBD

[SSSK00] Paul G. A. Sijben, John P. L. Segers, Louise F. A. Spergel, and Jack Kozik. Building the bridge: Devising an architecture to migrate voice-band calls to packet transport and multimedia services. Bell Labs Technical Journal, 5(3): 166–185, Autumn 2000. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Stoll:2007:RPW

[STB07] Dieter Stoll, Wolfgang Thomas, and Martin Belzner. The role of pseudo-wires for layer 2 services in intelligent transport networks. Bell Labs Technical Journal, 12(1): 207–220, Spring 2007. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Sundresh:2005:MCS

[Sun05] Tippure S. Sundresh. Macroscopic characterization of software and its relationship to reliability. Bell Labs Technical Journal, 10(1): 169–174, Spring 2005. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Sundresh:2006:SRM

[Sun06] Tippure S. Sundresh. Semantic reliability of multi-agent intelligent systems. Bell Labs Technical Journal, 11 (3):225–236, Autumn 2006. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Sijben:2002:MCM

[SvWdBvdG02] Paul Sijben, Willem van Willigenburg, Michel de Boer, and Sietse van der Gaast. Middleboxes: Controllable media firewalls. Bell Labs Technical Journal, 7(1):141– 157, Spring 2002. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Saniee:2009:DPR

[SW09] Iraj Saniee and Indra Widjaja. Design and performance of randomized schedules for time-domain wavelength interleaved networks. Bell Labs Technical Journal, 14(2):97– 111, Summer 2009. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Sweldens:2003:O

[Swe03] Wim Sweldens. Overview. Bell Labs Technical Journal, 8(3):1–2, Autumn 2003. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Storry:2008:ADS

[SŽVdLvW08] Charles Storry, Miroslav Živković, Jan Verlinden, and Adriaan J. de Lind van Wijngaarden. Aspects of dynamic spectrum management level 3. Bell Labs Technical Journal, 13(1):117– 127, Spring 2008. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Tushman:2004:MSI

[TA04] Michael Tushman and Philip Anderson, editors. Managing strategic innovation and change: a collection of readings. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, second edition, 2004. ISBN 0-19-513577-6 , 0-19-513578-4 (paperback). xix + 635 pp. LCCN HD31 .M29425 2004.

URL http://www.loc.gov/
catdir/enhancements/fy0615/
2003061008-d.html;   http:
//www.loc.gov/catdir/enhancements/
fy0636/2003061008-t.html;
 http://www.loc.gov/catdir/
enhancements/fy0723/2003061008-
b.html.

Tahat:2007:DMS

[TAS^{+}07] Luay H. Tahat, Erik E. Anderlind, Ashok P. Shah, Anthony J. Wycklendt, Edward A. Clark, and Dupyo Choi. Dual mode service: Learning from integrating an IP multimedia subsystem service into existing live networks. Bell Labs Technical Journal, 11 (4):135–150, Winter 2007. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Trickey:2004:ACA

[TB04] Hemi Trickey and Alvin Barshefsky. An access control architecture for managing large-scale network applications. Bell Labs Technical Journal, 8(4):29– 38, Winter 2004. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Tuijn:2008:SMS

[TB08] Jasper Aartse Tuijn and Dennis Bijwaard. Spanning a multimedia session across multiple devices. Bell Labs Technical Journal, 12 (4):179–193, Winter 2008. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

TJoens:2008:MHA

[TBVV08] Yves T’Joens, Tom Bostoen, Sven Van den Bosch, and Piet Vandaele. Managing home and access domains in modern broadband networks. Bell Labs Technical Journal, 13(1):247–262, Spring 2008. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Thompson:2007:PMD

[TF07] Gerald R. Thompson and Lori A. Flynn. Polymorphic malware detection and identification via contextfree grammar homomorphism. Bell Labs Technical Journal, 12(3):139– 147, Autumn 2007. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Trappeniers:2008:CME

[TGC^{+}08] Lieven Trappeniers, Marc Godon, Laurence Claeys, Olivier Martinot, and Emmanuel Marilly. Crossmedia experiences: Ambient community interactions in the city. Bell Labs Technical Journal, 13(2):5– 11, Summer 2008. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Tsay:2005:MCA

[TM05] David Tsay and Elissa Pariser Matthews. Metrics for comparative analysis of operations competency. Bell Labs Technical Journal, 10(1): 175–179, Spring 2005. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Tonello:2002:AMM

[Ton02] Andrea M. Tonello. Asynchronous multicarrier multiple access: Optimal and sub-optimal detection and decoding. Bell Labs Technical Journal, 7(3):191–217, Autumn 2002. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Torabi:2000:SMN

[Tor00] Mohammad Torabi. A shift in the mobile network service provisioning paradigm. Bell Labs Technical Journal, 5(3): 112–129, Autumn 2000. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Torabi:2002:MVR

[Tor02] Mohammad Torabi. Mobile virtual reality services. Bell Labs Technical Journal, 7(2): 185–193, Summer 2002. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Trimble:2002:JCH

[Tri02] William F. Trimble. Jerome C. Hunsaker and the rise of American aeronautics. Smithsonian history of aviation and spaceflight series. Smithsonian Institution Press, Washington, DC, USA, 2002. ISBN 1-58834-006-6. xi + 284 pp. LCCN TL540.H85 T75 2002. URL http:/

/www.loc.gov/catdir/toc/
fy032/2001042038.html.

Tsay:2000:ASO

[Tsa00] David Tsay. Analysis of service operations cost for TDM and packet tandem networks. Bell Labs Technical Journal, 5(4):97–112, Winter 2000. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Thottan:2003:SSB

[TSC^{+}03] Marina K. Thottan, George K. Swanson, Michael Cantone, Tin Kam Ho, Jennifer Ren, and Sanjoy Paul. SEQUIN: An SNMP-based MPLS network monitoring system. Bell Labs Technical Journal, 8(1): 95–111, Spring 2003. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Unmehopa:2001:SMI

[UGDS01] Musa R. Unmehopa, Michel L. F. Grech, Janusz A. Dobrowolski, and John J. Stanaway Jr. The support of mobile Internet applications in UMTS networks through the open service access. Bell Labs Technical Journal, 6(2): 47–64, Summer 2001. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Urrutia-Valdes:2004:SWN

[UVMES^{+}04] Carlos Urrutia-Valdés, Amit Mukhopadhyay, Mohamed El-Sayed, Zulfiquar Sayeed, and Dong Sun. SIP in 3G wireless networks: Service models, architecture, and network design. Bell Labs Technical Journal, 9(3):101– 126, Autumn 2004. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Urrutia-Valdes:2006:PAI

[UVMES06] Carlos Urrutia-Valdés, Amit Mukhopadhyay, and Mohamed El-Sayed. Presence and availability with IMS: Applications architecture, traffic analysis, and capacity impacts. Bell Labs Technical Journal, 10(4): 101–107, Winter 2006. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Vanguri:2005:WPS

[Van05] Raghava Rao V. Vanguri. Wireless provisioning service solution and deployment — A real-life experience in global project management. Bell Labs Technical Journal, 9(4):35–48, Winter 2005. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

vanderGaast:2008:EPC

[vdGB08] Sietse van der Gaast and Dennis Bijwaard. Efficiency of personalized content distribution. Bell Labs Technical Journal, 13(2):135– 145, Summer 2008. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

vanderGaast:2004:QSS

[vdGHM04] Sietse van der Gaast, Abdelkader Hajjaoui, and Erik Meeuwissen. Quality of service for SIP sessions in 3GPP-based networks. Bell Labs Technical Journal, 9(3): 127–134, Autumn 2004. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Verhoeyen:2008:CSA

[VDR08] Marc Verhoeyen, Danny De Vleeschauwer, and Dave Robinson. Content storage architectures for boosted IPTV service. Bell Labs Technical Journal, 13(3):29– 43, Autumn 2008. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Verhoeyen:2009:PIN

[VDR09] Marc Verhoeyen, Danny De Vleeschauwer, and Dave Robinson. Public IP network infrastructure evolutions to support emerging digital video services. Bell Labs Technical Journal, 14 (2):39–55, Summer 2009. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Vetter:2008:RMS

[VFRO08] Peter Vetter, François Fredricx, Govinda Rajan, and Karsten Oberle. Recommendations for a multi-service access architecture from the European MUSE project. Bell Labs Technical Journal, 13 (1):11–28, Spring 2008. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Valenzuela:2005:FWC

[VK05] Reinaldo A. Valenzuela and Kristin F. Kocan. Future wireless communications. Bell Labs Technical Journal, 10(2):1–3, Summer 2005. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Vlachos:2003:NNP

[Vla03] Kyriakos G. Vlachos. A novel network processor for security applications in highspeed data networks. Bell Labs Technical Journal, 8(1): 131–149, Spring 2003. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Caenegem:2008:MVQ

[VSL^{+}08] Tom N. M. Van Caenegem, Kris O. Struyve, Koen Laevens, Danny De Vleeschauwer, and Randy Sharpe. Maintaining video quality and optimizing video delivery over the bandwidth constrained DSL last mile through intelligent packet drop. Bell Labs Technical Journal, 13(1):53–68, Spring 2008. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Viswanathan:2009:NSA

[VVL^{+}09] Harish Viswanathan, Tom Van Leeuwen, Werner Liekens, Bruno Van Bogaert, and Willem Acke. Network and service architecture for emerging services based on home sensor networks. Bell Labs Technical Journal, 14 (2):235–249, Summer 2009. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

vanVeen:2005:ATE

[vVWBH05] Doutje T. van Veen, Marcus K. Weldon, Charles C. Bahr, and Edward E. Harstead. An analysis of the technical and economic essentials for providing video over fiber-to-the-premises networks. Bell Labs Technical Journal, 10 (1):181–200, Spring 2005. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Viswanathan:2008:PCA

[VW08] Ramesh Viswanathan and Thomas L. Wood. Portable call agent: A model for rapid development and emulation of network services. Bell Labs Technical Journal, 12(4):159–172, Winter 2008. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Vasireddy:2004:SPC

[VWC^{+}04] Rao Vasireddy, Steven Wolter, Uma Chandrashekhar, Robert J. Thornberry Jr., and Andrew R. McGee. Security posture for civilian and noncivilian networks. Bell Labs Technical Journal, 8(4):187– 202, Winter 2004. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Whiting:2008:PRD

[WAB^{+}08] Phil Whiting, Alexei Ashikhmin, Sem Borst, Jean Jennen, Gerhard Kramer, Adriaan J. de Lind van Wijngaarden, and Miroslav Živković. Performance results for digital subscriber line precoders. Bell Labs Technical Journal, 13(1):147–161, Spring 2008. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Wang:2000:RCW

[Wan00] Yih-Chen S. Wang. Reducing the complexity of wireless Q3 adapter development. Bell Labs Technical Journal, 5(2): 181–188, Summer 2000. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Wang:2001:TGI

[Wan01]  Jay J. Wang.  A traf-

fic generator implementation framework for packet switched networks. Bell Labs Technical Journal, 6(1): 152–162, Spring 2001. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Wang:2006:WCM

[Wan06] Hongzhou Wang. Warranty cost models considering imperfect repair and preventive maintenance. Bell Labs Technical Journal, 11(3): 147–159, Autumn 2006. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Wang:2007:XBT

[Wan07] Daping Wang. An XMLbased testing strategy for probing security vulnerabilities in the diameter protocol. Bell Labs Technical Journal, 12(3):79–93, Autumn 2007. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

White:2008:PMU

[WC08] Susan A. White and Mark M.

Clougherty. PSTN migration using IMS. Bell Labs Technical Journal, 13(1): 199–219, Spring 2008. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Wang:2005:FSN

[WDC^{+}05] Jian Jim Wang, Xuegong Deng, Lei Chen, Paul F. Sciortino Jr., Feng Liu, Stephen Tai, Xiaoming Liu, Anguel Nikolov, and Barry J. Weinbaum. Free-space nanooptical devices and integration: Design, fabrication, and manufacturing. Bell Labs Technical Journal, 10 (3):107–127, Autumn 2005. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Weimann:2009:IDC

[WHK^{+}09] Nils Weimann, Vincent Houtsma, Rose Kopf, Yves Baeyens, Joseph Weiner, Alaric Tate, John Frackoviak, Young-Kai Chen, Gregory Raybon, Jean Godin, Muriel Riet, Virginie Nodjiadjim, Agnieszka Konczykowska, Jean-Yves Dupuy, Filipe Jorge, André Scavennec, and Gabriel Charlet. InP DHBT circuits: From device physics to 40Gb/s and 100Gb/s transmission system experiments. Bell Labs Technical Journal, 14(3):43– 62, Autumn 2009. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Wischoeffer:2004:TDP

[Wis04] W. Shawn Wischoeffer. The transcend data production solution. Bell Labs Technical Journal, 9(1):197– 203, Spring 2004. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Wang:2007:QCS

[WPK^{+}07] Yalou Wang, Tony Putman, Nidal N. Khrais, Ching-Roung Chou, Gopal Kumar, and Freddy Goh. QoS control for streaming service in UMTS networks. Bell Labs Technical Journal, 11 (4):185–200, Winter 2007. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Wahl:2009:SIA

[WRL^{+}09] Stefan Wahl, Konrad Rieck, Pavel Laskov, Peter Domschitz, and Klaus-Robert Müller. Securing IMS against novel threats. Bell Labs Technical Journal, 14(1): 243–257, Spring 2009. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Ward:2006:DMI

[WSS^{+}06] Denise M. Ward, David P.

Strand, Wayne A. Senneke, Barbara L. Knolinski, Randall J. Scheer, and Alan J. McBride. Domain management of IMS. Bell Labs Technical Journal, 10(4): 233–254, Winter 2006. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Weldon:2008:NGA

[WvLS08] Marcus K. Weldon, Thierry van Landegem, and Edward S. Szurkowski. Nextgeneration access networks: A preview. Bell Labs Technical Journal, 13(1): 1–10, Spring 2008. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Wang:2008:CTO

[WXT^{+}08] Qingwen Wang, Cao Xingang, Zhong Tao, Jin Qianfu, and Wang Bin. China Telecom operators: Applications platform overview. Bell Labs Technical Journal, 13 (2):223–235, Summer 2008. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Weldon:2003:EFH

[WZ03] Marcus K. Weldon and Francis Zane. The economics of fiber to the home revisited. Bell Labs Technical Journal, 8(1):181–206, Spring 2003. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Xu:2001:GMB

[XLVN01] Yangguang Xu, Patrice N. Lamy, Eve L. Varma, and Ramesh Nagarajan. Generalized MPLS-based distributed control architecture for automatically switched transport networks. Bell Labs Technical Journal, 6(1):33– 49, Spring 2001. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Yan:2002:MCD

[Yan02] Bolin Yan. A multi-criteria decision-making model and its application in system solution engineering. Bell Labs Technical Journal, 7(1):197– 214, Spring 2002. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Yue:2006:ARO

[YJY06] Chen Wen Yue, Yu Jie, and Jiang Wen Yuan. Achieving reliability in OEM/ODM CPE products. Bell Labs Technical Journal, 11(1): 209–213, Spring 2006. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Zivkovic:2005:AAH

[ŽBLvB05] Miroslav Živković, Milind M.

Buddhikot, Ko Lagerberg, and Jeroen van Bemmel. Authentication across heterogeneous networks. Bell Labs Technical Journal, 10(2):39– 56, Summer 2005. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Zenner:2004:SIP

[Zen04] Guy J. Zenner. Session Initiation Protocol. Bell Labs Technical Journal, 9(3): 1–3, Autumn 2004. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Zenner:2003:EUS

[ZJP03] Guy J. Zenner, Mark H.

Jones, and Amit A. Patel. Emerging uses of SIP in service provider networks. Bell Labs Technical Journal, 8(1): 43–63, Spring 2003. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Zivkovic:2008:PDS

[ŽKN^{+}08] Miroslav Živković, Gerhard Kramer, Carl J. Nuzman, Carl Posthuma, James Wheeler, Phil Whiting, and Adriaan J. de Lind van Wijngaarden. Performance of digital subscriber line spectrum optimization algorithms. Bell Labs Technical Journal, 13 (1):129–146, Spring 2008. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Zhang:2008:IDI

[ZMS08] Peng Zhang, Juntao Ma, and Xiaorong Sun. Intelligent delivery of interactive advertisement content. Bell Labs Technical Journal, 13 (3):143–158, Autumn 2008. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Zeng:2008:EVS

[ZS08] Yugu Zeng and Thomas Strauss. Enhanced video streaming network with hybrid P2P technology. Bell Labs Technical Journal, 13 (3):45–58, Autumn 2008. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Zysman:2000:TEM

[ZTH^{+}00] George I. Zysman, Joseph A. Tarallo, Richard E. Howard, John Freidenfelds, Reinaldo A. Valenzuela, and Paul M. Mankiewich. Technology evolution for mobile and personal communications. Bell Labs Technical Journal, 5(1): 107–129, Spring 2000. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Zitouni:2004:MSS

[ZZLD04] Imed Zitouni, Qiru Zhou, Minkyu Lee, and Peter J. Danielsen. Media services in SIP networks. Bell Labs Technical Journal, 9(3):73– 85, Autumn 2004. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).